### v1

# Representative Action Selection for Large Action-Space Meta-Bandits

Abstract
We study the problem of selecting a subset from a large action space shared by a family of bandits, with the goal of achieving performance nearly matching that of using the full action space. We assume that similar actions tend to have related payoffs, modeled by a Gaussian process. To exploit this structure, we propose a simple $\epsilon$-net algorithm to select a representative subset. We provide theoretical guarantees for its performance and compare it empirically to Thompson Sampling and Upper Confidence Bound.
1 Introduction
We study a family of bandits that share a common but extremely large action space. We aim to understand whether it is possible—and how—to select a smaller set of representative actions that performs nearly as well as the full action space across all bandit instances. To build intuition, imagine a pharmacy preparing its inventory for the upcoming season. The available drugs (actions) are nearly infinite, and each customer (bandit) has unique characteristics. If two drugs share similar ingredients, their effects on a patient are likely to be similar. Likewise, if two patients have comparable health indices, a drug is likely to have similar effects on both. By modeling the expected outcome of each drug for each patient as a Gaussian process, we can capture these correlations. Now, consider medicine demand: If a drug treats a very rare illness, it is unlikely to be needed frequently, so the store can exclude it to optimize storage space. Conversely, if flu season is approaching, stocking several flu medications is a wise choice.
Different from prior approaches in multi-armed bandits (MAB) (dani2008stochastic, ) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits. This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
Consider the following MAB setting: In a bandit, if a decision-maker plays an action $a\in\mathcal{A}_{full}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$. We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle,$ where this bandit instance $\theta\in\mathbb{R}^{n}$ follows a multivariate Gaussian distribution. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ forms a Gaussian process (vershynin2018high, ). Our underlying setting aligns with the one considered in contextual bandit (dani2008stochastic, ; bietti2021contextual, ); however, the key difference lies in the objective.
Now consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{full}$. For a given bandit instance $\theta$, if the best action over the full space lies within $\mathcal{A}$, the decision-maker benefits from a reduced set of suboptimal actions to explore. Conversely, if the best action lies outside $\mathcal{A}$, regret arises from being unable to select this best action. Thus, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
$\textnormal{Regret}:=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$ (1)
which depends on the sampled bandit instance and is therefore a random variable. Our objective is to identify a small subset $\mathcal{A}$ that minimizes the expected regret $\mathbb{E}_{\theta}\left[\operatorname*{Regret}\right]$ over all possible bandit instances, making the underlying optimization both stochastic and combinatorial.
This objective is motivated by practical considerations. MAB regret in finite-horizon settings (auer2002using, ; auer2002nonstochastic, ) typically scales with the number of available actions. However, if a subset $\mathcal{A}$ is carefully chosen, the resulting MAB regret can be lower than when using the full action space, due to a reduced set of suboptimal actions. This is especially beneficial when there are a very large number of actions, as even the initialization phase can be computationally expensive. This improvement arises from a natural connection between the regret in Equation (1) and MAB regret, as discussed in Supplementary A.
The remainder of this section is literature review. The subsequent sections present our main contributions and are organized as follows:
  • Section 2: Subset Selection Framework and Algorithm. We introduce a framework that models expected outcome using a Gaussian process to capture correlations across similar actions and bandit instances. Based on this structure, we construct geometric $\epsilon$-nets and refine them using an importance-based selection mechanism. This mechanism is formalized as a variant of the $\epsilon$-net algorithm (haussler1986epsilon, ), tailored to the bandit setting.
  • Section 3: Regret Analysis. We provide theoretical guarantees, including upper and lower bounds on the expected regret that reflect the geometric and importance structure of the action space.
  • Section 4: Examples and Empirical Validation. We extend our analysis to settings where outcome (payoff) functions are sampled from a Gaussian process with a kernel, and empirically compare the performance to that of Thompson Sampling and Upper Confidence Bound.
The code used for the experiments is available at https://github.com/Quan-Zhou/MAB.
1.1 Related Works
Multi-Armed Bandits The classic MAB problem (lai1985asymptotically, ; auer2002finite, ) is defined by a set of arms, each deliver outcomes that are independently drawn from a fixed and unknown distribution. The decision-maker sequentially selects an arm, observes its outcome, and aims to maximize cumulative outcomes over time. The policy is to choose the next arm based on the sequence of past plays and obtained outcomes. Popular policies include the Upper Confidence Bound (auer2002finite, ), Thompson Sampling (agrawal2012analysis, ), and EXP3 (auer2002nonstochastic, ).
Contextual Bandits Unlike classic MAB, which assumes action outcomes are independent, contextual bandits (auer2002using, ) incorporate observed features (contexts) at each round. The decision-maker selects actions based on context, aiming to learn a policy that maps contexts to high-reward actions. Linear models such as LinUCB (chu2011contextual, ) and linear Thompson sampling (agrawal2013thompson, ) assume linear reward functions, enabling efficient learning. Early work includes LinRel (auer2002using, ), while later efforts extended to nonlinear models (filippi2010parametric, ; jun2017scalable, ). Nonlinear payoff structures have also been explored using Gaussian processes (srinivas2009gaussian, ) and RKHS methods (valko2013finite, ; chowdhury2017kernelized, ). Recent studies examine richer structures: budget-constrained linear contextual bandits (agrawal2016linear, ), reward feedback across contexts (balseiro2019contextual, ), and adversarially corrupted outcomes (he2022nearly, ). Empirical evaluations of common algorithms appear in bietti2021contextual .
Online Combinatorial Optimization Several works have extended stochastic linear optimization (dani2008stochastic, ; rusmevichientong2010linearly, ) to combinatorial settings. This line of work applies the Follow-the-Perturbed-Leader (FPL) algorithm (hannan1957approximation, ; kalai2005efficient, ), which introduces random perturbations to the estimated cumulative loss of each action, resulting in simple and efficient algorithms for various combinatorial structures. This idea has also been extended to online convex optimization. For instance, (hazan2007logarithmic, ) develop an FPL-style algorithm that achieves logarithmic regret for strictly convex payoff functions, particularly in settings where the decision point lies in the simplex. Follow-the-Regularized-Leader (FTRL) (shalev2006convex, ; shalev2012online, ) generalizes Follow-the-Leader by adding a regularization term to stabilize updates. (mcmahan2011follow, ) show that many online learning algorithms, including mirror descent, can be interpreted as instances of FTRL. In the context of online combinatorial optimization, (audibert2011minimax, ; audibert2014regret, ) study minimax regret bounds and propose FTRL-based algorithms alongside other strategies. (ito2021hybrid, ) further develop a hybrid algorithm based on FTRL.
Combinatorial Bandits Combinatorial bandits (cesa2012combinatorial, ) are a subclass of online combinatorial optimization where the decision-maker selects a subset of base arms from $\mathcal{A}_{full}$ in each round, forming a super arm. Early work on online shortest path problems with adversarial edge weights (gyorgy2007line, ; uchiya2010algorithms, ; liu2012adaptive, ) adapts the EXP3 framework (auer2002nonstochastic, ). Subsequent studies establish regret bounds for Thompson Sampling in structured settings (gopalan2014thompson, ) and propose general models with non-linear rewards and semi-bandit feedback (chen2016combinatorial, ; kveton2015tight, ). The matroid bandit model (kveton2014matroid, ) captures independence constraints via matroid optimization (whitney1992abstract, ). Later work reframes the shortest path problem as an adversarial MDP (rosenberg2019online, ) and introduces improved UCB variants with adaptive confidence estimation (zhou2024towards, ).
Epsilon net We identified two definitions of $\epsilon$-nets, with further discussion in Section 2.1. The first net of a given set, referring to (vershynin2018high, ), ensures that closed balls of radius $\epsilon$ centered at net points cover the given set. This relates to the covering number, that is the smallest possible cardinality of an $\epsilon$-net. This concept extends to function classes as a complexity measure, known as metric entropy, with further details in wainwright2019high , and has been used in russo2013eluder on the class of outcome functions. The second definition, as given in (matousek2013lectures, ), relates to importance sampling, requiring that every significant subset of a set intersects with the net. The classic $\epsilon$-net algorithm, introduced by haussler1986epsilon , operates via random sampling, yielding a net of size $\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\epsilon})$. As noted in matousek2013lectures , the importance measure extends to general probability measures, inspiring our approach. Further works aim to reduce net size (mustafa2019computing, ; pach2011tight, ; rabani2009explicit, ) and address online settings (bhore2024online, ).
2 Subset Selection Framework
We consider the problem of selecting a small number of representative action from a large (potentially infinite but compact) action space, $\mathcal{A}_{full}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$. Define the expected outcome of playing action $a$ in bandit instance $\theta$ as
$\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{full},\theta\sim\mathcal{N}(0,I),$ (2)
where the bandit instance $\theta$ follows a multivariate normal distribution with zero mean and identity covariance matrix $I$. Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ forms a canonical Gaussian process (vershynin2018high, ). In fact, any constant-mean Gaussian process can be represented as the canonical process; see Supplementary B for details.
Let $\operatorname*{Conv}(\mathcal{A}_{full})$ denote the convex hull that consists of all convex combinations of finitely many actions in $\mathcal{A}_{full}$. The extreme points of the convex hull are those points $x\in\operatorname*{Conv}(\mathcal{A}_{full})$ for which there do not exist distinct points $a,a^{\prime}\in\operatorname*{Conv}(\mathcal{A}_{full})$ and $\lambda\in(0,1)$ such that $x=\lambda a+(1-\lambda)a^{\prime}$.
By the extreme point theorem, i.e., the optimum of the linear function on a convex set is attained at an extreme point—it follows that the optimal action for any bandit instance is always an extreme point of the full action space. Therefore, if we use all the extreme points as representatives of the full action space, the regret (as defined in Equation (1)) is zero. See Supplementary C for details.
This example highlights how a geometric approach can be used to solve the stochastic combination problem. However, even if one only needs the extreme points, the set of extreme points may still be large or even infinite, e.g, the extreme points of a Euclidean ball is infinite. To address this, we will later introduce the notions of $\epsilon$-nets. Without loss of generality, we assume $\mathcal{A}_{full}$ consists only of the extreme points of $\operatorname*{Conv}(\mathcal{A}_{full})$, as they are the only points of interest.
2.1 Epsilon Nets
There are at least two definitions of $\epsilon$-net: one from a geometric perspective (vershynin2018high, ) and another from a measure-theoretic perspective (matousek2013lectures, ). Let $\|\cdot\|_{2}$ denote the Euclidean norm.
  • A subset $\mathcal{A}\subseteq\mathcal{A}_{full}$ is called a Geometric $\epsilon$-net if, for all $a\in\mathcal{A}_{full}$, there exists $a^{\prime}\in\mathcal{A}$ such that
    $\|a-a^{\prime}\|_{2}<\epsilon.$
  • Let $\mathcal{R}$ be a finite partition of the extreme points into disjoint clusters such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{full}$. Given a measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. A subset $\mathcal{A}\subseteq\mathcal{A}_{full}$ is called a Measure-Theoretic $\epsilon$-net with respect to measure $q$ if, for any cluster $r\in\mathcal{R}$, we have:
    $r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
Intuitively, the geometric definition ensures point-wise closeness, while the measure-theoretic definition ensures representative sampling. In well-behaved spaces, the two coincide. For example, let $\mathcal{A}_{full}=[0,1]$ and $q$ be the uniform distribution over this interval. A grid with spacing $\epsilon$ forms a geometric $\epsilon$-net and also ensures any subinterval of measure at least $\epsilon$ contains at least a grid point.
Both definitions are useful. A geometric $\epsilon$-net ensures small regret, as it guarantees that an action close to the optimal one always exists in $\mathcal{A}$. However, it suffers from the curse of dimensionality: for $[0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$, exponentially in $n$. To address this, we incorporate the measure-theoretic $\epsilon$-net to restrict the geometric construction to only the important clusters—as determined by the $q$-measure.
The partition $\mathcal{R}$ is the central concept that bridges the two definitions of $\epsilon$-nets. If two actions $a,a^{\prime}$ are in the same cluster, they are close in the Euclidean sense, and the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., in expected squared difference). This holds because, for the canonical Gaussian process, the increments of the process correspond to the Euclidean distance:
$\|\mu_{a}-\mu_{a^{\prime}}\|_{L^{2}}=\left(\mathbb{E}(\mu_{a}-\mu_{a^{\prime}})^{2}\right)^{1/2}=\left(\mathbb{E}(a-a^{\prime})^{T}\theta\theta^{T}(a-a^{\prime})\right)^{1/2}=\|a-a^{\prime}\|_{2},\;\forall a,a^{\prime}\in\mathcal{A}_{full},$
where $\theta^{T}$ denotes the transpose of $\theta$. The middle & right equalities use Equation (2) and $\mathbb{E}\theta\theta^{T}=I$. Then, selecting one point from each cluster $r\in\mathcal{R}$ forms a geometric $\epsilon$-net (which will later be formally defined as a reference subset), where $\epsilon$ is the maximum diameter of all the clusters. This construction yields bounded expected regret, as each selected point serves as a representative for its cluster, and all other points within the same cluster have bounded deviation from the representative. Furthermore, clusters in $\mathcal{A}_{full}$ that are deemed unimportant for expected regret can be ignored: even if no representative points are selected from these clusters, the resulting expected regret will not increase significantly compared to that of the reference subset.
2.2 Epsilon Net Algorithm
We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by haussler1986epsilon . It selects $K$ i.i.d. random actions, aligned with the distribution of bandit instances. Since repetitions are allowed, the resulting subset $\mathcal{A}$ may have fewer than $K$ distinct actions.
Algorithm 1 Epsilon Net Algorithm
1:Input: Action space $\mathcal{A}_{full}$, Sample size $K$, Distribution of bandits.
2:Output: A subset of actions $\mathcal{A}$.
3:$\mathcal{A}\leftarrow\emptyset$
4:for $1,\dots,K$ do
5:     Sample a bandit instance $\theta$.
6:     Find the optimal action $a^{*}(\theta):=\operatorname*{argmax}_{a\in\mathcal{A}_{full}}\langle a,\theta\rangle$.
7:     $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$triangleright Repetitions of actions are allowed
8:end for
We define the optimal action in a bandit instance $\theta$ as $a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{full}}\mu_{a}(\theta)$.
Assumption 1 (Unique optimal action).
The optimal action $a^{*}(\theta)$ is unique with probability $1$ for all bandit instances.
This assumption does not exclude the case where the action space is a polytope, and there may be specific bandit instances for which the optimal actions lie on the edges, resulting in an infinite number of optimal actions. Since the bandit instances follow a continuous distribution, the probability of encountering these specific bandits is zero.
Then, define the measure $q$:
[…]
Assumption 2 (Compact support).
the support of measure $q$ is compact.
The compactness assumption ensures that $\mathbb{E}_{\theta}[\max_{a\in\mathcal{A}_{full}}\mu_{a}]$ is finite and that the unique optimal action exists. Without loss of generality, let $\mathcal{A}_{full}$ be the support of measure $q$.
By definition, Algorithm 1 samples $K$ i.i.d. extreme points from clusters in $\mathcal{R}$ according to the probability measure $q$. If a cluster has a higher measure under $q$, its elements are more likely to be included in the output of Algorithm 1. In fact, with high probability, Algorithm 1 outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{full}$ with respect to measure $q$. (This is a simplified version of the $\epsilon$-net theorem; see Theorem 10.2.4 of (matousek2013lectures, ).)
Lemma 3 (Measure-theoretic $\epsilon$-net).
Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ samples. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, it holds that for any cluster $r\in\mathcal{R}$,
$r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
3 Regret Analysis
In this section, we begin by analyzing a special class of geometric $\epsilon$-nets, constructed by partitioning the action space into clusters and selecting a single representative action from each cluster. We then extend the analysis to obtain algorithm-dependent bounds for the output of Algorithm 1.
For any cluster $r\in\mathcal{R}$, we define its diameter as $\operatorname{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$. Throughout the following, we set $\epsilon$ to be the maximum diameter over all clusters, i.e., $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
Definition 1 (Reference subsets).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$, and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
We impose the following assumption on the partition, used only for lower bounds. This means that when the optimal action falls within an arbitrary cluster $r$, any action in this cluster is better than any action outside of it. It ensures that the clusters are relatively well-separated.
Assumption 4 (Effective partition).
For any $r\in\mathcal{R}$, whenever the optimal action lies in cluster $r$, i.e., $a^{*}(\theta)\in r$, then $\mu_{a}\geq\max_{a^{\prime}\in\mathcal{A}_{full}\setminus r}\mu_{a^{\prime}},\;\forall a\in r$.
Theorem 5 (Regret bounds of reference subsets).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, there is an absolute constant $C>0$, such that
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
If the partition $\mathcal{R}$ satisfies Assumption 4, then
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
Proof sketch. For each cluster $\ell\leq m$, define a simple Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, regret is upper bounded by $Y_{\ell}$ (or equal to it under Assumption 4). Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$. Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded using concentration property of Gaussian process. $\square$
Theorem 5 reveals that the expected regret grows linearly with the cluster diameter but logarithmically with the number of clusters. If the action space has a large overall diameter but exhibits a clustered structure—that is, actions are naturally grouped into a few tight clusters—then the reference subset can achieve significantly lower regret. In contrast, if the action space has a scattered or uniform structure, using a moderate number of clusters would lead to large cluster diameters, which in turn increases the regret of the reference subset.
As shown in the next section, Theorem 5 paves the way for analyzing Algorithm 1 by bounding its expected regret in terms of a reference subset, plus an additional term—since the proof relies on the possibility that its output matches a reference subset. Thus, the structure-regret relationship applies to Algorithm 1’s output.
In Figure 1, we study the effect of the cluster diameters (controlled by a spread parameter) on regret. We structure the clustered action space: Five center points are fixed on the unit sphere in $\mathbb{R}^{3}$, and around each center, 200 points are sampled to form five clusters. Each point is obtained by adding Gaussian noise (mean zero, standard deviation equal to the spread parameter) to the center direction, followed by projection back onto the unit sphere. Bandits are sampled from a 3-dimensional standard Gaussian distribution, i.e., $\theta\sim\mathcal{N}(0,I)$. The left subplot shows the expected regret of Algorithm 1 with $K=10$, computed using $10^{4}$ additional bandits, as the spread varies from 0.01 to 0.5. The curves and error shade represent the mean $\pm$ one standard deviation of expected regret over 30 repetitions. The middle and right subplots display example action spaces for spread values of 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
Figure 1: Illustration of clustered action spaces on unit sphere in $\mathbb{R}^{3}$ and the effect of cluster diameters on regret. Five clusters are formed by generating 5 fixed center points, with 200 points sampled around each using Gaussian noise (spread controls the variance). Bandits are drawn from $\mathcal{N}(0,I)$. The left subplot shows the mean $\pm$ standard deviation of the expected regret (over 30 trials) as the spread varies from 0.01 to 0.5, using $10^{4}$ additional bandits. The middle and right subplots show example action spaces (blue dots) for spread values 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
3.1 Regret Bounds of Algorithm
The expected regret in previous results is taken over bandit instances $\theta$. In contrast, since the output of Algorithm 1 is random, the expected regret analyzed in this section is taken with respect to both the algorithm’s randomness (i.e., the sampled $\mathcal{A}$) and the distribution over $\theta$.
The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds. The key difference is that the algorithm may omit some clusters, whereas the reference subset includes a representative from each cluster. However, although it may select a smaller subset, the algorithm achieves regret comparable to that of the reference subset, as it tends to omit clusters that contribute minimally to the expected regret.
Theorem 6 (Upper bound).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be output of Algorithm 1. For the same constant $C>0$ in Theorem 5,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof sketch. If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, choose $a_{\ell}\in r_{\ell}\cap\mathcal{A}$ as the representative point. If $r_{\ell}\cap\mathcal{A}=\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$. The new set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset. Then, for each $\ell\leq m$, define a simple Gaussian process $\{Z_{a}\}_{a\in r_{\ell}}$, where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and let $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, we consider two cases:
  • If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, the regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
  • If $r_{\ell}\cap\mathcal{A}=\emptyset$, the regret is still bounded by $\max_{a\in\mathcal{A}_{full}}\mu_{a}$. This is justified by the assumption of $0\in\mathrm{conv}(\mathcal{A})$, since $\mathbb{E}_{\theta}[\max_{a\in\mathcal{A}}\mu_{a}]$ is invariant under affine unitary transformations of $\mathcal{A}$. $\square$
Comparing Theorem 6 with Theorem 5, we observe that the algorithm-dependent upper bound exceeds that of the reference subset by an additional term:
$\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.$
Note that the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ behaves similarly to entropy. To illustrate, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$. This expression is small when $q(r)$ is small, and decays rapidly when $q(r)$ is large, making the overall term negligible when $q$ is concentrated on few clusters.
In general, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified. We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{full},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{full}$:
$N(\mathcal{A}_{full},\epsilon)=\min\left\{m\in\mathbb{N}:\exists\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{full},\ \forall a\in\mathcal{A}_{full},\ \exists\ell,\ \|a-a_{\ell}\|_{2}\leq\epsilon\right\}.$
Theorem 7 (Partition-independent upper bound).
Under Assumption 2, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{full}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 5, and another absolute constant $c>0$,
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)},\quad\text{where }K\geq c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{full},\epsilon).$
As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
The partition-independent upper bound in Theorem 7 depends on a positive constant $\epsilon$, which serves as a tolerance parameter up to the diameter of the action space. Given a $\epsilon$, the required number of samples $K$ scales with the square of the diameter-to-$\epsilon$ ratio, multiplied by the covering number—a geometric property of the action space. The resulting expected regret is then bounded in terms of $\epsilon$, the dimensionality, and the logarithm of the covering number. As $\epsilon$ decreases, more samples are needed, but the regret bound becomes tighter.
Further, we give a corresponding lower bound of Algorithm 1:
Theorem 8 (Lower bound).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 4. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C$ in Theorem 5 and another constant $c>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).\end{split}$
4 Examples
In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings. Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling. We therefore adopt a dimension-free view by treating the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ as a family of random variables indexed by an abstract set $\mathcal{A}_{full}$. In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{full}$:
$f(a):=\mu_{a},\quad\forall a\in\mathcal{A}_{full}.$ (4)
[…]
$k(a,a^{\prime})=\sum\nolimits_{i\leq\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(a^{\prime}),$
where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues. Let outcome $f$ be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{full}$ and a weight vector $\alpha\in\mathbb{R}^{N}$. As shown in (williams2006gaussian, ; wainwright2019high, ), the function $f$ can also be written as
$\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
where $\mathbf{f}$ is a vector of coefficients, and $\Phi(a)$ usually refereed as a feature map, has entries $\Phi_{i}(a)=\sqrt{\lambda_{i}}\phi_{i}(a)$. If the coefficients $\{\mathbf{f}_{i}\}_{i=1}^{\infty}$ are i.i.d. Gaussian, this formulation can be represented as the canonical Gaussian process model in Equation (2). Note that in this case, the actual action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{full}$.
Limitations: Gaussian assumptions may limit applicability in practice. We plan to extend this work to general random process with sub-Gaussian property.
4.1 Varying-dependence Actions
We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel (williams2006gaussian, ).
[…]
where $l$ is a length-scale parameter and $l(a):=0.1+0.9\cdot\exp(-\|a\|^{2})$ is a location-dependent length-scale function. Both of them control the dependence between actions. But, unlike the stationary RBF kernels, the Gibbs kernel allows some regions of the action space to be highly correlated over large distances, and others to have short-range correlations. That is, the correlation depends not only on the distance between actions, but also on where $a,a^{\prime}$ are located. When $l(a)=l$ is a constant, the Gibbs kernel reduces to the RBF kernel.
Sampling Outcome Functions from a Kernel: We first construct the kernel matrix $\mathbf{K}$, where each entry is given by $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$, for $a,a^{\prime}\in\mathcal{A}_{full}$, depending on the choice of kernel. We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$. Under either kernels defined in Equation (5), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{full}$.
Figure 2: Experiments with outcome functions sampled from RBF/Gibbs kernels in Equation (5). Left: Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in $[0,2]$ (blue curves, right y-axis). The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points. Right: Performance on solving stochastic combinatorial optimization in Equation (6), selecting $K=5$ distinct actions from 15 grid points in [0,2], with outcome functions sampled from RBF kernel at varying length-scales. Solid curves and shaded regions represent the mean and $\pm$ standard deviation of the expected regret over 50 repetitions. Results are presented for our method (green), Thompson Sampling (yellow), and UCB (orange), with TS/UCB run for 3000 rounds per repetition.
To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval $[0,2]$, using Gibbs kernel in Equation (5) to sample outcome functions. To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid. The left subplot of Figure 2 provides examples of sampled outcome functions (blue curves, with the y-axis on the right-hand side), which become smoother as the actions approach the left end of the interval—indicating stronger correlations among function values in that region. We run Algorithm 1 on this action space with $K=5000$ to select actions and record the frequency of each action being selected. The resulting histogram (purple bars, with the y-axis on the left) reflects the importance measure $q$, highlighting that Algorithm 1 tends to select more actions from regions where the outcome functions are rougher—i.e., where action outcomes are less correlated and their features $\Phi(a)$ are are farther apart. Another interesting aspect of this subplot is the two high bars at the edges. Recall that the actual action space consists of feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{ful}$. For actions indexed closer to 0, their feature vectors become more densely packed compared to those indexed closer to 2, resulting in more correlated outcomes. The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
As an extreme case, suppose $\mathcal{A}_{full}$ is a set of independent actions. This corresponds to the canonical process with the action space given by the orthogonal basis in $\mathbb{R}^{n}$ where $n=|\mathcal{A}_{full}|$. To see this, we associate each action $a$ with a unit vector $e_{a}$, which is $1$ at the $a^{th}$ entry but is zero otherwise, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$. With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ consists of mutually independent random variables. Then, as discussed in Supplementary I, for any subset $\mathcal{A}\subset\mathcal{A}_{full}$, including the output of Algorithm 1, the regret suffers from the lower bound
$\mathbb{E}_{\theta}[\textnormal{Regret}]\geq\frac{\sqrt{\log|\mathcal{A}_{full}|}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}.$
As a result, any algorithm, including Algorithm 1, would perform poorly unless the subset size $|\mathcal{A}|$ is sufficiently large.
4.2 Combinatorial Optimization under Uncertainty
Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in Equation (1). Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{full}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
$\max_{\mathcal{A}\subseteq\mathcal{A}_{full}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (6)
In general, this problem is non-trivial because the objective function involves an expectation over a collection of random variables whose distribution is neither specified nor directly accessible. Moreover, the expression inside the expectation is the maximum over a subset of potentially correlated random variables. As a result, small changes in $\mathcal{A}$ can lead to non-smooth variations in the maximum.
Our method is computationally efficient, avoiding the combinatorial complexity of the optimization problem in Equation (6). To assess the effectiveness in identifying near-optimal solutions, we compare it against Thompson Sampling (TS) and Upper Confidence Bound (UCB) on a simple example: We consider a fixed action space consisting of 15 grid points in $[0,2]$, using an RBF kernel in Equation (5) to sample outcome functions while varying the length-scale $l\in\{0.5,1,\dots,4\}$. In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where in each round, the optimal action is founded via an exhaustive search over the action space. In TS/UCB methods, we treat each $K$-tuple of actions as a super arm and adopt a bandit feedback setting: at each round, the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}$, and updates its policy accordingly, repeating this process for 3000 rounds. Thus, they are tailored to maximize the same objective. Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB, with the algorithms summarized in Supplementary J, along with examples of selected subsets and sampled outcome functions. We evaluate expected regret of all methods over $10^{5}$ randomly sampled outcome functions.
The right subplot of Figure 2 reports the expected regret of our method (green), TS (yellow) and UCB (orange), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions. As shown in this subplot, the expected regret decreases as the length-scale increases. This is because the length-scale $l$ of the RBF kernel controls the number of approximately independent actions in the process $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$: When actions $a$ and $a^{\prime}$ are far apart relative to the length-scale, the kernel value becomes very small, so the covariance between the function values $\mu_{a}$ and $\mu_{a^{\prime}}$ is also close to zero.
Appendix A Connection to Bayesian Regret of Multi-Armed Bandits
Given a subset $\mathcal{A}\subseteq\mathcal{A}_{full}$, let $\mu^{*}(\mathcal{A},\theta):=\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ be the highest expected outcome in a bandit instance $\theta$. We can equivalently rewrite the regret in Equation (1):
$\textnormal{Regret}(\theta)=\mu^{*}(\mathcal{A}_{full},\theta)-\mu^{*}(\mathcal{A},\theta).$
Bandit Regret: In the classic multi-armed bandit setting [7, 32], outcome of an arm $a\in\mathcal{A}_{full}$ in a bandit instance $\theta$ is generated i.i.d. according to an unknown but fixed law with expectation $\mu_{a}(\theta)$. A decision-maker would follow a policy $\pi$ that chooses the next arm based on the sequence of rounds and obtained outcomes. The bandit regret of policy $\pi$ in this bandit environment $\theta$ over $N$ rounds is defined as the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action and following a policy $\pi$ to choose actions. Bayesian (bandit) regret [2, 32] is the expectation of this bandit regret, taken over the randomness in the distribution of bandit instances, and the action chosen by a policy:
$\begin{split}\textnormal{BanditRegret}(\theta):=&\mu^{*}(\mathcal{A}_{full},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta),\\ \textnormal{BayesianRegret}:=&\mathbb{E}\left[\mu^{*}(\mathcal{A}_{full},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta)\right],\end{split}$
where $A_{t}$ is the action chosen by policy $\pi$ in round $t$. Since we consider settings where only a subset of the full action space is accessible, the action chosen by the policy can only from the subset $\mathcal{A}\subseteq\mathcal{A}_{full}$. This quantifies the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action from the full action set $\mathcal{A}_{full}$ and following policy $\pi$ to choose actions from the restricted subset $\mathcal{A}$.
Next, we establish the connection between the regret defined in this paper and the Bayesian (bandit) regret.
A.1 Optimal policy
If $\pi$ is the optimal policy, it selects the best action in $\mathcal{A}$ in every round. Therefore, the regret of only having access to $\mathcal{A}$ instead of $\mathcal{A}_{full}$ for $N$ rounds, is determined by the optimal actions in $\mathcal{A}$ and $\mathcal{A}_{full}$:
[…]
A.2 Thompson sampling policy
If $\pi$ is the Thompson sampling policy, the Bayesian regret is bounded by
[…]
$=$ $\mathbb{E}\left[\sum_{t=1}^{N}\left(\mu^{*}(\mathcal{A})-\mu_{A_{t}}+\mu^{*}(\mathcal{A}_{full})-\mu^{*}(\mathcal{A})\right)\right]$
$\leq$ $C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}],$
where $|\mathcal{A}|$ denotes the cardinality, and $C>0$ is a constant. The inequality uses the Bayesian (bandit) regret of Thompson sampling (Theorem 36.1 in [32]). Note that with access to the full action space, the Bayesian regret is bounded by $C\sqrt{|\mathcal{A}_{full}|\cdot N\log N}$. However, this can be worse than restricting to a subset $\mathcal{A}$ when $\mathbb{E}_{\theta}[\textnormal{Regret}]$ is sufficiently small and the time horizon $N$ is finite. Intuitively speaking, using a well-chosen subset of representative actions instead of the full action space can reduce Bayesian regret in multi-armed bandits.
Appendix B Canonical Representation of a Gaussian Process
The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, such that:
$\mu_{a}:=\langle a,g\rangle+c,\quad\;\forall\;a\in\mathcal{A}_{full},\;g\sim\mathcal{N}(0,\Sigma).$ (7)
We show how to transform it into the canonical Gaussian process. Note that
$g\equiv\Sigma^{1/2}\theta,\;\theta\sim\mathcal{N}(0,I),$
where $\equiv$ denotes equivalence in distribution. Let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$. Thus, $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$, and
$\mu_{a}=\langle a,g\rangle=\sum_{j\leq n}a_{j}g_{j}=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
Therefore, the setting in Equation (7) is equivalent to
$\mu_{a}=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{full},\;\theta\sim\mathcal{N}(0,I),$
where $\Sigma^{1/2}\mathcal{A}_{full}$ is the projected set of $\mathcal{A}_{full}$ by the matrix $\Sigma^{1/2}$.
Appendix C Convex hull dominated by extreme points
Lemma 9.
Given a subset $\mathcal{A}\subset\mathcal{A}_{full}$. If $\mathcal{A}$ includes all extreme points of $\operatorname*{Conv}(\mathcal{A}_{full})$, i.e., $\mathcal{A}_{full}\subseteq\operatorname*{Conv}(\mathcal{A})$, then it always holds
$\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=0\;\;\forall\theta\in\mathbb{R}^{n}.$
Proof.
Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{K}\}$. Since $\mathcal{A}_{full}\subseteq\operatorname*{Conv}(\mathcal{A})$, for any $a\in\mathcal{A}_{full}$, there exist non-negative numbers $\lambda_{1},\dots,\lambda_{K}$ that sum up to $1$ and such that $a=\lambda_{1}a_{1}+\dots+\lambda_{K}a_{K}$. Thus, by Equation (2), for any $\theta\in\mathbb{R}^{n}$
$\mu_{a}=\langle a,\theta\rangle=\left\langle\sum_{i\in 1:K}\lambda_{i}a_{i},\theta\right\rangle=\sum_{i\in 1:K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i\in 1:K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.$
It yields
$\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{full}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$
where the first inequality uses $\mathcal{A}\subset\mathcal{A}_{full}$. Thus, the two quantities are equal. ∎
Appendix D Proof of Lemma 3
Statement: Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ samples. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, it holds that for any cluster $r\in\mathcal{R}$,
$r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
Proof.
By definition, the algorithm draws $K$ i.i.d. samples from clusters in $\mathcal{R}$ according to the distribution $q$. Define a set of typical clusters $R_{\epsilon}:=\{r\in\mathcal{R}:q(r)>\epsilon\}$. Then,
$\Pr[r\cap\mathcal{A}\!=\!\emptyset]=(1-q(r))^{K}<(1-\epsilon)^{K}\leq\exp(-K\epsilon),\;\forall r\in R_{\epsilon},$ (8)
where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$. The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [46]. Therefore,
$\Pr[r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon]=$ $1-\Pr[r\cap\mathcal{A}\neq\emptyset\quad\exists r\in R_{\epsilon}]$
$\geq$ $1-\sum_{r\in R_{\epsilon}}\Pr[r\cap\mathcal{A}\neq\emptyset]$
$>$ $1-\sum_{r\in R_{\epsilon}}\exp(-K\epsilon)$
$=$ $1-|R_{\epsilon}|\exp(-K\epsilon)\geq 1-1/\epsilon\cdot\exp(-K\epsilon),$
where the first inequality uses union bound, the second inequality uses Equation (8). The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$. ∎
Appendix E Proof of Theorem 5
Lemma 10 (Regret decomposition of reference subsets).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space and an arbitrary reference subset $\mathcal{A}$. For each $\ell\leq m$, define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Under Assumption 4, $\operatorname*{Regret}=Y_{\ell}$ where the cluster $r_{\ell}$ contains the optimal action. Hence,
$\mathbb{E}_{\theta}\left[\operatorname*{Regret}\right]=\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}_{\theta}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right].$
Proof.
Let the cluster $r_{\ell}$ contains the optimal action, i.e., $a^{*}(\theta)\in r_{\ell}$, then
$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell},$
where the middle equality uses that when $a^{*}(\theta)\in r_{\ell}$, $\max_{a\in\mathcal{A}_{full}}\mu_{a}=\max_{a\in r_{\ell}}\mu_{a}$ and further with Assumption 4, $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\geq\mu_{a_{\ell}}$ when $a^{*}(\theta)\in r_{\ell}$. Then,
$\mathbb{E}\left[\operatorname*{Regret}\right]=\sum_{\ell\leq m}\Pr[a^{*}(\theta)\in r_{\ell}]\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]=\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],$
where the left equality uses tower rule, the right equality uses the definition of $q$ measure. ∎
To bound the expected regret of reference subsets, we need a few more lemmas:
Lemma 11 (Expectation integral identity).
Given a non-negative random variables $X$. If $\Pr[X\geq u]\leq c\exp{\left(-\frac{u^{2}}{\epsilon^{2}}\right)}$ for any $u>0$, then $\mathbb{E}X\leq C\epsilon\sqrt{\log c}$, where $\epsilon,c,C$ are positive constants.
Proof.
Let $u_{0}:=\epsilon\sqrt{\log c}$, then we have
[…]
$=$ $u_{0}+\exp\left(-\frac{u_{0}^{2}}{\epsilon^{2}}\right)\cdot\frac{c\epsilon^{2}}{2u_{0}}$
$=$ $\epsilon\sqrt{\log c}+\frac{\epsilon}{2\sqrt{\log c}}\leq C\epsilon\sqrt{\log c}.$
Lemma 12 (Borell-TIS inequality; Lemma 2.4.7 [45]).
Given a set $\mathcal{S}$, and a zero-mean Gaussian process $\left(X_{a}\right)_{a\in\mathcal{S}}$. Let $\epsilon:=\sup_{a\in\mathcal{S}}\left(\mathbb{E}X_{a}^{2}\right)^{\frac{1}{2}}$. Then for $u>0$, we have
$\Pr\left[\left|\sup_{a\in\mathcal{S}}X_{a}-\mathbb{E}\sup_{a\in\mathcal{S}}X_{a}\right|\geq u\right]\leq 2\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
It means that the size of the fluctuations of $\mathbb{E}\sup_{a\in\mathcal{S}}X_{a}$ is governed by the size of the individual random variables $X_{a}$.
Statement of Theorem 5:
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, for some constant $C>0$,
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
If the partition $\mathcal{R}$ satisfies Assumption 4, then
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
Proof.
Fix $\ell$, define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{full}$, it holds that $\mathbb{E}Y_{\ell}=\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}$.
When the cluster $r_{\ell}$ contains the optimal action, i.e., $a^{*}(\theta)\in r_{\ell}$, we have
$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell}\leq\max_{\ell^{\prime}\leq m}Y_{\ell^{\prime}},$
where the first equality follows from the definition of regret in Equation (1), and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$. The inequality holds because $a_{\ell}\in\mathcal{A}$, and hence $\max_{a\in\mathcal{A}}\mu_{a}\geq\mu_{a_{\ell}}$. The final equality follows from the definition of $Y_{\ell}$. On the other hand, according to Lemma 10, under Assumption 4, $\operatorname*{Regret}=Y_{\ell}$, and thus $\operatorname*{Regret}\geq\min_{\ell^{\prime}\leq m}Y_{\ell^{\prime}}$, where the cluster $r_{\ell}$ contains the optimal action.
Therefore, we reach the important conclusion that
$\min_{\ell\leq m}Y_{\ell}\leq\operatorname*{Regret}\leq\max_{\ell\leq m}Y_{\ell},$ (9)
where the left inequality holds under Assumption 4. Further, by definition $r_{\ell}\subset B(a_{\ell},\epsilon)$, such that $\mathbb{E}Z_{a}^{2}=\mathbb{E}(\mu_{a}-\mu_{a_{\ell}})^{2}=\|a-a_{\ell}\|^{2}_{2}\leq\epsilon^{2}$. Using Lemma 12 on the process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$, we have
$\Pr\left[\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq u\right]\leq 2\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
By union bound, we have
$\Pr\left[\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq u\right]\leq 2m\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
Using Lemma 11, we have for some absolute constant $C>0$
$\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\leq C\epsilon\sqrt{\log m}.$ (10)
Upper bound:
$\mathbb{E}\operatorname*{Regret}\leq\mathbb{E}\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\leq\max_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}+C\epsilon\sqrt{\log m},$ (11)
where the first inequality uses Equation (9). The second inequality uses $\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\leq\mathbb{E}Y_{\ell}+\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses Equation (10) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
Lower bound:
$\mathbb{E}\operatorname*{Regret}\geq\mathbb{E}\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq\min_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}-C\epsilon\sqrt{\log m},$ (12)
where the first inequality uses Equation (9) under Assumption 4. The second inequality uses $\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\geq\mathbb{E}Y_{\ell}-\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$. The last inequality uses Equation (10) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$. ∎
Appendix F Proof of Theorem 6
Lemma 13 (Transformation invariance).
Given any vector $x\in\mathbb{R}^{n}$, let $S+x:=\{s+x:s\in\mathcal{S}\}$. If $\mathbb{E}[\theta]=0$, then $\mathbb{E}\left[\max_{a\in\mathcal{S}+x}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
Proof.
$\mathbb{E}\left[\max_{a\in\mathcal{S}+x}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime}+x,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right]+\mathbb{E}\left[\langle x,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
where the last equality uses $\mathbb{E}\left[\langle x,\theta\rangle\right]=\left\langle x,\mathbb{E}[\theta]\right\rangle=0$. ∎
Lemma 14.
Consider a partition $\mathcal{R}$ of the full action space. Let $\mathcal{A}$ be the output of Algorithm 1. Then, the event that the optimal action falls in a cluster $r$, i.e., $\{a^{*}(\theta)\in r\}$, is independent of whether the subset $\mathcal{A}$ intersects with the cluster. It holds that
$\Pr[a^{*}(\theta)\in r,\,r\cap\mathcal{A}\!=\!\emptyset]=$ $q(r)(1-q(r))^{K},$
$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!\neq\!\emptyset]\leq$ $q(r).$
Proof.
Since $\mathcal{A}$ is the output of Algorithm 1, the event $\{a^{*}(\theta)\in r\}$ is independent from $\{r\cap\mathcal{A}\!\neq\!\emptyset\}$ or $\{r\cap\mathcal{A}\!=\!\emptyset\}$. We have
$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!=\!\emptyset]=\Pr[a^{*}(\theta)\in r]\Pr[r\cap\mathcal{A}\!=\!\emptyset]=q(r)\Pr[r\cap\mathcal{A}\!=\!\emptyset]=q(r)(1-q(r))^{K},$
where the first equality uses independence between $\{a^{*}(\theta)\in r\}$ and $\{r\cap\mathcal{A}\!\neq\!\emptyset\}$, the second equality uses the definition of measure $q$, the third equality use the probability of missing cluster $r$ in $K$ i.i.d. samplings. Similarly,
$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!\neq\!\emptyset]=q(r)\cdot\Pr[r\cap\mathcal{A}\!\neq\!\emptyset].$
Using $\Pr[r\cap\mathcal{A}\!\neq\!\emptyset]\leq 1$, we complete the proof. ∎
Statement of Theorem 6:
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 5,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof.
If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$. If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative. The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1. The cluster $r_{\ell}$ is contained in a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$. Define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{full}$, we have $\mathbb{E}Y_{\ell}=\mathbb{E}\sup_{a\in r_{\ell}}\mu_{a}$.
Consider the case that $a^{*}(\theta)\in r_{\ell}$. We have
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\leq&\mathbb{E}\left[Y_{\ell}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ =&\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\\ \leq&\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (13)
where the first inequality uses $\operatorname*{Regret}\leq Y_{\ell}$ if $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$, the equality uses that $Y_{\ell}$ is independent of $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$. On the other hand, we assume that $0\in\operatorname*{Conv}(\mathcal{A})$ because even if it doesn’t hold, we can always find a vector $x\in\mathbb{R}^{n}$ such that $0\in\operatorname*{Conv}(\mathcal{A}+x)$, without changing the value of $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ (c.f. Lemma 13). Therefore,
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\leq\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (14)
where the inequality uses $\operatorname*{Regret}\leq\max_{a\in\mathcal{A}_{full}}\mu_{a}$, as a consequence of $0\in\operatorname*{Conv}(\mathcal{A})$, and that $\max_{a\in\mathcal{A}_{full}}\mu_{a}$ is independent of $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
$\mathbb{E}[\operatorname*{Regret}]=$ $\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]$
$+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]$
$\leq$ $\sum_{\ell\leq m}q(r_{\ell})\cdot\left(\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]+(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\right)$
$=$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r\right]$
$\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}\max_{a\in r}\mu_{a}+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2},$
where the first equality uses tower rule. The first inequality uses Equations (13-14), and Lemma 14. The last equality uses tower rule again. The last inequality uses Equation (11) in the proof of Theorem 5 on the summand $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$. The other summand is bounded via Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$. ∎
Appendix G Proof of Theorem 7
Lemma 15.
Let $q$ denote a discrete probability distribution over a finite support $\mathcal{R}$. Define
$M:=\max_{q}\mathbb{E}_{q}\left[(1-q(r))^{K}\right]\quad\text{s.t. }\sum_{r\in\mathcal{R}}q(r)=1,\;q(r)\in[0,1)\;\forall r\in\mathcal{R}.$ (15)
When $|\mathcal{R}|\geq K+1$, the maximum is $M=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$, and it is attained when $q$ is uniform.
When $|\mathcal{R}|<K+1$, the maximum is upper bounded by $M\leq\frac{|\mathcal{R}|}{K+1}\left(\frac{K}{K+1}\right)^{K}$, and there exists a feasible distribution $q^{\prime}$ such that $\mathbb{E}_{q^{\prime}}\left[(1-q^{\prime}(r))^{K}\right]\geq\frac{|\mathcal{R}|-1}{K+1}\left(\frac{K}{K+1}\right)^{K}$.
Proof.
Let =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=|\mathcal{R}|$. Let $q_{i}:=q(r_{i})$ denote the probability mass at each support point. Define the function $f(q_{\ell}):=q_{\ell}\cdot(1-q_{\ell})^{K}$.
Case I $m<K+1$: The first derivative of $f$ is
$f^{\prime}(q_{\ell})=(1-q_{\ell})^{K-1}\big{(}1-(K+1)q_{\ell}\big{)},$
which is positive on the interval $\left[0,\frac{1}{K+1}\right)$ and negative on $\left(\frac{1}{K+1},1\right)$. Therefore, $f(q_{\ell})$ attains its maximum over $[0,1)$ at
$q^{*}_{\ell}=\frac{1}{K+1},$
with the corresponding maximum value
$f(q^{*}_{\ell})\leq\frac{1}{K+1}\left(\frac{K}{K+1}\right)^{K}.$
Since there are $m$ support points, this yields the upper bound $M\leq\sum_{\ell=1}^{m}f(q^{*}_{\ell})\leq\frac{m}{K+1}\left(\frac{K}{K+1}\right)^{K}$. Also, since $\frac{1}{K+1}<\frac{1}{m}$, the solution $q_{1}=\dots=q_{m-1}=\frac{1}{K+1}$ and $q_{m}=\frac{K-m+2}{K+1}$ is feasible. Thus,
$M\geq\sum_{\ell=1}^{m-1}f\left(\frac{1}{K+1}\right)+f\left(\frac{K-m+2}{K+1}\right)\geq\frac{m-1}{K+1}\left(\frac{K}{K+1}\right)^{K},$
where the right inequality uses that $f(q_{\ell})\geq 0$ for $q_{\ell}\in[0,1]$.
Case II $m\geq K+1$: Consider the relaxed maximization problem of Equation (15):
$\max_{q_{1},\dots,q_{m}\in[0,1)}\sum_{\ell=1}^{m}f(q_{\ell})\quad\text{subjectto }\sum_{\ell=1}^{m}q_{\ell}\leq 1.$ (16)
Let $\lambda\geq 0$ be the Lagrange multiplier associated with the constraint. Define the Lagrangian:
$\mathcal{L}(q_{1},\dots,q_{m},\lambda)=\sum_{\ell=1}^{m}q_{\ell}(1-q_{\ell})^{K}-\lambda\left(\sum_{\ell=1}^{m}q_{\ell}-1\right).$
For each $\ell=1,\dots,m$, compute the partial derivative of $\mathcal{L}$ with respect to $q_{\ell}$:
$\frac{\partial}{\partial q_{\ell}}\left[q_{\ell}(1-q_{\ell})^{K}\right]=(1-q_{\ell})^{K}-Kq_{\ell}(1-q_{\ell})^{K-1}.$
Setting this derivative equal to zero yields the stationary condition:
$(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda,\quad\text{with }\lambda\geq 0.$
To find critical points of Equation (16), we solve the system:
$(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda\geq 0\quad\forall\ell\leq m,\quad\text{and}\quad\sum_{\ell=1}^{m}q_{\ell}\leq 1.$
Define the function $g(q_{\ell}):=(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})$. For $g(q_{\ell})\geq 0$, it must hold that $1-(K+1)q_{\ell}\geq 0$, i.e., $q_{\ell}\leq\frac{1}{K+1}$. Therefore, any feasible solution to this system must satisfy $q_{\ell}\in\left[0,\frac{1}{K+1}\right]$ for all $\ell\leq m$.
Then, over the interval $q_{\ell}\in\left[0,\frac{1}{K+1}\right]$, both factors $(1-q_{\ell})^{K-1}$ and $1-(K+1)q_{\ell}$ are positive and decreasing. Hence, $g(q_{\ell})$ is positive and strictly decreasing, so the equation $g(q_{\ell})=\lambda\geq 0$ has at most one solution. Therefore, all $q_{\ell}$’s must be equal at a critical point. Let $q_{\ell}=c$ for all $\ell\leq m$. Due to the assumption of $m\geq K+1$, we have $\frac{1}{m}\leq\frac{1}{K+1}$, so any choice of $c\in\left[0,\frac{1}{m}\right]$ satisfies the constraint $\sum_{\ell=1}^{m}q_{\ell}\leq 1$ and is feasible for the system.
Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of Equation (16). The corresponding objective value is:
$\sum_{\ell=1}^{m}f(c)=m\cdot c\cdot(1-c)^{K},$
which is increasing in $c$ over the interval $\left[0,\frac{1}{m}\right]$. Hence, the maximum is attained at $c=\frac{1}{m}$, and the optimal value is:
$\sum_{\ell=1}^{m}f\left(\frac{1}{m}\right)=\left(1-\frac{1}{m}\right)^{K}=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K},$
achieved when $q_{\ell}=\frac{1}{m}$ for all $\ell$.
To confirm that this critical point is indeed a maximum, observe that the Hessian of the objective function $f(q_{\ell})$ is diagonal (since the function is separable), and the diagonal entries are:
$\frac{\partial^{2}}{\partial q_{\ell}^{2}}\left[q_{\ell}(1-q_{\ell})^{K}\right]=-(1-q_{\ell})^{K-2}\left(2K-K(K+1)q_{\ell}\right),$
which is negative for $q_{\ell}\leq\frac{1}{K+1}$, because then $02-(K+1)q_{\ell}>$ 0. Hence, the Hessian is negative definite, and the critical point is a local (and thus global) maximum.
Finally, note that Equation (16) is a relaxation of Equation (15). While the optimum of the original problem is upper bounded by that of the relaxed problem, the optimal solution to the relaxed problem also lies within the feasible region of the original problem. Therefore, the maximum of Equation (15) is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$. ∎
Statement of Theorem 7
Under Assumption 2, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{full}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n\in\mathbb{N}$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 5, we have:
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)},\quad\text{where }K\geq\frac{1}{2}\left(\frac{M^{2}N(\mathcal{A}_{full},\epsilon)}{\epsilon^{2}e}-1\right).$
As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
Proof.
By Lemma 13, we can shift the action space to be centered at the origin. So, without loss of generality, we assume that $\mathcal{A}_{full}\subset M\cdot B_{2}^{n}$, the scaled unit Euclidean ball in $\mathbb{R}^{n}$. Then,
$\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}^{2}\right]\leq M^{2}\cdot\mathbb{E}\|\theta\|_{2}^{2}=M^{2}n,$ (17)
where the inequality follows from $\mu_{a}=\langle\theta,a\rangle\leq M\|\theta\|_{2}$ for all $a\in MB_{2}^{n}$, and the equality uses that $\theta\sim\mathcal{N}(0,I)$, so each coordinate has variance 1 and $\|\theta\|_{2}^{2}=\sum_{i=1}^{n}\theta_{i}^{2}$ has expectation $n$.
Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{full}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{full},\epsilon)$. Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
$r_{\ell}=\left\{a\in\mathcal{A}_{full}:\pi(a)=a_{\ell}\right\}.$
Then,
$\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}\right]\leq\mathbb{E}_{\theta}\left[\max_{a\in B(a_{\ell},\epsilon)}\mu_{a}\right]=\mathbb{E}_{\theta}\left[\max_{a\in\epsilon B_{2}^{n}}\mu_{a}\right]\leq\epsilon\sqrt{n},$ (18)
where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 13 to shift; and the final bound uses Claim 3 from Supplementary K. Then,
$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2},$
$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}\cdot M\sqrt{n}$
$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\epsilon\sqrt{n},$
where the first inequality follows from the algorithm-dependent upper bound in Theorem 6. The second inequality follows from Equations (17) and (18), and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma 15). The last inequality comes from
$\log\left(\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}M\right)=$ $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)+K\log\left(\tfrac{2K}{2K+1}\right)+\log M$
$\leq$ $\tfrac{1}{2}\log m-\tfrac{1}{2}\log(2K+1)-\tfrac{1}{2}+\log M$
$\leq$ $\tfrac{1}{2}\log m-\tfrac{1}{2}\log\left(\frac{mM^{2}}{\epsilon^{2}e}\right)-\tfrac{1}{2}+\log M\leq\log\epsilon,$
where the first inequality uses $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)=\tfrac{1}{2}(\log m-\log(2K+1))$. Also, $\log\left(\tfrac{2K}{2K+1}\right)=\log\left(1-\tfrac{1}{2K+1}\right)\leq-\tfrac{1}{2K+1}$ due to the inequality $\log(1-X)\leq-X$ for $X<1$, and for a large $K$, we approximate $-\tfrac{K}{2K+1}\approx-\frac{1}{2}$. The last inequality uses the assumption $K\geq\frac{1}{2}\left(\frac{mM^{2}}{\epsilon^{2}e}-1\right)$.
Finally, since $\mathcal{A}_{full}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{full},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of [49]. Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$. ∎
Appendix H Proof of Theorem 8
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 4. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. Then, for the same constant $C$ in Theorem 5 and another constant $c>0$, for any reference subset $\mathcal{A}^{\prime}$:
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).\end{split}$
Proof.
Choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative, such that the subset $\mathcal{A}^{\prime}=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1. Fix $\ell$. As in Lemma 10, define a Gaussian process $(Z_{a})_{a\in r_{\ell}}$ by setting $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and define random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. The key idea is that $\operatorname*{Regret}\geq Y_{\ell}$ whenever $r_{\ell}\cap\mathcal{A}=\emptyset$ and $a^{*}(\theta)\in r_{\ell}$, such that
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]=&\mathbb{E}\left[\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ \geq&\mathbb{E}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ =&\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (19)
where the inequality follows from Assumption 4. The last equality follows from $Y_{\ell}$ is independent from $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\right]=&\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]\\ &+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]\\ \geq&\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right],\\ \geq&\sum_{\ell\leq m}q(r_{\ell})\cdot(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}^{2}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\right)^{1/2}\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\end{split}$
where the equality uses the tower rule. The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 14. The second inequality follows from Equation (19). The third inequality uses the Pólya-Szegő inequality [17], corresponding to a constant $c>0$:
$c:=\frac{1}{2}\left(\sqrt{\frac{M_{1}M_{2}}{m_{1}m_{2}}}+\sqrt{\frac{m_{1}m_{2}}{M_{1}M_{2}}}\right),$
where $m_{1},m_{2},M_{1},M_{2}$ are constants such that
[…]
Note that $m_{2}$ is positive due to the assumption that each cluster contains more than one action. The fourth inequality follows from Jensen’s inequality that for a random variable $X$, $\mathbb{E}[X^{2}]\geq\mathbb{E}[X]^{2}$, and Lemma 10. Since the reference subset $\mathcal{A}^{\prime}$ is arbitrary, the lower bound $c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]$ holds for any reference subset. The last inequality uses Theorem 5. ∎
Appendix I Example of orthogonal basis
In general, even when the distribution of bandit instances $\theta$ is well-behaved, e.g., the standard Gaussian distribution in $\mathbb{R}^{n}$, the importance measure $q$ is usually non-trivial. In this section, we consider an example of the action space being orthogonal basis.
Let the set $\mathcal{A}_{full}=\{e_{i}:i=1,\dots,n\}$ consists of the $n$ unit vectors aligned with the coordinate axes in $\mathbb{R}^{n}$. Hence, $\operatorname{diam}(\mathcal{A}_{full})=\sqrt{2}$. In this case, $\max_{a\in\mathcal{A}_{full}}\langle a,\theta\rangle$ is equivalent to the maximum among the $n$ entries of $\theta$, where by definition, each entry follows i.i.d. from the standard normal distribution $\mathcal{N}(0,1)$, such that each entry has the same probability of having the maximum value. Let each cluster contain only one unit vector, thus the importance measure of clusters $q$ is a uniform distribution.
Expected maximum in $\mathcal{A}_{full}$: Let $X_{i}$, $i=1,\dots,n$ be i.i.d. samples from $\mathcal{N}(0,1)$, then
$\mathbb{E}\max_{a\in\mathcal{A}_{full}}\mu_{a}=\mathbb{E}_{\theta}\max_{i=1,\dots,n}\theta_{i}=\mathbb{E}\max_{i=1,\dots,n}X_{i},$
where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary L, we have
$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}_{full}}\mu_{a}\leq\sqrt{2\log n},\quad\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}}\mu_{a}\leq\sqrt{2\log|\mathcal{A}|}.$
Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in Equation (1),
$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}\leq\mathbb{E}_{\theta}[\textnormal{Regret}]\leq\sqrt{2\log n}-\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}.$
Upper bound of expected regret of Algorithm 1: Since the importance measure of clusters $q$ is a uniform distribution, we have
$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\left(1-\frac{1}{n}\right)^{K}\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right]\leq\left(1-\frac{1}{n}\right)^{K}\sqrt{2\log n},$
where the left inequality uses Corollary 16 and that if each cluster contains one action, $\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]=0$ for each cluster and $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)=0$.
Corollary 16.
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where the $q$ measure is uniform over all clusters. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 5, it holds that
$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right].$
Under Assumption 4,
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).$
Proof.
Consider the variables $Y_{\ell}$ defined in the proof of Theorem 6,
$\mathbb{E}[\operatorname*{Regret}]\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\sum_{r\in\mathcal{R}}q(r)\cdot\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r\right]$
$=$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right],$
where the first inequality comes from the proof of Theorem 6, the equality uses $q(r)=1/|\mathcal{R}|$ for all $r\in\mathcal{R}$. Then, using Equation (11) in the proof of Theorem 5, we derive the upper bound.
Consider the variables $Y_{\ell}$ defined in the proof of Theorem 8,
$\mathbb{E}[\operatorname*{Regret}]\geq$ $\sum_{\ell\leq m}q(r_{\ell})\cdot\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]$
$=$ $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\geq\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right],$
where the first inequality comes from the proof of Theorem 8, the equality uses $q(r)=1/|\mathcal{R}|$ for all $r\in\mathcal{R}$. The last inequality uses the definition of $q$ measure and $Y_{\ell}\geq\min_{\ell\leq m}Y_{\ell}$. Then, using Equation (12) in the proof of Theorem 5, we derive the lower bound. ∎
Appendix J Comparisons with UCB & TS
To understand why Algorithm 1 outperforms UCB and TS, we still consider the fixed action space of 15 grid points over the interval $[0,2]$ and indexed by $0,1,\ldots,14$. Outcome functions are generated using an RBF kernel with a fixed length-scale $l=1$. The middle upper subplot of Figure 4 displays ten example outcome functions sampled from this setup. In each trial, Algorithms 1, 2 (UCB), and 3 (TS) are run to select $K=5$ distinct representative actions. UCB and TS are each executed for 3000 rounds (excluding initialization), with each round corresponding to a bandit instance. The number of rounds is chosen to approximately match the number of possible super arms, given by $\binom{15}{5}=3003$.
The upper part of Figure 3 visualizes the representative actions selected by each algorithm in three individual trials. Each row corresponds to one trial. The lower part shows the frequency with which each of the 15 actions is selected across 50 independent trials. The frequencies for actions indexed 1 through 13 are relatively uniform across algorithms, but a notable distinction is that Algorithm 1 (green) exhibits a strong preference toward the boundary actions—those indexed 0 and 14.
To understand this preference of Algorithm 1, recall that its output frequencies directly reflect the underlying measure $q$, since the algorithm samples actions according to $q$ by design. In Figure 4, we vary the RBF kernel length-scale across values 0.1, 1, and 4. For each setting, we sample ten example outcome functions (shown in the upper part) and run Algorithm 1 for 50 trials, recording the frequency of each selected action as an approximation of the $q$-measure (shown in the lower part).
In Figure 4, when the length-scale is 4, the outcome functions are very smooth—almost linear—so the maximum is typically attained at the boundary actions (indices 0 and 14), causing the $q$-measure to concentrate heavily on them. In contrast, with a length-scale of 0.1, the functions are very rough, with outcomes nearly independent across actions, resulting in an approximately uniform $q$-measure over the action space. Further, since the outcome functions are generated using a stationary RBF kernel, the corresponding feature representations $\Phi(a)$, for $a\in\mathcal{A}_{full}$, are uniformly distributed. When action outcomes are nearly independent (e.g., with a very small length-scale), the feature vectors $\Phi(a)$ form an orthogonal basis, with each pair $(\Phi(a),\Phi(a^{\prime}))$ equidistant. As discussed in Supplementary I, this leads to a uniform $q$-measure when each cluster contains a single action. Indeed, the lower-left part of Figure 4 confirms this: for length-scale $l=0.1$, the action selection frequencies are nearly uniform. As the length-scale increases, action outcomes become more correlated, and the pairwise distances between $\Phi(a)$ vectors shrink. The boundary actions (indices 0 and 14), remain the farthest apart and thus are more likely to be optimal in a given bandit instance.
This explains why Algorithm 1 performs better: it closely aligned with the $q$-measure, whereas UCB and TS select actions more uniformly. Further, in any given round $\theta$ of TS/UCB, the reward $\max_{a\in\mathcal{A}}\mu_{a}$ for optimal and suboptimal super arms can be indistinguishable if both include the optimal base action $a^{*}(\theta)$. This ambiguity slows down the convergence of the learning algorithms because they cannot effectively differentiate between good and bad super arms—a problem exacerbated by the combinatorial number of possible super arms.
Figure 3: Comparison of representative actions selected by three algorithms in a synthetic setting. Outcome functions are generated using an RBF kernel (length-scale $l=1$) over 15 grid points in $[0,2]$. In each trial, $K=5$ distinct actions are selected per algorithm (Algorithm 1 in green, TS in yellow, and UCB in red); UCB and TS are run for 3000 rounds. Upper: Representative actions selected in three individual trials; each row shows one trial. Lower: Action selection frequencies of Algorithm 1, TS, and UCB over 50 trials. Algorithm 1 consistently selects boundary actions (indices 0 and 14), while UCB and TS select more uniformly.
Figure 4: Effect of kernel length-scale on importance measure $q$. Upper: Example outcome functions generated using RBF kernels with varying length-scales $l\in\{0.1,1,4\}$, over 15 grid points in $[0,2]$, with actions indexed from 0 to 14. Lower: Action selection frequencies of Algorithm 1 over 50 trials for each length-scale setting. These frequencies approximate the importance measure $q$. As the length-scale increases, the outcome functions become smoother and more linear, causing $q$-measure to concentrate at the boundary actions (indices 0 and 14).
Algorithm 2 Upper Confidence Bound with Gaussian Rewards
1:Input: Base actions $\mathcal{A}_{full}$, subset cardinality $K$, kernel matrix $\mathbf{K}$, time horizon $T=3000$
2:Initialize: Super arms $\{\mathcal{A}_{i}\}_{i\leq M}$ with $M=\binom{|\mathcal{A}_{full}|}{K}$, reward variance $\sigma^{2}=1$
3:Initialize: $N_{i}\leftarrow 0$, $\hat{\mu}_{i}\leftarrow 0$ for all $i=1,\dots,K$
4:for $t=1$ to $T+M$ do $\triangleright$ Each round corresponds to a bandit
5:     Sample an outcome function $f\sim\mathcal{N}(0,\mathbf{K})$
6:     if $t\leq M$ then
7:         Select arm $i_{t}=t$
8:     else
9:         Select arm $i_{t}=\arg\max_{i}\left(\hat{\mu}_{i}+\sqrt{\frac{2\log t}{N_{i}+\epsilon}}\right)$
10:     end if
11:     Pull arm $i_{t}$, observe reward $r_{t}=\max_{a\in\mathcal{A}_{i_{t}}}f(a)$
12:     Update:
$N_{i_{t}}\leftarrow N_{i_{t}}+1$
$\hat{\mu}_{i_{t}}\leftarrow\hat{\mu}_{i_{t}}+\frac{r_{t}-\hat{\mu}_{i_{t}}}{N_{i_{t}}}$
13:end for
Algorithm 3 Thompson Sampling with Gaussian Priors
1:Input: Base actions $\mathcal{A}_{full}$, subset cardinality $K$, kernel matrix $\mathbf{K}$, time horizon $T=3000$
2:Initialize: Super arms $\{\mathcal{A}_{i}\}_{i\leq M}$ with $M=\binom{|\mathcal{A}_{full}|}{K}$, reward variance $\sigma^{2}=1$
3:Initialize: prior mean $\mu_{i}\leftarrow 0$, prior precision $\tau_{i}\leftarrow\epsilon$ for all $i=1,\dots,K$
4:for $t=1$ to $T$ do $\triangleright$ Each round corresponds to a bandit
5:     Sample an outcome function $f\sim\mathcal{N}(0,\mathbf{K})$
6:     for $i=1$ to $M$ do
7:         Sample $\tilde{\mu}_{i}\sim\mathcal{N}(\mu_{i},\sigma^{2}/\tau_{i})$
8:     end for
9:     Select arm $i_{t}=\arg\max_{i}\tilde{\mu}_{i}$
10:     Pull arm $i_{t}$, observe reward $r_{t}=\max_{a\in\mathcal{A}_{i_{t}}}f(a)$
11:     Update:
$\tau_{i_{t}}\leftarrow\tau_{i_{t}}+1$
$\mu_{i_{t}}\leftarrow\mu_{i_{t}}+\frac{r_{t}-\mu_{i_{t}}}{\tau_{i_{t}}}$
12:end for
Appendix K Properties of Gaussian width
Given a set $\mathcal{S}\in\mathbb{R}^{n}$, the term $\mathbb{E}[\max_{a\in\mathcal{S}}\mu_{a}]$ where $\theta\sim\mathcal{N}(0,I)$ is called Gaussian (mean) width.
Claim 1: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
$\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,-\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle-a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
where the first equality uses $-\theta$ and $\theta$ are identically distributed. The third equality uses for any $a\in\mathcal{S}$, it holds that $-a\in-\mathcal{S}$.
Claim 2: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]\leq\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$.
Let $a^{*}(-\mathcal{S},\theta)$ denote the optimal action in $\mathcal{S}$ for bandit instance $\theta$.
$2\cdot\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]+\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$
$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta),\;\theta\rangle\right]+\mathbb{E}\left[\langle a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
$\leq$ $\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right],$
where the first equality uses Claim 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$. The third equality uses linearity of expectation. The inequality uses that the vector $a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta)$ belongs to the set of vectors $\{a-a^{\prime}:a,a^{\prime}\in\mathcal{S}\}$. In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need this inequality to prove the next claim.
Claim 3: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n}$.
$\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]$
$\leq$ $\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$
$\leq$ $\frac{1}{2}\mathbb{E}\max_{a,a^{\prime}\in\mathcal{S}}\|\theta\|_{2}\|a-a^{\prime}\|_{2}$
$\leq$ $\frac{1}{2}\mathbb{E}\operatorname{diam}(\mathcal{S})\|\theta\|_{2}\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n},$
where the first equality uses the definition of $\mu_{a}$. The first inequality uses Claim 2. The second inequality uses Cauchy-Schwarz inequality. The third inequality uses the definition of $\operatorname{diam}(\cdot)$. The last inequality uses $\mathbb{E}\|\theta\|_{2}\leq\sqrt{n}$.
Appendix L Bounds of expected maximum of Gaussian
Let $X_{1},\dots,X_{N}$ be $N$ random Gaussian variables (no necessarily independent) with zero mean and variance of marginals smaller than $\sigma^{2}$, then
$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]\leq\sigma\sqrt{2\log N}.$
Proof.
for any $\delta>0$,
$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]=$ $\frac{1}{\delta}\mathbb{E}\left[\log\exp(\delta\max_{i=1,\dots,N}X_{i})\right]\leq\frac{1}{\delta}\log\mathbb{E}\left[\exp(\delta\max_{i=1,\dots,N}X_{i})\right]$
$=$ $\frac{1}{\delta}\log\mathbb{E}\left[\max_{i=1,\dots,N}\exp(\delta X_{i})\right]\leq\frac{1}{\delta}\log\sum_{i=1}^{N}\mathbb{E}\left[\exp(\delta X_{i})\right]$
$\leq$ $\frac{1}{\delta}\log\sum_{i=1}^{N}\exp(\sigma^{2}\delta^{2}/2)=\frac{\log N}{\delta}+\frac{\sigma^{2}\delta}{2},$
where the first inequality uses Jensen’s inequality. Taking $\delta:=\sqrt{2(\log N)/\sigma^{2}}$ yields the results. ∎
Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [28]:
$\mathbb{E}_{\theta}\left[\max_{i=1,\dots,N}X_{i}\right]\geq\frac{\sigma\sqrt{\log N}}{\sqrt{\pi\log 2}}.$

### v2

@@
-Representative Action Selection for Large Action-Space Meta-Bandits
+Representative Action Selection for Large Action Space Meta-Bandits
 Abstract
@@
 Conversely, if flu season is approaching, stocking several flu medications is a wise choice.
-Different from prior approaches in multi-armed bandits (MAB) (dani2008stochastic, ) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits.
+Different from prior approaches in multi-armed bandits (MAB) (Dani et al., 2008) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits.
 This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
-Consider the following MAB setting: In a bandit, if a decision-maker plays an action $a\in\mathcal{A}_{full}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$.
-We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle,$ where this bandit instance $\theta\in\mathbb{R}^{n}$ follows a multivariate Gaussian distribution.
-Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ forms a Gaussian process (vershynin2018high, ).
-Our underlying setting aligns with the one considered in contextual bandit (dani2008stochastic, ; bietti2021contextual, ); however, the key difference lies in the objective.
-Now consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{full}$.
-For a given bandit instance $\theta$, if the best action over the full space lies within $\mathcal{A}$, the decision-maker benefits from a reduced set of suboptimal actions to explore.
+Consider the following MAB setting: In a bandit, if a decision-maker plays an action with a fixed but unknown feature vector $a\in\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$.
+We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle$ where this bandit instance $\theta\in\mathbb{R}^{n}$ is drawn from an unknown multivariate Gaussian distribution.
+Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ forms a Gaussian process (GP) (Vershynin, 2018, Chapter 7), and we also provide results under general sub-Gaussian assumption.
+Our underlying setting aligns with the one considered in contextual bandit (Dani et al., 2008); however, the key difference lies in the objective.
+Now consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$.
+For a given bandit instance $\theta$, if the best action over the full space lies within $\mathcal{A}$, the decision-maker benefits from reduced suboptimal actions to explore.
 Conversely, if the best action lies outside $\mathcal{A}$, regret arises from being unable to select this best action.
 Thus, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
-$\textnormal{Regret}:=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$ (1)
+$\boxed{\textnormal{Regret}:=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}},}$ (1)
 which depends on the sampled bandit instance and is therefore a random variable.
@@
 This objective is motivated by practical considerations.
-MAB regret in finite-horizon settings (auer2002using, ; auer2002nonstochastic, ) typically scales with the number of available actions.
-However, if a subset $\mathcal{A}$ is carefully chosen, the resulting MAB regret can be lower than when using the full action space, due to a reduced set of suboptimal actions.
-This is especially beneficial when there are a very large number of actions, as even the initialization phase can be computationally expensive.
-This improvement arises from a natural connection between the regret in Equation (1) and MAB regret, as discussed in Supplementary A.
-The remainder of this section is literature review. The subsequent sections present our main contributions and are organized as follows:
-  • Section 2: Subset Selection Framework and Algorithm.
-We introduce a framework that models expected outcome using a Gaussian process to capture correlations across similar actions and bandit instances.
-Based on this structure, we construct geometric $\epsilon$-nets and refine them using an importance-based selection mechanism.
-This mechanism is formalized as a variant of the $\epsilon$-net algorithm (haussler1986epsilon, ), tailored to the bandit setting.
-  • Section 3: Regret Analysis. We provide theoretical guarantees, including upper and lower bounds on the expected regret that reflect the geometric and importance structure of the action space.
-  • Section 4: Examples and Empirical Validation.
-We extend our analysis to settings where outcome (payoff) functions are sampled from a Gaussian process with a kernel, and empirically compare the performance to that of Thompson Sampling and Upper Confidence Bound.
-The code used for the experiments is available at https://github.com/Quan-Zhou/MAB.
+The classic Bayesian regret in the bandit literature (Agrawal and Goyal, 2012) typically scales with the number of available actions.
+However, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower due to the reduced number of suboptimal actions.
+This is especially beneficial when the action space is large, as even the initialization phase can be computationally expensive.
+To see this, we can decompose the bandit regret as follows:
+[…]
+$=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}}\mu_{a}(\theta)\!-\!\mu_{A_{t}}+\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)\!-\!\max_{a\in\mathcal{A}}\mu_{a}(\theta)\right]$
+$\leq C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}].$
+In the first equality, $A_{t}\in\mathcal{A}$ denotes the action chosen by a policy, e.g., Thompson Sampling (Agrawal and Goyal, 2012), in round $t$, and $N$ is the number of rounds.
+The expectation is taken over the randomness in the distribution of bandit instances and in actions selected by the policy.
+The inequality follows from the well-known regret bounds for Thompson Sampling (Lattimore and Szepesvári, 2020).
+$|\mathcal{A}|$ denotes the cardinality of the action set and $C>0$ is a constant.
+Note that if the policy has access to the full action space, the Bayesian bandit regret is instead bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$.
+Our main contributions are organized as follows:
+  • Meta-Bandits Framework.
+We propose a meta-bandits framework that specifically tackles combinatorial action selection by leveraging correlations across similar actions and bandit instances.
+To the best of our knowledge, this is the first such framework.
+  • $\epsilon$-Net Algorithm.
+We introduce a simple algorithm within this framework.
+It starts with the intuitive idea of placing a grid over the action space, then refines it using an importance-based selection mechanism.
+  • Regret Analysis.
+We provide theoretical guarantees for both the grid and the algorithm’s output, including upper and lower bounds on expected regret, along with results under general sub-Gaussian processes.
+We also discuss the cost of not using a grid, which depends on the importance-structure of the action space.
+  • Generalization and Empirical Validation.
+We extend the analysis to settings where outcome functions are sampled from a reproducing kernel Hilbert space (RKHS), and empirically compare our algorithm to Thompson Sampling (TS) and Upper Confidence Bound (UCB).
 1.1 Related Works
-Multi-Armed Bandits The classic MAB problem (lai1985asymptotically, ; auer2002finite, ) is defined by a set of arms, each deliver outcomes that are independently drawn from a fixed and unknown distribution.
-The decision-maker sequentially selects an arm, observes its outcome, and aims to maximize cumulative outcomes over time.
-The policy is to choose the next arm based on the sequence of past plays and obtained outcomes.
-Popular policies include the Upper Confidence Bound (auer2002finite, ), Thompson Sampling (agrawal2012analysis, ), and EXP3 (auer2002nonstochastic, ).
-Contextual Bandits Unlike classic MAB, which assumes action outcomes are independent, contextual bandits (auer2002using, ) incorporate observed features (contexts) at each round.
-The decision-maker selects actions based on context, aiming to learn a policy that maps contexts to high-reward actions.
-Linear models such as LinUCB (chu2011contextual, ) and linear Thompson sampling (agrawal2013thompson, ) assume linear reward functions, enabling efficient learning.
-Early work includes LinRel (auer2002using, ), while later efforts extended to nonlinear models (filippi2010parametric, ; jun2017scalable, ).
-Nonlinear payoff structures have also been explored using Gaussian processes (srinivas2009gaussian, ) and RKHS methods (valko2013finite, ; chowdhury2017kernelized, ).
-Recent studies examine richer structures: budget-constrained linear contextual bandits (agrawal2016linear, ), reward feedback across contexts (balseiro2019contextual, ), and adversarially corrupted outcomes (he2022nearly, ).
-Empirical evaluations of common algorithms appear in bietti2021contextual .
-Online Combinatorial Optimization Several works have extended stochastic linear optimization (dani2008stochastic, ; rusmevichientong2010linearly, ) to combinatorial settings.
-This line of work applies the Follow-the-Perturbed-Leader (FPL) algorithm (hannan1957approximation, ; kalai2005efficient, ), which introduces random perturbations to the estimated cumulative loss of each action, resulting in simple and efficient algorithms for various combinatorial structures.
-This idea has also been extended to online convex optimization.
-For instance, (hazan2007logarithmic, ) develop an FPL-style algorithm that achieves logarithmic regret for strictly convex payoff functions, particularly in settings where the decision point lies in the simplex.
-Follow-the-Regularized-Leader (FTRL) (shalev2006convex, ; shalev2012online, ) generalizes Follow-the-Leader by adding a regularization term to stabilize updates.
-(mcmahan2011follow, ) show that many online learning algorithms, including mirror descent, can be interpreted as instances of FTRL.
-In the context of online combinatorial optimization, (audibert2011minimax, ; audibert2014regret, ) study minimax regret bounds and propose FTRL-based algorithms alongside other strategies.
-(ito2021hybrid, ) further develop a hybrid algorithm based on FTRL.
-Combinatorial Bandits Combinatorial bandits (cesa2012combinatorial, ) are a subclass of online combinatorial optimization where the decision-maker selects a subset of base arms from $\mathcal{A}_{full}$ in each round, forming a super arm.
-Early work on online shortest path problems with adversarial edge weights (gyorgy2007line, ; uchiya2010algorithms, ; liu2012adaptive, ) adapts the EXP3 framework (auer2002nonstochastic, ).
-Subsequent studies establish regret bounds for Thompson Sampling in structured settings (gopalan2014thompson, ) and propose general models with non-linear rewards and semi-bandit feedback (chen2016combinatorial, ; kveton2015tight, ).
-The matroid bandit model (kveton2014matroid, ) captures independence constraints via matroid optimization (whitney1992abstract, ).
-Later work reframes the shortest path problem as an adversarial MDP (rosenberg2019online, ) and introduces improved UCB variants with adaptive confidence estimation (zhou2024towards, ).
-Epsilon net We identified two definitions of $\epsilon$-nets, with further discussion in Section 2.1.
-The first net of a given set, referring to (vershynin2018high, ), ensures that closed balls of radius $\epsilon$ centered at net points cover the given set.
-This relates to the covering number, that is the smallest possible cardinality of an $\epsilon$-net.
-This concept extends to function classes as a complexity measure, known as metric entropy, with further details in wainwright2019high , and has been used in russo2013eluder on the class of outcome functions.
-The second definition, as given in (matousek2013lectures, ), relates to importance sampling, requiring that every significant subset of a set intersects with the net.
-The classic $\epsilon$-net algorithm, introduced by haussler1986epsilon , operates via random sampling, yielding a net of size $\mathcal{O}(\frac{1}{\epsilon}\log\frac{1}{\epsilon})$.
-As noted in matousek2013lectures , the importance measure extends to general probability measures, inspiring our approach.
-Further works aim to reduce net size (mustafa2019computing, ; pach2011tight, ; rabani2009explicit, ) and address online settings (bhore2024online, ).
+Multi-Armed Bandits (Lai and Robbins, 1985; Auer et al., 2002a) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution.
+The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time.
+Popular methods include the UCB (Auer et al., 2002a), TS (Agrawal and Goyal, 2012), and EXP3 (Auer et al., 2002b) for adversarial settings.
+Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few independent samples as possible (Jamieson and Nowak, 2014; Kaufmann et al., 2016).
+Popular methods in fixed confidence setting include Action Elimination Even-Dar et al.
+(2006); Karnin et al.
+(2013), UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum.
+In fixed budget setting, there is Successive halving Karnin et al.
+(2013), successive reject Audibert and Bubeck (2010).
+Our algorithm includes a step for identifying the optimal action.
+While this line of work assumes (non-centered) sub-Gaussian outcome distributions within a bandit instance, we model each instance with a linear structure, allowing for further simplification.
+Stochastic Linear Optimization assumes that the expected outcome of each action depends through the inner product between a context $\theta$ and an action $a\in\mathcal{A}_{\operatorname*{full}}$ Auer (2002); Dani et al.
+(2008); Rusmevichientong and Tsitsiklis (2010), and our model of each bandit instance align with this setting.
+This is a sequential method for maximizing a linear function—i.e., $f(a):=\mu_{a}(\theta)$ with fixed $\theta$—based on noisy observations of the function values.
+Soare et al.
+(2014) study the sample complexity of optimal action identification in this setting.
+However, this line of work assumes the action feature vectors are known.
+GP Optimization can deal with the case that the feature vectors of actions are unknown.
+Srinivas et al.
+(2009) models the outcome function as a sample from a GP prior with a kernel function (Williams and Rasmussen, 2006).
+Their Bayesian algorithm, GP-UCB, uses this GP prior to first select points $x_{t}$ that enable global estimation of the function, and then plays the point with the highest posterior mean.
+The widespread adoption of this method in bandit settings (Valko et al., 2013; Li and Scarlett, 2022), as well as in continuous action spaces (Chowdhury and Gopalan, 2017), highlights the practicality of assuming that action outcomes are correlated.
+This also motivates our extension, where the outcome function of each bandit instance is modeled as a sample from a RKHS.
+Before discussing two seemingly relevant lines of work that select a subset from the action space, we first highlight a advantage of our approach.
+These methods assume a fixed $|\mathcal{A}|=K$, which is often unclear in practice and requires restarting the algorithm when changed.
+In contrast, our algorithm can adapt the subset size on the fly.
+For example, if 100 runs all return the same action, choosing this one may suffice ($|\mathcal{A}|=1$); if each run yields a different action, we may need more iterations to ensure a small regret ($|\mathcal{A}|>100$).
+[…]
+Example 1.
+Consider three actions: $a_{1}=[1,0]$, $a_{2}=[0.9,0.1]$, and $a_{3}=[-0.1,1]$, and suppose bandits are sampled uniformly from $\theta_{1}=[1,0]$ and $\theta_{2}=[0,1]$. Then,
+$\mathbb{E}\langle a_{1},\theta\rangle=\mathbb{E}\langle a_{2},\theta\rangle=0.5,\quad\mathbb{E}\langle a_{3},\theta\rangle=0.45.$
+So under the Best-222-Action perspective, $a_{1}$ and $a_{2}$ would be selected. However, this is suboptimal in our framework, since $\mu_{a_{1}}$ and $\mu_{a_{2}}$ are positively correlated:
+$\mathbb{E}\max_{a\in\{a_{1},a_{2}\}}\langle a,\theta\rangle=0.55,\quad\mathbb{E}\max_{a\in\{a_{1},a_{3}\}}\langle a,\theta\rangle=1.$
+In contrast, our algorithm, if run until it selects two distinct actions, would output $a_{1}$ and $a_{3}$, which is the true optimal.
+Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super arm $\mathcal{A}$, with $|\mathcal{A}|=K$.
+Popular methods include CUCB Chen et al.
+(2016), CTS Wang and Chen (2018).
+We argue that this line of work is not applicable to our framework: (1) It assumes that the expected outcome of a super arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition.
+In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions.
+(2) It assumes independence across base arms, whereas we explicitly model correlations.
+Ignoring these correlations misses the core challenge—an issue illustrated in Example 1.
+[…]
+Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$.
+It is an important topic in high-dimensional probability (Vershynin, 2018).
+The sharpest known bounds are due to Talagrand (2014).
 2 Subset Selection Framework
-We consider the problem of selecting a small number of representative action from a large (potentially infinite but compact) action space, $\mathcal{A}_{full}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$.
-Define the expected outcome of playing action $a$ in bandit instance $\theta$ as
-$\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{full},\theta\sim\mathcal{N}(0,I),$ (2)
-where the bandit instance $\theta$ follows a multivariate normal distribution with zero mean and identity covariance matrix $I$.
-Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ forms a canonical Gaussian process (vershynin2018high, ).
-In fact, any constant-mean Gaussian process can be represented as the canonical process; see Supplementary B for details.
-Let $\operatorname*{Conv}(\mathcal{A}_{full})$ denote the convex hull that consists of all convex combinations of finitely many actions in $\mathcal{A}_{full}$.
-The extreme points of the convex hull are those points $x\in\operatorname*{Conv}(\mathcal{A}_{full})$ for which there do not exist distinct points $a,a^{\prime}\in\operatorname*{Conv}(\mathcal{A}_{full})$ and $\lambda\in(0,1)$ such that $x=\lambda a+(1-\lambda)a^{\prime}$.
-By the extreme point theorem, i.e., the optimum of the linear function on a convex set is attained at an extreme point—it follows that the optimal action for any bandit instance is always an extreme point of the full action space.
-Therefore, if we use all the extreme points as representatives of the full action space, the regret (as defined in Equation (1)) is zero.
-See Supplementary C for details.
+We consider the problem of selecting a small number of representative actions from a large action space $\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, where $n\in\mathbb{N}$.
+(This framework applies to the case $n=+\infty$, with the additional assumption $\sum_{i\geq 1}a^{2}_{i}\!<\!\infty$.) The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, and includes a constant $c\in\mathbb{R}$:
+$\mu_{a}:=\langle a,g\rangle+c,\quad\forall,a\in\mathcal{A}_{\operatorname*{full}},;g\sim\mathcal{N}(0,\Sigma),$ (2)
+where $\Sigma$ is a positive semi-definite matrix.
+Let $\theta\in\mathbb{R}^{n}$ follows a multivariate normal distribution with zero mean and identity covariance matrix.
+The distribution of $g$ is in fact equivalent to $\Sigma^{1/2}\theta$.
+Now, let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$.
+We have $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$ and
+$\langle a,g\rangle=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
+Therefore, the setting in Equation (2) is equivalent to
+$\mu_{a}:=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}},\;\theta\sim\mathcal{N}(0,I),$
+where $\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}}$ denotes the image of $\mathcal{A}_{\operatorname*{full}}$ under the linear transformation $\Sigma^{1/2}$.
+Since the constant $c$ does not affect the regret (as defined in Equation (1)), we can, without loss of generality, focus on this canonical Gaussian process (Vershynin, 2018, Chapter 7) in the remainder:
+$\boxed{\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{\operatorname*{full}},\theta\sim\mathcal{N}(0,I).}$ (3)
+We define the extreme points as those $x\in\mathcal{A}_{\operatorname*{full}}$ for which there do not exist distinct $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ and $\lambda\in(0,1)$ such that $x=\lambda a+(1-\lambda)a^{\prime}$.
+By the extreme point theorem, if we select all extreme points—denoted $\mathcal{A}=\{a_{1},\dots,a_{K}\}$—as representatives of the full action space, the regret is zero.
+This is because any $a\in\mathcal{A}_{\operatorname*{full}}$ can be expressed as a convex combination of the extreme points: $a=\lambda_{1}a_{1}+\dots+\lambda_{K}a_{K}$, where $\lambda_{i}\geq 0$ and $\sum_{i=1}^{K}\lambda_{i}=1$.
+Thus, for any $\theta\in\mathbb{R}^{n}$:
+$\begin{split}\langle a,\theta\rangle=\sum_{i=1}^{K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i=1}^{K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.\end{split}$
+By Equation (3), it yields
+$\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$
+where the left inequality uses $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$.
+Thus, the two quantities $\max_{a\in\mathcal{A}}\mu_{a}$ and $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ are equal.
 This example highlights how a geometric approach can be used to solve the stochastic combination problem.
-However, even if one only needs the extreme points, the set of extreme points may still be large or even infinite, e.g, the extreme points of a Euclidean ball is infinite.
+However, even if one only needs the extreme points, the set of extreme points may still be large, e.g, the extreme points of a Euclidean ball is infinite.
 To address this, we will later introduce the notions of $\epsilon$-nets.
-Without loss of generality, we assume $\mathcal{A}_{full}$ consists only of the extreme points of $\operatorname*{Conv}(\mathcal{A}_{full})$, as they are the only points of interest.
+Without loss of generality, we assume $\mathcal{A}_{\operatorname*{full}}$ consists only of the extreme points of $\mathcal{A}_{\operatorname*{full}}$, as they are the only points of interest.
 2.1 Epsilon Nets
-There are at least two definitions of $\epsilon$-net: one from a geometric perspective (vershynin2018high, ) and another from a measure-theoretic perspective (matousek2013lectures, ).
+If the set of extreme points $\mathcal{A}_{\operatorname*{full}}$ is still large, a natural approach is to construct a grid over the action space, where the grid points serve as representative actions.
+This ensures that for every action in the full space, there exists a representative that is close to it.
+This idea is formally captured by the notion of a (geometric) $\epsilon$-net.
+To proceed, we clarify what we mean by an $\epsilon$-net, as there are at least two definitions: one from a geometric perspective (Vershynin, 2018, Chapter 4) and another from a measure-theoretic perspective (Matousek, 2013, Chapter 10).
 Let $\|\cdot\|_{2}$ denote the Euclidean norm.
-  • A subset $\mathcal{A}\subseteq\mathcal{A}_{full}$ is called a Geometric $\epsilon$-net if, for all $a\in\mathcal{A}_{full}$, there exists $a^{\prime}\in\mathcal{A}$ such that
+Define the diameter of a compact set $r\in\mathbb{R}^{n}$ as $\mathrm{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$.
+  • A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a Geometric $\epsilon$-net if, for all $a\in\mathcal{A}_{\operatorname*{full}}$, there exists $a^{\prime}\in\mathcal{A}$ such that
     $\|a-a^{\prime}\|_{2}<\epsilon.$
-  • Let $\mathcal{R}$ be a finite partition of the extreme points into disjoint clusters such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{full}$.
+  • Let $\mathcal{R}$ be a finite partition of the extreme points into disjoint clusters such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{\operatorname*{full}}$.
 Given a measure $q$ assigning a value to each cluster $r\in\mathcal{R}$.
-A subset $\mathcal{A}\subseteq\mathcal{A}_{full}$ is called a Measure-Theoretic $\epsilon$-net with respect to measure $q$ if, for any cluster $r\in\mathcal{R}$, we have:
+A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a Measure-Theoretic $\epsilon$-net with respect to measure $q$ if, for any cluster $r\in\mathcal{R}$, we have:
     $r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
-Intuitively, the geometric definition ensures point-wise closeness, while the measure-theoretic definition ensures representative sampling.
-In well-behaved spaces, the two coincide.
-For example, let $\mathcal{A}_{full}=[0,1]$ and $q$ be the uniform distribution over this interval.
-A grid with spacing $\epsilon$ forms a geometric $\epsilon$-net and also ensures any subinterval of measure at least $\epsilon$ contains at least a grid point.
-Both definitions are useful.
-A geometric $\epsilon$-net ensures small regret, as it guarantees that an action close to the optimal one always exists in $\mathcal{A}$.
-However, it suffers from the curse of dimensionality: for $[0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$, exponentially in $n$.
-To address this, we incorporate the measure-theoretic $\epsilon$-net to restrict the geometric construction to only the important clusters—as determined by the $q$-measure.
-The partition $\mathcal{R}$ is the central concept that bridges the two definitions of $\epsilon$-nets.
-If two actions $a,a^{\prime}$ are in the same cluster, they are close in the Euclidean sense, and the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., in expected squared difference).
-This holds because, for the canonical Gaussian process, the increments of the process correspond to the Euclidean distance:
-$\|\mu_{a}-\mu_{a^{\prime}}\|_{L^{2}}=\left(\mathbb{E}(\mu_{a}-\mu_{a^{\prime}})^{2}\right)^{1/2}=\left(\mathbb{E}(a-a^{\prime})^{T}\theta\theta^{T}(a-a^{\prime})\right)^{1/2}=\|a-a^{\prime}\|_{2},\;\forall a,a^{\prime}\in\mathcal{A}_{full},$
-where $\theta^{T}$ denotes the transpose of $\theta$.
-The middle & right equalities use Equation (2) and $\mathbb{E}\theta\theta^{T}=I$.
-Then, selecting one point from each cluster $r\in\mathcal{R}$ forms a geometric $\epsilon$-net (which will later be formally defined as a reference subset), where $\epsilon$ is the maximum diameter of all the clusters.
-This construction yields bounded expected regret, as each selected point serves as a representative for its cluster, and all other points within the same cluster have bounded deviation from the representative.
-Furthermore, clusters in $\mathcal{A}_{full}$ that are deemed unimportant for expected regret can be ignored: even if no representative points are selected from these clusters, the resulting expected regret will not increase significantly compared to that of the reference subset.
+A geometric $\epsilon$-net ensures small regret because if two actions $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ are close in the Euclidean sense, then the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., their expected squared difference is small):
+$\|\mu_{a}\!-\!\mu_{a^{\prime}}\|_{L^{2}}=\!\left(\mathbb{E}(a\!-\!a^{\prime})^{\top}\theta\theta^{\top}(a\!-\!a^{\prime})\right)^{1/2}\!=\|a\!-\!a^{\prime}\|_{2},$
+where $\theta^{\top}$ denotes the transpose of $\theta$.
+The equalities use Equation (3) and $\mathbb{E}\theta\theta^{\top}=I$.
+Therefore, by definition, a geometric $\epsilon$-net guarantees the existence of an action $a\in\mathcal{A}$ whose expected outcome $\mu_{a}$ is close to that of the optimal action for any given bandit instance.
+However, this net suffers from the curse of dimension: e.g., for $[0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$.
+The measure-theoretic $\epsilon$-net addresses this issue.
+Put simply, the measure-theoretic $\epsilon$-net restricts the grid construction to only the most important clusters $r\in\mathcal{R}$, as determined by the $q$-measure.
 2.2 Epsilon Net Algorithm
-We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by haussler1986epsilon .
+We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by Haussler and Welzl (1986).
 It selects $K$ i.i.d.
@@
 Algorithm 1 Epsilon Net Algorithm
-1:Input: Action space $\mathcal{A}_{full}$, Sample size $K$, Distribution of bandits.
+1:Input: Action space $\mathcal{A}_{\operatorname*{full}}$, Sample size $K$.
 2:Output: A subset of actions $\mathcal{A}$.
-3:$\mathcal{A}\leftarrow\emptyset$
+[…]
 4:for $1,\dots,K$ do
 5:  Sample a bandit instance $\theta$.
-6:     Find the optimal action $a^{*}(\theta):=\operatorname*{argmax}_{a\in\mathcal{A}_{full}}\langle a,\theta\rangle$.
-7:     $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$triangleright Repetitions of actions are allowed
+6:  Find optimal action $a^{*}(\theta)\!:=\!\operatorname*{argmax}_{a\in\mathcal{A}_{\operatorname*{full}}}\langle a,\theta\rangle$.
+7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$triangleright Repetition of actions is allowed
 8:end for
-We define the optimal action in a bandit instance $\theta$ as $a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{full}}\mu_{a}(\theta)$.
+We define the optimal action in a bandit instance $\theta$ as
+$a^{*}(\theta):=\operatorname*{argmax}_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta).$
 Assumption 2 (Unique optimal action).
-The optimal action $a^{*}(\theta)$ is unique with probability $1$ for all bandit instances.
-This assumption does not exclude the case where the action space is a polytope, and there may be specific bandit instances for which the optimal actions lie on the edges, resulting in an infinite number of optimal actions.
-Since the bandit instances follow a continuous distribution, the probability of encountering these specific bandits is zero.
-Then, define the measure $q$:
+The optimal action $a^{*}(\theta)$ is unique with probability 111 for all bandit instances.
+Define the Importance Measure $q$ over a partition $\mathcal{R}$:
+$\boxed{q(r):=\Pr[a^{*}(\theta)\in r]=\!\int\!\mathbb{1}\{a^{*}(\theta)\in r\}\;p(\theta)d\theta,}$ (4)
 […]
-[…]
-Assumption 2 (Compact support).
+Assumption 3.
 The support of measure $q$ is compact.
-The compactness assumption ensures that $\mathbb{E}_{\theta}[\max_{a\in\mathcal{A}_{full}}\mu_{a}]$ is finite and that the unique optimal action exists.
-Without loss of generality, let $\mathcal{A}_{full}$ be the support of measure $q$.
+The compactness assumption ensures that the term $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\right]$ is finite and guarantees the attainment of a unique optimal action.
+Without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}$ is the support of the measure $q$.
 By definition, Algorithm 1 samples $K$ i.i.d.
-extreme points from clusters in $\mathcal{R}$ according to the probability measure $q$.
-If a cluster has a higher measure under $q$, its elements are more likely to be included in the output of Algorithm 1.
-In fact, with high probability, Algorithm 1 outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{full}$ with respect to measure $q$.
-(This is a simplified version of the $\epsilon$-net theorem; see Theorem 10.2.4 of (matousek2013lectures, ).)
-Lemma 3 (Measure-theoretic $\epsilon$-net).
+extreme points from clusters in $\mathcal{R}$ according to measure $q$.
+If a cluster $r\in\mathcal{R}$ has a higher measure $q(r)$, its elements are more likely to be included in the output.
+In fact, with high probability, this algorithm outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$ with respect to measure $q$.
+(This is a simplified version of Theorem 10.2.4 of Matousek (2013).)
+Lemma 4.
 Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$.
@@
 $r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
+The partition $\mathcal{R}$ bridges the two definitions of $\epsilon$-net: choosing one point from each cluster gives an $\epsilon$-net in both senses, though with different values of $\epsilon$.
+Geometrically, $\epsilon$ is the largest cluster diameter $\epsilon:=\max_{r\in\mathcal{R}}\mathrm{diam}(r)$; measure-theoretically, $\epsilon$ is the smallest cluster measure $\epsilon:=\min_{r\in\mathcal{R}}q(r)$.
+Hence, we have the following corollary:
+Corollary 5.
+Under the same conditions as Lemma 4, let $\epsilon:=\min_{r\in\mathcal{R}}q(r)$. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, the following holds:
+$\forall a\in\mathcal{A}_{\operatorname*{full}},\ \exists a^{\prime}\in\mathcal{A}\text{ such that }\|a-a^{\prime}\|_{2}<\max_{r\in\mathcal{R}}\operatorname{diam}(r).$
 3 Regret Analysis
@@
 We then extend the analysis to obtain algorithm-dependent bounds for the output of Algorithm 1.
-For any cluster $r\in\mathcal{R}$, we define its diameter as $\operatorname{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$.
-Throughout the following, we set $\epsilon$ to be the maximum diameter over all clusters, i.e., $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
 Definition 1 (Reference subsets).
@@
 A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$, and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
-We impose the following assumption on the partition, used only for lower bounds.
-This means that when the optimal action falls within an arbitrary cluster $r$, any action in this cluster is better than any action outside of it.
-It ensures that the clusters are relatively well-separated.
+We impose the following assumption on the partition, used only for lower bounds: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it.
+This ensures clusters are well-separated.
 Assumption 6 (Effective partition).
-For any $r\in\mathcal{R}$, whenever the optimal action lies in cluster $r$, i.e., $a^{*}(\theta)\in r$, then $\mu_{a}\geq\max_{a^{\prime}\in\mathcal{A}_{full}\setminus r}\mu_{a^{\prime}},\;\forall a\in r$.
+For any $r\in\mathcal{R}$, whenever the optimal action lies in cluster $r$, i.e., $a^{*}(\theta)\in r$, then $\mu_{a}\geq\max_{a^{\prime}\in\mathcal{A}_{\operatorname*{full}}\setminus r}\mu_{a^{\prime}},\;\forall a\in r$.
 Theorem 7 (Regret bounds of reference subsets).
@@
 Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$.
-Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded using concentration property of Gaussian process.
+Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded via concentration property of Gaussian process.
 $\square$
-Theorem 5 reveals that the expected regret grows linearly with the cluster diameter but logarithmically with the number of clusters.
-If the action space has a large overall diameter but exhibits a clustered structure—that is, actions are naturally grouped into a few tight clusters—then the reference subset can achieve significantly lower regret.
-In contrast, if the action space has a scattered or uniform structure, using a moderate number of clusters would lead to large cluster diameters, which in turn increases the regret of the reference subset.
-As shown in the next section, Theorem 5 paves the way for analyzing Algorithm 1 by bounding its expected regret in terms of a reference subset, plus an additional term—since the proof relies on the possibility that its output matches a reference subset.
-Thus, the structure-regret relationship applies to Algorithm 1’s output.
-In Figure 1, we study the effect of the cluster diameters (controlled by a spread parameter) on regret.
-We structure the clustered action space: Five center points are fixed on the unit sphere in $\mathbb{R}^{3}$, and around each center, 200 points are sampled to form five clusters.
-Each point is obtained by adding Gaussian noise (mean zero, standard deviation equal to the spread parameter) to the center direction, followed by projection back onto the unit sphere.
-Bandits are sampled from a 3-dimensional standard Gaussian distribution, i.e., $\theta\sim\mathcal{N}(0,I)$.
-The left subplot shows the expected regret of Algorithm 1 with $K=10$, computed using $10^{4}$ additional bandits, as the spread varies from 0.01 to 0.5.
-The curves and error shade represent the mean $\pm$ one standard deviation of expected regret over 30 repetitions.
-The middle and right subplots display example action spaces for spread values of 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
-Figure 1: Illustration of clustered action spaces on unit sphere in $\mathbb{R}^{3}$ and the effect of cluster diameters on regret.
-Five clusters are formed by generating 5 fixed center points, with 200 points sampled around each using Gaussian noise (spread controls the variance).
-Bandits are drawn from $\mathcal{N}(0,I)$.
-The left subplot shows the mean $\pm$ standard deviation of the expected regret (over 30 trials) as the spread varies from 0.01 to 0.5, using $10^{4}$ additional bandits.
-The middle and right subplots show example action spaces (blue dots) for spread values 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
 3.1 Regret Bounds of Algorithm
+The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds.
+The key difference is that whereas the reference subset includes a representative from each cluster, the algorithm may miss some clusters.
+However, although it may select a smaller subset, the algorithm achieves regret comparable to that of the reference subset, as it tends to miss clusters that contribute minimally to the expected regret.
 The expected regret in previous results is taken over bandit instances $\theta$.
 In contrast, since the output of Algorithm 1 is random, the expected regret analyzed in this section is taken with respect to both the algorithm’s randomness (i.e., the sampled $\mathcal{A}$) and the distribution over $\theta$.
-The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds.
-The key difference is that the algorithm may omit some clusters, whereas the reference subset includes a representative from each cluster.
-However, although it may select a smaller subset, the algorithm achieves regret comparable to that of the reference subset, as it tends to omit clusters that contribute minimally to the expected regret.
-Theorem 6 (Upper bound).
+Upper Bound.
+Theorem 8.
 Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
@@
 For the same constant $C>0$ in Theorem 7,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
 Proof sketch.
@@
 When $a^{*}(\theta)\in r_{\ell}$, we consider two cases:
-  • If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, the regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
-  • If $r_{\ell}\cap\mathcal{A}=\emptyset$, the regret is still bounded by $\max_{a\in\mathcal{A}_{full}}\mu_{a}$.
-This is justified by the assumption of $0\in\mathrm{conv}(\mathcal{A})$, since $\mathbb{E}_{\theta}[\max_{a\in\mathcal{A}}\mu_{a}]$ is invariant under affine unitary transformations of $\mathcal{A}$.
-$\square$
-Comparing Theorem 6 with Theorem 5, we observe that the algorithm-dependent upper bound exceeds that of the reference subset by an additional term:
-$\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.$
-Note that the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ behaves similarly to entropy.
-To illustrate, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$.
-This expression is small when $q(r)$ is small, and decays rapidly when $q(r)$ is large, making the overall term negligible when $q$ is concentrated on few clusters.
-In general, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified.
-We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{full},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{full}$:
-$N(\mathcal{A}_{full},\epsilon)=\min\left\{m\in\mathbb{N}:\exists\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{full},\ \forall a\in\mathcal{A}_{full},\ \exists\ell,\ \|a-a_{\ell}\|_{2}\leq\epsilon\right\}.$
-Theorem 7 (Partition-independent upper bound).
-Under Assumption 2, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{full}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$.
+  • $r_{\ell}\cap\mathcal{A}\neq\emptyset$: The regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
+  • $r_{\ell}\cap\mathcal{A}=\emptyset$: The regret is bounded by $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\!\mu_{a}$. $\square$
+The term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ in Theorem 8 behaves similarly to entropy: it is maximized when $q$ is uniform, provided $|\mathcal{R}|\geq 2K+1$ (since we can choose any partition); see Lemma 18 in Supplementary.
+Intuitively, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$.
+This expression is small when $q(r)$ is small, and decays faster when $q(r)$ gets larger, making the overall term negligible if $q$ is highly concentrated.
+The connection between Theorem 7 and 8 is insightful:
+Remark 9.
+Consider placing a grid over the action space and defining a partition $\mathcal{R}$ by assigning each point in the space to its nearest grid point.
+In this way, the grid acts as a reference subset with respect to the partition $\mathcal{R}$.
+The regret upper bound in Theorem 7 applies to this grid, as it holds for any reference subset.
+Meanwhile, Theorem 8 applies to any partition, including $\mathcal{R}$.
+As a result, we obtain a regret bound for Algorithm 1 that exceeds the grid’s bound by only one additional term:
+$\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
+where the part $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]$ is simply a constant, as the action space $\mathcal{A}_{\operatorname*{full}}$ is fixed.
+If the distribution $q$ is highly concentrated, then the expectation $\mathbb{E}_{q}[(1-q(r))^{2K}]$ is small, making the extra term negligible.
+In this case, the benefit of constructing an explicit grid—often a non-trivial task when the action feature vectors are unknown or high-dimensional—is limited.
+On the other hand, if the measure $q$ is close to uniform, then $\min_{r\in\mathcal{R}}q(r)$ is large.
+In this case, with relatively high probability, Algorithm 1 outputs a geometric $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$-net (see Corollary 5).
+Moreover, since the partition is constructed using a grid, the cluster diameters are similar, so $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$ closely matches the spacing of the original grid—making the output resemble a slightly coarser grid.
+Partition-Independent Upper Bound.
+Generally speaking, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified.
+We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$:
+$N(\mathcal{A}_{\operatorname*{full}},\!\epsilon)=\min\Big{\{}m\!\in\!\mathbb{N}:$$\;\exists\{a_{\ell}\}_{\ell\leq m}\subseteq\mathcal{A}_{\operatorname*{full}},$
+$\forall a\in\mathcal{A}_{\operatorname*{full}},\exists\ell,\|a\!-\!a_{\ell}\|_{2}\leq\epsilon\Big{\}}.$
+Theorem 10.
+Under Assumption 3, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$.
 Let the action space have dimension $n$, and fix a constant $0<\epsilon<M$.
@@
 For the same constant $C>0$ in Theorem 7, and another absolute constant $c>0$,
-$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)},\quad\text{where }K\geq c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{full},\epsilon).$
+$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;$ $2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},$
+$\text{where }K\geq$ $c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon).$
 As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
-The partition-independent upper bound in Theorem 7 depends on a positive constant $\epsilon$, which serves as a tolerance parameter up to the diameter of the action space.
-Given a $\epsilon$, the required number of samples $K$ scales with the square of the diameter-to-$\epsilon$ ratio, multiplied by the covering number—a geometric property of the action space.
-The resulting expected regret is then bounded in terms of $\epsilon$, the dimensionality, and the logarithm of the covering number.
-As $\epsilon$ decreases, more samples are needed, but the regret bound becomes tighter.
-Further, we give a corresponding lower bound of Algorithm 1:
-Theorem 8 (Lower bound).
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 4.
-Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
-Let $\mathcal{A}$ be the output of Algorithm 1.
-For the same constant $C$ in Theorem 5 and another constant $c>0$,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).\end{split}$
-4 Examples
+[…]
+Lower Bound.
+Theorem 11.
+Under the same condition of Theorem 8. Let each cluster contains more than one action and satisfies Assumption 6. For the same constant $C$ in Theorem 7 and another absolute constant $c>0$,
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\\ &\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
+3.2 Regret Bound for Sub-Gaussian Process
+In fact, the Gaussian process assumption can be relaxed to a sub-Gaussian one, where the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
+$\forall u>0,\quad\Pr\left[|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|_{2}^{2}}\right).$
+Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional (Talagrand, 2014) for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm.
+It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$.
+Using this functional, we establish a regret bound for our algorithm:
+Theorem 12.
+Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process.
+Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space.
+Let $\mathcal{A}$ be output of Algorithm 1.
+Then, for a constant $C>0$,
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+4 Generalization and Empirical Validation
 In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings.
 Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling.
-We therefore adopt a dimension-free view by treating the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ as a family of random variables indexed by an abstract set $\mathcal{A}_{full}$.
-In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{full}$:
-$f(a):=\mu_{a},\quad\forall a\in\mathcal{A}_{full}.$ (4)
+We therefore adopt a dimension-free view by treating the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set $\mathcal{A}_{\operatorname*{full}}$.
+In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$:
+$f(a):=\mu_{a},\quad\forall a\in\mathcal{A}_{\operatorname*{full}}.$
 […]
@@
 where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues.
-Let outcome $f$ be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{full}$ and a weight vector $\alpha\in\mathbb{R}^{N}$.
-As shown in (williams2006gaussian, ; wainwright2019high, ), the function $f$ can also be written as
+Let outcome $f$ be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{\operatorname*{full}}$ and a weight vector $\alpha\in\mathbb{R}^{N}$.
+The function $f$ can be rewritten as
 $\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
@@
 Gaussian, this formulation can be represented as the canonical Gaussian process model in Equation (3).
-Note that in this case, the actual action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{full}$.
-Limitations: Gaussian assumptions may limit applicability in practice. We plan to extend this work to general random process with sub-Gaussian property.
+Note that in this case, the actual action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{\operatorname*{full}}$.
 4.1 Varying-dependence Actions
-We study the effect of varying action dependence using a Gaussian process with a kernel.
-To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel (williams2006gaussian, ).
+Assuming the outcome function is generated from a kernel inherently implies that the expected outcomes of actions are correlated.
+To study the effect of varying correlation levels among actions, we conduct the following experiments:
+Sampling Outcome Functions from a Kernel.
+We use the stationary RBF kernel (Williams and Rasmussen, 2006), and provide additional illustrations using the non-stationary Gibbs kernel in Supplementary:
 […]
-where $l$ is a length-scale parameter and $l(a):=0.1+0.9\cdot\exp(-\|a\|^{2})$ is a location-dependent length-scale function.
-Both of them control the dependence between actions.
-But, unlike the stationary RBF kernels, the Gibbs kernel allows some regions of the action space to be highly correlated over large distances, and others to have short-range correlations.
-That is, the correlation depends not only on the distance between actions, but also on where $a,a^{\prime}$ are located.
-When $l(a)=l$ is a constant, the Gibbs kernel reduces to the RBF kernel.
-Sampling Outcome Functions from a Kernel: We first construct the kernel matrix $\mathbf{K}$, where each entry is given by $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$, for $a,a^{\prime}\in\mathcal{A}_{full}$, depending on the choice of kernel.
-We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$.
-Under either kernels defined in Equation (5), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{full}$.
-Figure 2: Experiments with outcome functions sampled from RBF/Gibbs kernels in Equation (5).
-Left: Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in $[0,2]$ (blue curves, right y-axis).
-The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points.
-Right: Performance on solving stochastic combinatorial optimization in Equation (6), selecting $K=5$ distinct actions from 15 grid points in [0,2], with outcome functions sampled from RBF kernel at varying length-scales.
-Solid curves and shaded regions represent the mean and $\pm$ standard deviation of the expected regret over 50 repetitions.
-Results are presented for our method (green), Thompson Sampling (yellow), and UCB (orange), with TS/UCB run for 3000 rounds per repetition.
-To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval $[0,2]$, using Gibbs kernel in Equation (5) to sample outcome functions.
-To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid.
-The left subplot of Figure 2 provides examples of sampled outcome functions (blue curves, with the y-axis on the right-hand side), which become smoother as the actions approach the left end of the interval—indicating stronger correlations among function values in that region.
-We run Algorithm 1 on this action space with $K=5000$ to select actions and record the frequency of each action being selected.
-The resulting histogram (purple bars, with the y-axis on the left) reflects the importance measure $q$, highlighting that Algorithm 1 tends to select more actions from regions where the outcome functions are rougher—i.e., where action outcomes are less correlated and their features $\Phi(a)$ are are farther apart.
-Another interesting aspect of this subplot is the two high bars at the edges.
-Recall that the actual action space consists of feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{ful}$.
-For actions indexed closer to 0, their feature vectors become more densely packed compared to those indexed closer to 2, resulting in more correlated outcomes.
-The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
-As an extreme case, suppose $\mathcal{A}_{full}$ is a set of independent actions.
-This corresponds to the canonical process with the action space given by the orthogonal basis in $\mathbb{R}^{n}$ where $n=|\mathcal{A}_{full}|$.
-To see this, we associate each action $a$ with a unit vector $e_{a}$, which is $1$ at the $a^{th}$ entry but is zero otherwise, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$.
-With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ consists of mutually independent random variables.
-Then, as discussed in Supplementary I, for any subset $\mathcal{A}\subset\mathcal{A}_{full}$, including the output of Algorithm 1, the regret suffers from the lower bound
-$\mathbb{E}_{\theta}[\textnormal{Regret}]\geq\frac{\sqrt{\log|\mathcal{A}_{full}|}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}.$
-As a result, any algorithm, including Algorithm 1, would perform poorly unless the subset size $|\mathcal{A}|$ is sufficiently large.
-4.2 Combinatorial Optimization under Uncertainty
+where $l$ is a length-scale parameter that control the dependence between actions.
+By (Kanagawa et al., 2018, Theorem 4.12), we sample functions from a RKHS by first constructing the kernel matrix $\mathbf{K}$ with entries $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, and then drawing $f\sim\mathcal{N}(0,\mathbf{K})$.
+Figure 1: Performance comparison on stochastic combinatorial optimization in Equation (6), selecting $K=5$ distinct actions from 15 grid points, with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales.
+Left: Expected regret for our method (green), Thompson Sampling (yellow), UCB (orange), and Successive Halving (SH, blue), averaged over 50 repetitions at each length-scale.
+TS/UCB are run for 3000 rounds; SH is given a budget of 37,000 pulls.
+Right: SH’s expected regret (blue) and pull counts (yellow bars) per round.
+Our method’s performance (green) is shown for comparison.
+SH requires nearly 10,000 pulls—about 3× the number of super arms—to match our method, making it infeasible for large $\mathcal{A}_{\operatorname*{full}}$.
 Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in Equation (1).
-Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{full}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
-$\max_{\mathcal{A}\subseteq\mathcal{A}_{full}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (6)
+Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
+$\max_{\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (6)
 In general, this problem is non-trivial because the objective function involves an expectation over a collection of random variables whose distribution is neither specified nor directly accessible.
@@
 Our method is computationally efficient, avoiding the combinatorial complexity of the optimization problem in Equation (6).
-To assess the effectiveness in identifying near-optimal solutions, we compare it against Thompson Sampling (TS) and Upper Confidence Bound (UCB) on a simple example: We consider a fixed action space consisting of 15 grid points in $[0,2]$, using an RBF kernel in Equation (5) to sample outcome functions while varying the length-scale $l\in\{0.5,1,\dots,4\}$.
-In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where in each round, the optimal action is founded via an exhaustive search over the action space.
-In TS/UCB methods, we treat each $K$-tuple of actions as a super arm and adopt a bandit feedback setting: at each round, the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}$, and updates its policy accordingly, repeating this process for 3000 rounds.
-Thus, they are tailored to maximize the same objective.
-Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB, with the algorithms summarized in Supplementary J, along with examples of selected subsets and sampled outcome functions.
+To assess effectiveness in identifying near-optimal solutions, we compare our method against Thompson Sampling (TS) and Upper Confidence Bound (UCB) in a simple example, using Successive Halving (SH) as a reference for approximating the optimal subset—though SH becomes impractical when the full action space is large.
+We consider a fixed action space consisting of 15 grid points in [0,2][0,2][ 0 , 2 ], using an RBF kernel in Equation (5) to sample outcome functions while varying the length-scale $l\in\{0.5,1,\dots,4\}$.
+In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where the payoff of each action in a round is given by $\mu_{a}$.
+We use exhaustive search over the action space to find the optimal action $\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$.
+For other methods, we treat each $K$-tuple of actions as a super arm, and the payoff of each super arm in a round $\mathcal{A}$ is given by $\max_{a\in\mathcal{A}}\mu_{a}$.
+Thus, these methods are tailored to find the best super arm that maximizes $\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]$, aligning with the same objective.
+In TS/UCB methods, we adopt a bandit feedback setting: at each round, the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}$, and updates its policy accordingly, repeating this process for 3000 rounds, chosen to roughly match the number of super arms ($N=3003$).
+Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB.
+The algorithms are summarized in Supplementary, with the prior of TS set to $\mathcal{N}(0,1)$.
+In the SH method, all super arms are evaluated using a fixed budget of arm pulls (we use the minimum budget required in Karnin et al.
+(2013), $N\log_{2}N\approx 37{,}000$) over a few rounds.
+In each round, the number of remaining super arms is halved.
+This process continues until only the best-performing super arm remains or the budget is exhausted.
 We evaluate expected regret of all methods over $10^{5}$ randomly sampled outcome functions.
-The right subplot of Figure 2 reports the expected regret of our method (green), TS (yellow) and UCB (orange), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions.
+The left subplot of Figure 1 reports the expected regret of our method (green), TS (yellow), UCB (orange) and SH (blue), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions.
 As shown in this subplot, the expected regret decreases as the length-scale increases.
-This is because the length-scale $l$ of the RBF kernel controls the number of approximately independent actions in the process $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$: When actions $a$ and $a^{\prime}$ are far apart relative to the length-scale, the kernel value becomes very small, so the covariance between the function values $\mu_{a}$ and $\mu_{a^{\prime}}$ is also close to zero.
+This is because the length-scale $l$ of the RBF kernel controls the number of approximately independent actions in the process $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$: When actions $a$ and $a^{\prime}$ are far apart relative to the length-scale, the kernel value becomes very small, so the covariance between the function values $\mu_{a}$ and $\mu_{a^{\prime}}$ is close to zero.
+Regarding computational complexity: in a given bandit instance $\theta$, identifying the optimal action $a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)$ is a well-studied problem in both best-arm identification and GP optimization.
+However, if computing $a^{*}(\theta)$ is not computationally efficient, then identifying the optimal super arm, as in SH—whose search space grows combinatorially—is infeasible.
+Next, in the right subplot of Figure 1, we examine sample complexity.
+We run our method (green) and Successive Halving (SH, blue) for 50 additional repetitions, using a fixed length-scale $l=1$.
+Since SH proceeds in rounds, we track the current best super arm and the total number of super arm pulls so far after each round.
+The expected regret of the current best super arm, is evaluated over another $10^{5}$ randomly sampled outcome functions.
+The blue solid curves and shades for SH represent the mean $\pm$ one standard deviation of expected regret over 50 repetitions over each SH round.
+The yellow bars (corresponding to the right y-axis) show the number of super arm pulls by SH at the end of each round.
+For comparison, the green curve and shaded region represent the mean $\pm$ one standard deviation of expected regret for our method over the same 50 repetitions.
+As shown, SH’s performance roughly matches ours after three rounds, requiring 9005 super arm pulls—approximately three times the number of super arms.
+If the action space $|\mathcal{A}_{\operatorname*{full}}|$ is already large, running SH is intractable.
+4.2 Conclusion and Future Work
+We proposed a framework for selecting a subset of correlated actions, modeling payoff correlations with a Gaussian process and extending to the sub-Gaussian case.
+A simple algorithm was introduced and shown to effectively identify near-optimal subsets.
+A key direction for future work is developing a stopping criterion.
+When the subset size is flexible but sampling new bandit instances is costly, ideas from species discovery Roswell et al.
+(2021) may be useful—treating the discovery of a new action as analogous to discovering a new species.
+One could use Turing’s formula to estimate the probability of finding a new action, or frame the process as a one-armed bandit that stops when the discovery probability becomes sufficiently low.
 Appendix A Connection to Bayesian Regret of Multi-Armed Bandits
-Given a subset $\mathcal{A}\subseteq\mathcal{A}_{full}$, let $\mu^{*}(\mathcal{A},\theta):=\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ be the highest expected outcome in a bandit instance $\theta$.
+Given a subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$, let $\mu^{*}(\mathcal{A},\theta):=\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ be the highest expected outcome in a bandit instance $\theta$.
 We can equivalently rewrite the regret in Equation (1):
-$\textnormal{Regret}(\theta)=\mu^{*}(\mathcal{A}_{full},\theta)-\mu^{*}(\mathcal{A},\theta).$
-Bandit Regret: In the classic multi-armed bandit setting [7, 32], outcome of an arm $a\in\mathcal{A}_{full}$ in a bandit instance $\theta$ is generated i.i.d.
+$\textnormal{Regret}(\theta)=\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\mu^{*}(\mathcal{A},\theta).$
+Bandit Regret: In the classic multi-armed bandit setting [Auer et al., 2002a, Lattimore and Szepesvári, 2020], outcome of an arm $a\in\mathcal{A}_{\operatorname*{full}}$ in a bandit instance $\theta$ is generated i.i.d.
 according to an unknown but fixed law with expectation $\mu_{a}(\theta)$.
@@
 The bandit regret of policy $\pi$ in this bandit environment $\theta$ over $N$ rounds is defined as the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action and following a policy $\pi$ to choose actions.
-Bayesian (bandit) regret [2, 32] is the expectation of this bandit regret, taken over the randomness in the distribution of bandit instances, and the action chosen by a policy:
-$\begin{split}\textnormal{BanditRegret}(\theta):=&\mu^{*}(\mathcal{A}_{full},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta),\\ \textnormal{BayesianRegret}:=&\mathbb{E}\left[\mu^{*}(\mathcal{A}_{full},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta)\right],\end{split}$
+Bayesian bandit regret [Agrawal and Goyal, 2012, Lattimore and Szepesvári, 2020] is the expectation of this bandit regret, taken over the randomness in the distribution of bandit instances, and the action chosen by the policy:
+$\begin{split}\textnormal{BanditRegret}(\theta):=&\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta),\\ \textnormal{BayesianBanditRegret}:=&\mathbb{E}\!\left[\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta)\right],\end{split}$
 where $A_{t}$ is the action chosen by policy $\pi$ in round $t$.
-Since we consider settings where only a subset of the full action space is accessible, the action chosen by the policy can only from the subset $\mathcal{A}\subseteq\mathcal{A}_{full}$.
-This quantifies the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action from the full action set $\mathcal{A}_{full}$ and following policy $\pi$ to choose actions from the restricted subset $\mathcal{A}$.
-Next, we establish the connection between the regret defined in this paper and the Bayesian (bandit) regret.
+Since we consider settings where only a subset of the full action space is accessible, the action chosen by the policy can only from the subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$.
+This quantifies the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action from the full action set $\mathcal{A}_{\operatorname*{full}}$ and following policy $\pi$ to choose actions from the restricted subset $\mathcal{A}$.
+Next, we establish the connection between the regret defined in this paper and the Bayesian bandit regret.
 A.1 Optimal policy
 If $\pi$ is the optimal policy, it selects the best action in $\mathcal{A}$ in every round.
-Therefore, the regret of only having access to $\mathcal{A}$ instead of $\mathcal{A}_{full}$ for $N$ rounds, is determined by the optimal actions in $\mathcal{A}$ and $\mathcal{A}_{full}$:
+Therefore, the regret of only having access to $\mathcal{A}$ instead of $\mathcal{A}_{\operatorname*{full}}$ for $N$ rounds, is determined by the optimal actions in $\mathcal{A}$ and $\mathcal{A}_{\operatorname*{full}}$:
 […]
@@
 […]
-$=$ $\mathbb{E}\left[\sum_{t=1}^{N}\left(\mu^{*}(\mathcal{A})-\mu_{A_{t}}+\mu^{*}(\mathcal{A}_{full})-\mu^{*}(\mathcal{A})\right)\right]$
+$=$ $\mathbb{E}\left[\sum_{t=1}^{N}\left(\mu^{*}(\mathcal{A})-\mu_{A_{t}}+\mu^{*}(\mathcal{A}_{\operatorname*{full}})-\mu^{*}(\mathcal{A})\right)\right]$
 $\leq$ $C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}],$
 where $|\mathcal{A}|$ denotes the cardinality, and $C>0$ is a constant.
-The inequality uses the Bayesian (bandit) regret of Thompson sampling (Theorem 36.1 in [32]).
-Note that with access to the full action space, the Bayesian regret is bounded by $C\sqrt{|\mathcal{A}_{full}|\cdot N\log N}$.
+The inequality uses the Bayesian (bandit) regret of Thompson sampling (Theorem 36.1 in Lattimore and Szepesvári [2020]).
+Note that with access to the full action space, the Bayesian regret is bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$.
 However, this can be worse than restricting to a subset $\mathcal{A}$ when $\mathbb{E}_{\theta}[\textnormal{Regret}]$ is sufficiently small and the time horizon $N$ is finite.
 Intuitively speaking, using a well-chosen subset of representative actions instead of the full action space can reduce Bayesian regret in multi-armed bandits.
-Appendix B Canonical Representation of a Gaussian Process
-The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, such that:
-$\mu_{a}:=\langle a,g\rangle+c,\quad\;\forall\;a\in\mathcal{A}_{full},\;g\sim\mathcal{N}(0,\Sigma).$ (7)
-We show how to transform it into the canonical Gaussian process. Note that
-$g\equiv\Sigma^{1/2}\theta,\;\theta\sim\mathcal{N}(0,I),$
-where $\equiv$ denotes equivalence in distribution. Let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$. Thus, $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$, and
-$\mu_{a}=\langle a,g\rangle=\sum_{j\leq n}a_{j}g_{j}=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
-Therefore, the setting in Equation (7) is equivalent to
-$\mu_{a}=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{full},\;\theta\sim\mathcal{N}(0,I),$
-where $\Sigma^{1/2}\mathcal{A}_{full}$ is the projected set of $\mathcal{A}_{full}$ by the matrix $\Sigma^{1/2}$.
-Appendix C Convex hull dominated by extreme points
-Lemma 9.
-Given a subset $\mathcal{A}\subset\mathcal{A}_{full}$.
-If $\mathcal{A}$ includes all extreme points of $\operatorname*{Conv}(\mathcal{A}_{full})$, i.e., $\mathcal{A}_{full}\subseteq\operatorname*{Conv}(\mathcal{A})$, then it always holds
-$\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=0\;\;\forall\theta\in\mathbb{R}^{n}.$
-Proof.
-Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{K}\}$.
-Since $\mathcal{A}_{full}\subseteq\operatorname*{Conv}(\mathcal{A})$, for any $a\in\mathcal{A}_{full}$, there exist non-negative numbers $\lambda_{1},\dots,\lambda_{K}$ that sum up to $1$ and such that $a=\lambda_{1}a_{1}+\dots+\lambda_{K}a_{K}$.
-Thus, by Equation (2), for any $\theta\in\mathbb{R}^{n}$
-$\mu_{a}=\langle a,\theta\rangle=\left\langle\sum_{i\in 1:K}\lambda_{i}a_{i},\theta\right\rangle=\sum_{i\in 1:K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i\in 1:K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.$
-It yields
-$\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{full}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$
-where the first inequality uses $\mathcal{A}\subset\mathcal{A}_{full}$. Thus, the two quantities are equal. ∎
-Appendix D Proof of Lemma 3
+Appendix B Proof of Lemma 4
 Statement: Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$.
@@
 where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$.
-The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [46].
+The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [TOPSØE1, 2007].
 Therefore,
-$\Pr[r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon]=$ $1-\Pr[r\cap\mathcal{A}\neq\emptyset\quad\exists r\in R_{\epsilon}]$
+$\Pr[r\cap\mathcal{A}\neq\emptyset\;\text{whenever}\;q(r)>\epsilon]$
+$=$ $1-\Pr[r\cap\mathcal{A}\neq\emptyset\;\exists r\in R_{\epsilon}]$
 $\geq$ $1-\sum_{r\in R_{\epsilon}}\Pr[r\cap\mathcal{A}\neq\emptyset]$
@@
 ∎
-Appendix E Proof of Theorem 5
+Appendix C Proof of Theorem 7
 Lemma 13 (Regret decomposition of reference subsets).
@@
 Let the cluster $r_{\ell}$ contains the optimal action, i.e., $a^{*}(\theta)\in r_{\ell}$, then
-$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell},$
-where the middle equality uses that when $a^{*}(\theta)\in r_{\ell}$, $\max_{a\in\mathcal{A}_{full}}\mu_{a}=\max_{a\in r_{\ell}}\mu_{a}$ and further with Assumption 4, $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\geq\mu_{a_{\ell}}$ when $a^{*}(\theta)\in r_{\ell}$.
+$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell},$
+where the middle equality uses that when $a^{*}(\theta)\in r_{\ell}$, $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\max_{a\in r_{\ell}}\mu_{a}$ and further with Assumption 6, $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\geq\mu_{a_{\ell}}$ when $a^{*}(\theta)\in r_{\ell}$.
 Then,
@@
 Proof.
-Let $u_{0}:=\epsilon\sqrt{\log c}$, then we have
+[…]
 […]
@@
 $=$ $u_{0}+\exp\left(-\frac{u_{0}^{2}}{\epsilon^{2}}\right)\cdot\frac{c\epsilon^{2}}{2u_{0}}$
-$=$ $\epsilon\sqrt{\log c}+\frac{\epsilon}{2\sqrt{\log c}}\leq C\epsilon\sqrt{\log c}.$
-Lemma 12 (Borell-TIS inequality; Lemma 2.4.7 [45]).
+$=$ $\epsilon\sqrt{\log c}+\frac{\epsilon}{2\sqrt{\log c}}\leq C\epsilon\sqrt{\log c},$
+where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of Vershynin [2018]). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
+Lemma 15 (Borell-TIS inequality; Lemma 2.4.7 of [Talagrand, 2014]).
 Given a set $\mathcal{S}$, and a zero-mean Gaussian process $\left(X_{a}\right)_{a\in\mathcal{S}}$.
@@
 $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
-Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{full}$, it holds that $\mathbb{E}Y_{\ell}=\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}$.
+Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, it holds that $\mathbb{E}Y_{\ell}=\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}$.
 When the cluster $r_{\ell}$ contains the optimal action, i.e., $a^{*}(\theta)\in r_{\ell}$, we have
-$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell}\leq\max_{\ell^{\prime}\leq m}Y_{\ell^{\prime}},$
+$\operatorname*{Regret}=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell}\leq\max_{\ell^{\prime}\leq m}Y_{\ell^{\prime}},$
 where the first equality follows from the definition of regret in Equation (1), and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$.
@@
 ∎
-Appendix F Proof of Theorem 6
+Appendix D Proof of Theorem 8
 Lemma 16 (Transformation invariance).
@@
 Proof.
-$\mathbb{E}\left[\max_{a\in\mathcal{S}+x}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime}+x,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right]+\mathbb{E}\left[\langle x,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
+$\mathbb{E}\left[\max_{a\in\mathcal{S}+x}\langle a,\theta\rangle\right]=$ $\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime}+x,\theta\rangle\right]$
+$=$ $\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right]+\mathbb{E}\left[\langle x,\theta\rangle\right]$
+$=$ $\mathbb{E}\left[\max_{a^{\prime}\in\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
 where the last equality uses $\mathbb{E}\left[\langle x,\theta\rangle\right]=\left\langle x,\mathbb{E}[\theta]\right\rangle=0$. ∎
@@
 Since $\mathcal{A}$ is the output of Algorithm 1, the event $\{a^{*}(\theta)\in r\}$ is independent from $\{r\cap\mathcal{A}\!\neq\!\emptyset\}$ or $\{r\cap\mathcal{A}\!=\!\emptyset\}$. We have
-$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!=\!\emptyset]=\Pr[a^{*}(\theta)\in r]\Pr[r\cap\mathcal{A}\!=\!\emptyset]=q(r)\Pr[r\cap\mathcal{A}\!=\!\emptyset]=q(r)(1-q(r))^{K},$
+$\Pr[a^{*}(\theta)\in r,r\cap\mathcal{A}\!=\!\emptyset]=$ $\Pr[a^{*}(\theta)\in r]\Pr[r\cap\mathcal{A}\!=\!\emptyset]$
+$=$ $q(r)\Pr[r\cap\mathcal{A}\!=\!\emptyset]$
+$=$ $q(r)(1-q(r))^{K},$
 where the first equality uses independence between $\{a^{*}(\theta)\in r\}$ and $\{r\cap\mathcal{A}\!\neq\!\emptyset\}$, the second equality uses the definition of measure $q$, the third equality use the probability of missing cluster $r$ in $K$ i.i.d.
@@
 Then, for the same constant $C>0$ in Theorem 7,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
 Proof.
@@
 $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
-Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{full}$, we have $\mathbb{E}Y_{\ell}=\mathbb{E}\sup_{a\in r_{\ell}}\mu_{a}$.
+Since $\mathbb{E}\mu_{a}=0$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, we have $\mathbb{E}Y_{\ell}=\mathbb{E}\sup_{a\in r_{\ell}}\mu_{a}$.
 Consider the case that $a^{*}(\theta)\in r_{\ell}$. We have
-$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\leq&\mathbb{E}\left[Y_{\ell}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ =&\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\\ \leq&\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (13)
+$\begin{split}&\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ \leq&\mathbb{E}\left[Y_{\ell}\Big{|}r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ =&\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\\ \leq&\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (12)
 where the first inequality uses $\operatorname*{Regret}\leq Y_{\ell}$ if $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$, the equality uses that $Y_{\ell}$ is independent of $r_{\ell}\cap\mathcal{A}\!\neq\!\emptyset$.
@@
 Therefore,
-$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\leq\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (14)
-where the inequality uses $\operatorname*{Regret}\leq\max_{a\in\mathcal{A}_{full}}\mu_{a}$, as a consequence of $0\in\operatorname*{Conv}(\mathcal{A})$, and that $\max_{a\in\mathcal{A}_{full}}\mu_{a}$ is independent of $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$.
+$\begin{split}&\mathbb{E}\left[\operatorname*{Regret}\Big{|}r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ \leq&\mathbb{E}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (13)
+where the inequality uses $\operatorname*{Regret}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$, as a consequence of $0\in\operatorname*{Conv}(\mathcal{A})$, and that $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ is independent of $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$.
 Further,
-$\mathbb{E}[\operatorname*{Regret}]=$ $\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]$
+$\mathbb{E}[\operatorname*{Regret}]=$ $\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]$ (14)
 $+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big{|}r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]$
-$\leq$ $\sum_{\ell\leq m}q(r_{\ell})\cdot\left(\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]+(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\right)$
-$=$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r\right]$
-$\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}\max_{a\in r}\mu_{a}+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2},$
+$\leq$ $\sum_{\ell\leq m}q(r_{\ell})\cdot\left(\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]+(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\right)$
+$=$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\Big{|}a^{*}(\theta)\in r\right]$
+$\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
 where the first equality uses tower rule.
-The first inequality uses Equations (13-14), and Lemma 14.
+The first inequality uses Equations (12-13), and Lemma 17.
 The last equality uses tower rule again.
-The last inequality uses Equation (11) in the proof of Theorem 5 on the summand $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$.
-The other summand is bounded via Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$.
+The last inequality uses Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$.
+Note that the Gaussian process assumption is not used in Equation (14); it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$.
+By applying Equation (10) in the proof of Theorem 7, we bound this term and complete the proof.
 ∎
-Appendix G Proof of Theorem 7
+Appendix E Proof of Theorem 10
 Lemma 18.
 Let $q$ denote a discrete probability distribution over a finite support $\mathcal{R}$. Define
-$M:=\max_{q}\mathbb{E}_{q}\left[(1-q(r))^{K}\right]\quad\text{s.t. }\sum_{r\in\mathcal{R}}q(r)=1,\;q(r)\in[0,1)\;\forall r\in\mathcal{R}.$ (15)
+$\begin{split}M:=\max_{q}&\;\mathbb{E}_{q}\left[(1-q(r))^{K}\right]\\ \textnormal{s.t.}&\sum_{r\in\mathcal{R}}q(r)=1,\;q(r)\in[0,1)\;\;\forall r\in\mathcal{R}.\end{split}$ (15)
 When $|\mathcal{R}|\geq K+1$, the maximum is $M=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$, and it is attained when $q$ is uniform.
@@
 $f^{\prime}(q_{\ell})=(1-q_{\ell})^{K-1}\big{(}1-(K+1)q_{\ell}\big{)},$
-which is positive on the interval $\left[0,\frac{1}{K+1}\right)$ and negative on $\left(\frac{1}{K+1},1\right)$. Therefore, $f(q_{\ell})$ attains its maximum over $[0,1)$ at
+which is positive on the interval $\left[0,\frac{1}{K+1}\right)$ and negative on the interval $\left(\frac{1}{K+1},1\right)$. Therefore, $f(q_{\ell})$ attains its maximum over [0,1)[0,1)[ 0 , 1 ) at
 $q^{*}_{\ell}=\frac{1}{K+1},$
@@
 Thus,
-$M\geq\sum_{\ell=1}^{m-1}f\left(\frac{1}{K+1}\right)+f\left(\frac{K-m+2}{K+1}\right)\geq\frac{m-1}{K+1}\left(\frac{K}{K+1}\right)^{K},$
+$M\geq$ $\sum_{\ell=1}^{m-1}f\left(\frac{1}{K+1}\right)+f\left(\frac{K-m+2}{K+1}\right)$
+$\geq$ $\frac{m-1}{K+1}\left(\frac{K}{K+1}\right)^{K},$
 where the right inequality uses that $f(q_{\ell})\geq 0$ for $q_{\ell}\in[0,1]$.
@@
 $\frac{\partial^{2}}{\partial q_{\ell}^{2}}\left[q_{\ell}(1-q_{\ell})^{K}\right]=-(1-q_{\ell})^{K-2}\left(2K-K(K+1)q_{\ell}\right),$
-which is negative for $q_{\ell}\leq\frac{1}{K+1}$, because then $02-(K+1)q_{\ell}>$ 0. Hence, the Hessian is negative definite, and the critical point is a local (and thus global) maximum.
+which is negative for $q_{\ell}\leq\frac{1}{K+1}$, because then $2-(K+1)q_{\ell}>0$. Hence, the Hessian is negative definite, and the critical point is a local (and thus global) maximum.
 Finally, note that Equation (16) is a relaxation of Equation (15).
@@
 Statement of Theorem 10
-Under Assumption 2, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{full}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$.
+Under Assumption 3, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$.
 Let the action space have dimension $n\in\mathbb{N}$, and fix a constant $0<\epsilon<M$.
@@
 For the same constant $C>0$ in Theorem 7, we have:
-$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)},\quad\text{where }K\geq\frac{1}{2}\left(\frac{M^{2}N(\mathcal{A}_{full},\epsilon)}{\epsilon^{2}e}-1\right).$
+$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\quad\text{where }K\geq\frac{1}{2}\left(\frac{M^{2}N(\mathcal{A}_{\operatorname*{full}},\epsilon)}{\epsilon^{2}e}-1\right).$
 As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
@@
 By Lemma 16, we can shift the action space to be centered at the origin.
-So, without loss of generality, we assume that $\mathcal{A}_{full}\subset M\cdot B_{2}^{n}$, the scaled unit Euclidean ball in $\mathbb{R}^{n}$.
+So, without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}\subset M\cdot B_{2}^{n}$, the scaled unit Euclidean ball in $\mathbb{R}^{n}$.
 Then,
-$\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}^{2}\right]\leq M^{2}\cdot\mathbb{E}\|\theta\|_{2}^{2}=M^{2}n,$ (17)
+$\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}^{2}\right]\leq M^{2}\cdot\mathbb{E}\|\theta\|_{2}^{2}=M^{2}n,$ (17)
 where the inequality follows from $\mu_{a}=\langle\theta,a\rangle\leq M\|\theta\|_{2}$ for all $a\in MB_{2}^{n}$, and the equality uses that $\theta\sim\mathcal{N}(0,I)$, so each coordinate has variance 1 and $\|\theta\|_{2}^{2}=\sum_{i=1}^{n}\theta_{i}^{2}$ has expectation $n$.
-Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{full}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{full},\epsilon)$.
+Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{\operatorname*{full}}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{\operatorname*{full}},\epsilon)$.
 Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
-$r_{\ell}=\left\{a\in\mathcal{A}_{full}:\pi(a)=a_{\ell}\right\}.$
+$r_{\ell}=\left\{a\in\mathcal{A}_{\operatorname*{full}}:\pi(a)=a_{\ell}\right\}.$
 Then,
@@
 where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 16 to shift; and the final bound uses Claim 3 from Supplementary K. Then,
-$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2},$
+$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
 $\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}\cdot M\sqrt{n}$
@@
 The last inequality uses the assumption $K\geq\frac{1}{2}\left(\frac{mM^{2}}{\epsilon^{2}e}-1\right)$.
-Finally, since $\mathcal{A}_{full}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{full},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of [49].
-Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$.
+Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of Vershynin [2018].
+Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$.
 ∎
-Appendix H Proof of Theorem 8
+Appendix F Proof of Theorem 11
 Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 6.
@@
 The second inequality follows from Equation (19).
-The third inequality uses the Pólya-Szegő inequality [17], corresponding to a constant $c>0$:
+The third inequality uses the Pólya-Szegő inequality [Dragomir, 2004], corresponding to a constant $c>0$:
 $c:=\frac{1}{2}\left(\sqrt{\frac{M_{1}M_{2}}{m_{1}m_{2}}}+\sqrt{\frac{m_{1}m_{2}}{M_{1}M_{2}}}\right),$
@@
 ∎
-Appendix I Example of orthogonal basis
-In general, even when the distribution of bandit instances $\theta$ is well-behaved, e.g., the standard Gaussian distribution in $\mathbb{R}^{n}$, the importance measure $q$ is usually non-trivial.
-In this section, we consider an example of the action space being orthogonal basis.
-Let the set $\mathcal{A}_{full}=\{e_{i}:i=1,\dots,n\}$ consists of the $n$ unit vectors aligned with the coordinate axes in $\mathbb{R}^{n}$.
-Hence, $\operatorname{diam}(\mathcal{A}_{full})=\sqrt{2}$.
-In this case, $\max_{a\in\mathcal{A}_{full}}\langle a,\theta\rangle$ is equivalent to the maximum among the $n$ entries of $\theta$, where by definition, each entry follows i.i.d.
-from the standard normal distribution $\mathcal{N}(0,1)$, such that each entry has the same probability of having the maximum value.
-Let each cluster contain only one unit vector, thus the importance measure of clusters $q$ is a uniform distribution.
-Expected maximum in $\mathcal{A}_{full}$: Let $X_{i}$, $i=1,\dots,n$ be i.i.d. samples from $\mathcal{N}(0,1)$, then
-$\mathbb{E}\max_{a\in\mathcal{A}_{full}}\mu_{a}=\mathbb{E}_{\theta}\max_{i=1,\dots,n}\theta_{i}=\mathbb{E}\max_{i=1,\dots,n}X_{i},$
+Appendix G Proof of Theorem 12
+Definition 2.
+The random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a mean-zero sub-Gaussian process if the process $\mathbb{E}\mu_{a}=0$ and has the increment condition:
+$\forall u>0,\Pr[|\mu_{a}-\mu_{a^{\prime}}|\geq u]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|^{2}_{2}}\right).$
+Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional [Talagrand, 2014] for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm.
+It is well-known that under the assumption that $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a Gaussian process, this gives the tightest bound, for a universal constant $L$:
+$\frac{1}{L}\gamma_{2}(\mathcal{S})\leq\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq L\gamma_{2}(\mathcal{S}).$
+Within the proof of the upper bound for general sub-Gaussian process, it also derives the deviation bound for the term $\max_{a\in\mathcal{S}}\mu_{a}$:
+Lemma 19 (Theorem 2.2.22 of Talagrand [2014]).
+Let $\{\mu_{a}\}_{a\in\mathcal{S}}$ be a mean-zero sub-Gaussian process, then there exists a constant $c>0$:
+$\Pr\left[\sup_{a,a^{\prime}\in\mathcal{S}}|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{cu^{2}}{\gamma^{2}_{2}(\mathcal{S})}\right).$
+Lemma 20.
+Let $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ be a mean-zero sub-Gaussian process.
+Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, and an arbitrary reference subset $\mathcal{A}:=\{a_{\ell}\}_{\ell\leq m}$.
+Fix $\ell$, define a non-negative random variable
+$Y_{\ell}:=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
+Then, for some constant $C>0$,
+$\mathbb{E}_{\theta}\left[\max_{\ell\leq m}Y_{\ell}\right]\leq C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell}).$
+Proof.
+Let $\epsilon:=\max_{\ell\leq m}\gamma_{2}(r_{\ell})$. For any $\ell\leq m$, we have
+$\Pr\left[|Y_{\ell}|\geq u\right]=$ $\Pr\left[\left|\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right|\geq u\right]$
+$\leq$ $\Pr\left[\sup_{a,a^{\prime}\in r_{\ell}}|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]$
+$\leq$ $2\exp\left(-\frac{cu^{2}}{\gamma^{2}_{2}(r_{\ell})}\right)$
+$\leq$ $2\exp\left(-\frac{cu^{2}}{\epsilon^{2}}\right),$
+where the equality uses definition of $Y_{\ell}$. The first inequality uses
+$\left|\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right|\leq\sup_{a\in r_{\ell}}|\mu_{a}-\mu_{a_{\ell}}|\leq\sup_{a,a^{\prime}\in r_{\ell}}|\mu_{a}-\mu_{a^{\prime}}|.$
+The second inequality uses the fact that $\{\mu_{a}\}_{a\in r_{\ell}}$ is a mean-zero sub-Gaussian process and applies Lemma 19.
+The last inequality uses the definition of $\epsilon$, and $c>0$ is a constant.
+By union bound, we have
+$\Pr\left[\max_{\ell\leq m}|Y_{\ell}|\geq u\right]\leq 2m\exp\left(-\frac{cu^{2}}{\epsilon^{2}}\right).$
+Further, using Lemma 14, we have for some absolute constant $C>0$:
+$\mathbb{E}\left[\max_{\ell\leq m}|Y_{\ell}|\right]\leq C\epsilon\sqrt{\log m}.$
+Statement of Theorem 12:
+Let $\mathcal{A}$ be output of Algorithm 1.
+Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process.
+Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space.
+Then, for the same constant $C>0$,
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+Proof.
+If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$.
+If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative.
+The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1.
+For each $\ell\leq m$, define a random variable
+$Y_{\ell}:=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
+Following the same reasoning used in the proof of Theorem 8:
+$\mathbb{E}[\operatorname*{Regret}]\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}$
+$\leq$ $C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
+The first inequality follows from Equation (14) in the proof of Theorem 8. The last inequality uses Lemma 20. ∎
+Appendix H The effect of clustering structure on regret
+In Figure 2, we study the effect of the cluster diameters (controlled by a spread parameter) on regret.
+We structure the clustered action space: Five center points are fixed on the unit sphere in $\mathbb{R}^{3}$, and around each center, 200 points are sampled to form five clusters.
+Each point is obtained by adding Gaussian noise (mean zero, standard deviation equal to the spread parameter) to the center direction, followed by projection back onto the unit sphere.
+Bandits are sampled from a 3-dimensional standard Gaussian distribution, i.e., $\theta\sim\mathcal{N}(0,I)$.
+The left subplot shows the expected regret of Algorithm 1 with $K=10$, computed using $10^{4}$ additional bandits, as the spread varies from 0.01 to 0.5.
+The curves and error shade represent the mean $\pm$ one standard deviation of expected regret over 30 repetitions.
+The middle and right subplots display example action spaces for spread values of 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
+Figure 2: Illustration of clustered action spaces on unit sphere in $\mathbb{R}^{3}$ and the effect of cluster diameters on regret.
+Five clusters are formed by generating 5 fixed center points, with 200 points sampled around each using Gaussian noise (spread controls the variance).
+Bandits are drawn from $\mathcal{N}(0,I)$.
+The left subplot shows the mean $\pm$ standard deviation of the expected regret (over 30 trials) as the spread varies from 0.01 to 0.5, using $10^{4}$ additional bandits.
+The middle and right subplots show example action spaces (blue dots) for spread values 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
+Appendix I Varying-dependence actions with RBF/Gibbs kernels
+We study the effect of varying action dependence using a Gaussian process with a kernel.
+To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel [Williams and Rasmussen, 2006].
+[…]
+where $l$ is a length-scale parameter and $l(a):=0.1+0.9\cdot\exp(-\|a\|^{2})$ is a location-dependent length-scale function.
+Both of them control the dependence between actions.
+But, unlike the stationary RBF kernels, the Gibbs kernel allows the correlation to depend not only on the distance between actions, but also on their locations.
+When $l(a)=l$ is a constant, the Gibbs kernel reduces to the RBF kernel.
+Sampling Outcome Functions from a Kernel: We first construct the kernel matrix $\mathbf{K}$, where each entry is given by $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$, for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, depending on the choice of kernel.
+We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$.
+Under either kernels defined in Equation (20), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{\operatorname*{full}}$.
+In this way, we sample functions from a RKHS function class; See [Kanagawa et al., 2018, Theorem 4.12].
+Figure 3: Experiments with outcome functions sampled from RBF/Gibbs kernels in Equation (20).
+Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in [0,2][0,2][ 0 , 2 ] (blue curves, right y-axis).
+The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points.
+To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval [0,2][0,2][ 0 , 2 ], using Gibbs kernel in Equation (20) to sample outcome functions.
+To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid.
+Figure 3 provides examples of sampled outcome functions (blue curves, with the y-axis on the right-hand side), which become smoother as the actions approach the left end of the interval—indicating stronger correlations among function values in that region.
+We run Algorithm 1 on this action space with $K=5000$ to select actions and record the frequency of each action being selected.
+The resulting histogram (purple bars, with the y-axis on the left) reflects the importance measure $q$, highlighting that Algorithm 1 tends to select more actions from regions where the outcome functions are rougher—i.e., where action outcomes are less correlated and their features $\Phi(a)$ are are farther apart.
+Another interesting aspect of this subplot is the two high bars at the edges.
+Recall that the actual action space consists of feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{ful}$.
+For actions indexed closer to 0, their feature vectors become more densely packed compared to those indexed closer to 2, resulting in more correlated outcomes.
+The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
+Appendix J Independent and identically distributed actions
+As an extreme case, suppose that $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ is a set of i.i.d.
+random variables.
+This corresponds to the canonical process in which the action space is given by the orthonormal basis of $\mathbb{R}^{n}$, where $n=|\mathcal{A}_{\operatorname*{full}}|$.
+To see this, we associate each action $a$ with a unit vector $e_{a}$, which has a value of 111 at the $a$th coordinate and 0 elsewhere, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$.
+With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ consists of mutually independent random variables.
+Let the set $\mathcal{A}_{\operatorname*{full}}=\{e_{i}:i=1,\dots,n\}$ denote the $n$ unit vectors aligned with the coordinate axes in $\mathbb{R}^{n}$.
+Hence, $\operatorname{diam}(\mathcal{A}_{\operatorname*{full}})=\sqrt{2}$.
+In this case, $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\langle a,\theta\rangle$ is equivalent to the maximum among the $n$ entries of $\theta$, where each entry is i.i.d.
+from the standard normal distribution $\mathcal{N}(0,1)$.
+By symmetry, each coordinate has the same probability of attaining the maximum value.
+If each cluster contains only one unit vector, then the importance measure $q$ over clusters is the uniform distribution.
+Expected maximum in $\mathcal{A}_{\operatorname*{full}}$: Let $X_{i}$, $i=1,\dots,n$ be i.i.d. samples from $\mathcal{N}(0,1)$, then
+$\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\mathbb{E}_{\theta}\max_{i=1,\dots,n}\theta_{i}=\mathbb{E}\max_{i=1,\dots,n}X_{i},$
 where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary L, we have
-$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}_{full}}\mu_{a}\leq\sqrt{2\log n},\quad\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}}\mu_{a}\leq\sqrt{2\log|\mathcal{A}|}.$
+$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\sqrt{2\log n},\quad\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}}\mu_{a}\leq\sqrt{2\log|\mathcal{A}|}.$
 Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in Equation (1),
 $\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}\leq\mathbb{E}_{\theta}[\textnormal{Regret}]\leq\sqrt{2\log n}-\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}.$
-Upper bound of expected regret of Algorithm 1: Since the importance measure of clusters $q$ is a uniform distribution, we have
-$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\left(1-\frac{1}{n}\right)^{K}\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right]\leq\left(1-\frac{1}{n}\right)^{K}\sqrt{2\log n},$
-where the left inequality uses Corollary 16 and that if each cluster contains one action, $\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]=0$ for each cluster and $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)=0$.
-Corollary 16.
-Let $\mathcal{A}$ be output of Algorithm 1.
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where the $q$ measure is uniform over all clusters.
-Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
-Then, for the same constant $C>0$ in Theorem 5, it holds that
-$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right].$
-Under Assumption 4,
-$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).$
-Proof.
-Consider the variables $Y_{\ell}$ defined in the proof of Theorem 6,
-$\mathbb{E}[\operatorname*{Regret}]\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\sum_{r\in\mathcal{R}}q(r)\cdot\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r\right]$
-$=$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right],$
-where the first inequality comes from the proof of Theorem 6, the equality uses $q(r)=1/|\mathcal{R}|$ for all $r\in\mathcal{R}$.
-Then, using Equation (11) in the proof of Theorem 5, we derive the upper bound.
-Consider the variables $Y_{\ell}$ defined in the proof of Theorem 8,
-$\mathbb{E}[\operatorname*{Regret}]\geq$ $\sum_{\ell\leq m}q(r_{\ell})\cdot\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]$
-$=$ $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\geq\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right],$
-where the first inequality comes from the proof of Theorem 8, the equality uses $q(r)=1/|\mathcal{R}|$ for all $r\in\mathcal{R}$.
-The last inequality uses the definition of $q$ measure and $Y_{\ell}\geq\min_{\ell\leq m}Y_{\ell}$.
-Then, using Equation (12) in the proof of Theorem 5, we derive the lower bound.
-∎
-Appendix J Comparisons with UCB & TS
-To understand why Algorithm 1 outperforms UCB and TS, we still consider the fixed action space of 15 grid points over the interval $[0,2]$ and indexed by $0,1,\ldots,14$.
-Outcome functions are generated using an RBF kernel with a fixed length-scale $l=1$.
-The middle upper subplot of Figure 4 displays ten example outcome functions sampled from this setup.
-In each trial, Algorithms 1, 2 (UCB), and 3 (TS) are run to select $K=5$ distinct representative actions.
-UCB and TS are each executed for 3000 rounds (excluding initialization), with each round corresponding to a bandit instance.
-The number of rounds is chosen to approximately match the number of possible super arms, given by $\binom{15}{5}=3003$.
-The upper part of Figure 3 visualizes the representative actions selected by each algorithm in three individual trials.
-Each row corresponds to one trial.
-The lower part shows the frequency with which each of the 15 actions is selected across 50 independent trials.
-The frequencies for actions indexed 1 through 13 are relatively uniform across algorithms, but a notable distinction is that Algorithm 1 (green) exhibits a strong preference toward the boundary actions—those indexed 0 and 14.
-To understand this preference of Algorithm 1, recall that its output frequencies directly reflect the underlying measure $q$, since the algorithm samples actions according to $q$ by design.
-In Figure 4, we vary the RBF kernel length-scale across values 0.1, 1, and 4.
-For each setting, we sample ten example outcome functions (shown in the upper part) and run Algorithm 1 for 50 trials, recording the frequency of each selected action as an approximation of the $q$-measure (shown in the lower part).
-In Figure 4, when the length-scale is 4, the outcome functions are very smooth—almost linear—so the maximum is typically attained at the boundary actions (indices 0 and 14), causing the $q$-measure to concentrate heavily on them.
-In contrast, with a length-scale of 0.1, the functions are very rough, with outcomes nearly independent across actions, resulting in an approximately uniform $q$-measure over the action space.
-Further, since the outcome functions are generated using a stationary RBF kernel, the corresponding feature representations $\Phi(a)$, for $a\in\mathcal{A}_{full}$, are uniformly distributed.
-When action outcomes are nearly independent (e.g., with a very small length-scale), the feature vectors $\Phi(a)$ form an orthogonal basis, with each pair $(\Phi(a),\Phi(a^{\prime}))$ equidistant.
-As discussed in Supplementary I, this leads to a uniform $q$-measure when each cluster contains a single action.
-Indeed, the lower-left part of Figure 4 confirms this: for length-scale $l=0.1$, the action selection frequencies are nearly uniform.
-As the length-scale increases, action outcomes become more correlated, and the pairwise distances between $\Phi(a)$ vectors shrink.
-The boundary actions (indices 0 and 14), remain the farthest apart and thus are more likely to be optimal in a given bandit instance.
-This explains why Algorithm 1 performs better: it closely aligned with the $q$-measure, whereas UCB and TS select actions more uniformly.
-Further, in any given round $\theta$ of TS/UCB, the reward $\max_{a\in\mathcal{A}}\mu_{a}$ for optimal and suboptimal super arms can be indistinguishable if both include the optimal base action $a^{*}(\theta)$.
-This ambiguity slows down the convergence of the learning algorithms because they cannot effectively differentiate between good and bad super arms—a problem exacerbated by the combinatorial number of possible super arms.
-Figure 3: Comparison of representative actions selected by three algorithms in a synthetic setting.
-Outcome functions are generated using an RBF kernel (length-scale $l=1$) over 15 grid points in $[0,2]$.
-In each trial, $K=5$ distinct actions are selected per algorithm (Algorithm 1 in green, TS in yellow, and UCB in red); UCB and TS are run for 3000 rounds.
-Upper: Representative actions selected in three individual trials; each row shows one trial.
-Lower: Action selection frequencies of Algorithm 1, TS, and UCB over 50 trials.
-Algorithm 1 consistently selects boundary actions (indices 0 and 14), while UCB and TS select more uniformly.
-Figure 4: Effect of kernel length-scale on importance measure $q$.
-Upper: Example outcome functions generated using RBF kernels with varying length-scales $l\in\{0.1,1,4\}$, over 15 grid points in $[0,2]$, with actions indexed from 0 to 14.
-Lower: Action selection frequencies of Algorithm 1 over 50 trials for each length-scale setting.
-These frequencies approximate the importance measure $q$.
-As the length-scale increases, the outcome functions become smoother and more linear, causing $q$-measure to concentrate at the boundary actions (indices 0 and 14).
-Algorithm 2 Upper Confidence Bound with Gaussian Rewards
-1:Input: Base actions $\mathcal{A}_{full}$, subset cardinality $K$, kernel matrix $\mathbf{K}$, time horizon $T=3000$
-2:Initialize: Super arms $\{\mathcal{A}_{i}\}_{i\leq M}$ with $M=\binom{|\mathcal{A}_{full}|}{K}$, reward variance $\sigma^{2}=1$
-3:Initialize: $N_{i}\leftarrow 0$, $\hat{\mu}_{i}\leftarrow 0$ for all $i=1,\dots,K$
-4:for $t=1$ to $T+M$ do $\triangleright$ Each round corresponds to a bandit
-5:     Sample an outcome function $f\sim\mathcal{N}(0,\mathbf{K})$
-6:     if $t\leq M$ then
-7:         Select arm $i_{t}=t$
-8:     else
-9:         Select arm $i_{t}=\arg\max_{i}\left(\hat{\mu}_{i}+\sqrt{\frac{2\log t}{N_{i}+\epsilon}}\right)$
-10:     end if
-11:     Pull arm $i_{t}$, observe reward $r_{t}=\max_{a\in\mathcal{A}_{i_{t}}}f(a)$
-12:     Update:
-$N_{i_{t}}\leftarrow N_{i_{t}}+1$
-$\hat{\mu}_{i_{t}}\leftarrow\hat{\mu}_{i_{t}}+\frac{r_{t}-\hat{\mu}_{i_{t}}}{N_{i_{t}}}$
-13:end for
-Algorithm 3 Thompson Sampling with Gaussian Priors
-1:Input: Base actions $\mathcal{A}_{full}$, subset cardinality $K$, kernel matrix $\mathbf{K}$, time horizon $T=3000$
-2:Initialize: Super arms $\{\mathcal{A}_{i}\}_{i\leq M}$ with $M=\binom{|\mathcal{A}_{full}|}{K}$, reward variance $\sigma^{2}=1$
-3:Initialize: prior mean $\mu_{i}\leftarrow 0$, prior precision $\tau_{i}\leftarrow\epsilon$ for all $i=1,\dots,K$
-4:for $t=1$ to $T$ do $\triangleright$ Each round corresponds to a bandit
-5:     Sample an outcome function $f\sim\mathcal{N}(0,\mathbf{K})$
-6:     for $i=1$ to $M$ do
-7:         Sample $\tilde{\mu}_{i}\sim\mathcal{N}(\mu_{i},\sigma^{2}/\tau_{i})$
-8:     end for
-9:     Select arm $i_{t}=\arg\max_{i}\tilde{\mu}_{i}$
-10:     Pull arm $i_{t}$, observe reward $r_{t}=\max_{a\in\mathcal{A}_{i_{t}}}f(a)$
-11:     Update:
-$\tau_{i_{t}}\leftarrow\tau_{i_{t}}+1$
-$\mu_{i_{t}}\leftarrow\mu_{i_{t}}+\frac{r_{t}-\mu_{i_{t}}}{\tau_{i_{t}}}$
-12:end for
+As a result, any algorithm, including Algorithm 1, would perform poorly unless the subset size $|\mathcal{A}|$ is sufficiently large.
 Appendix K Properties of Gaussian width
@@
 where the first inequality uses Jensen’s inequality. Taking $\delta:=\sqrt{2(\log N)/\sigma^{2}}$ yields the results. ∎
-Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [28]:
+Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [Kamath, 2015]:
 $\mathbb{E}_{\theta}\left[\max_{i=1,\dots,N}X_{i}\right]\geq\frac{\sigma\sqrt{\log N}}{\sqrt{\pi\log 2}}.$

### v3

@@
 By modeling the expected outcome of each drug for each patient as a Gaussian process, we can capture these correlations.
-Now, consider medicine demand: If a drug treats a very rare illness, it is unlikely to be needed frequently, so the store can exclude it to optimize storage space.
-Conversely, if flu season is approaching, stocking several flu medications is a wise choice.
-Different from prior approaches in multi-armed bandits (MAB) (Dani et al., 2008) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits.
+This correlation structure makes subset selection particularly interesting: if two drugs treat the same disease, choosing one may suffice, while drugs for distinct diseases might both be kept.
+In addition, demand also matters: drugs for rare illnesses can be excluded to save space, while flu medications should be stocked during flu season.
+Different from prior approaches in multi-armed bandits (MAB) (lai1985asymptotically, ) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits.
 This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
-Consider the following MAB setting: In a bandit, if a decision-maker plays an action with a fixed but unknown feature vector $a\in\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$.
+Consider the following setting: In a bandit, if a decision-maker plays an action with a fixed but unknown feature vector $a\in\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$.
 We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle$ where this bandit instance $\theta\in\mathbb{R}^{n}$ is drawn from an unknown multivariate Gaussian distribution.
-Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ forms a Gaussian process (GP) (Vershynin, 2018, Chapter 7), and we also provide results under general sub-Gaussian assumption.
-Our underlying setting aligns with the one considered in contextual bandit (Dani et al., 2008); however, the key difference lies in the objective.
-Now consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$.
-For a given bandit instance $\theta$, if the best action over the full space lies within $\mathcal{A}$, the decision-maker benefits from reduced suboptimal actions to explore.
-Conversely, if the best action lies outside $\mathcal{A}$, regret arises from being unable to select this best action.
-Thus, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
-$\boxed{\textnormal{Regret}:=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}},}$ (1)
+Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ forms a Gaussian process (GP) (vershynin2018high, , Chapter 7), while a more general sub-Gaussian assumption is considered in the Supplementary.
+Our setting aligns with that of contextual bandits (dani2008stochastic, ), and the restriction to linear functions may be partially mitigated by allowing $n$ to be unbounded.
+Consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$.
+For a given bandit instance $\theta$, if the optimal action over the full space lies within $\mathcal{A}$, the decision-maker benefits from reduced suboptimal actions to explore.
+Conversely, if the optimal action lies outside $\mathcal{A}$, regret arises from being unable to select this best action.
+Thus, for a bandit instance $\theta$, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
+$\boxed{\textnormal{Regret}(\theta):=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}}(\theta),}$ (1)
 which depends on the sampled bandit instance and is therefore a random variable.
@@
 This objective is motivated by practical considerations.
-The classic Bayesian regret in the bandit literature (Agrawal and Goyal, 2012) typically scales with the number of available actions.
-However, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower due to the reduced number of suboptimal actions.
+The classic Bayesian regret in the bandit literature (agrawal2012analysis, ) typically scales with the number of available actions.
+However, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower.
 This is especially beneficial when the action space is large, as even the initialization phase can be computationally expensive.
 To see this, we can decompose the bandit regret as follows:
-[…]
+$\textnormal{BayesianBanditRegret}:=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\mu_{A_{t}}(\theta)\right]$
 $=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}}\mu_{a}(\theta)\!-\!\mu_{A_{t}}+\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)\!-\!\max_{a\in\mathcal{A}}\mu_{a}(\theta)\right]$
 $\leq C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}].$
-In the first equality, $A_{t}\in\mathcal{A}$ denotes the action chosen by a policy, e.g., Thompson Sampling (Agrawal and Goyal, 2012), in round $t$, and $N$ is the number of rounds.
+In the first equality, $A_{t}\in\mathcal{A}$ denotes the action chosen by a policy, e.g., Thompson Sampling (agrawal2012analysis, ), in round $t$, and $N$ is the number of rounds.
 The expectation is taken over the randomness in the distribution of bandit instances and in actions selected by the policy.
-The inequality follows from the well-known regret bounds for Thompson Sampling (Lattimore and Szepesvári, 2020).
-$|\mathcal{A}|$ denotes the cardinality of the action set and $C>0$ is a constant.
+The inequality follows from the well-known regret bounds for Thompson Sampling (lattimore2020bandit, ) [^1] This bound could be relaxed if the feature vector $a$ is known, or if $a$ can be incorporated as the input of a kernel function, as discussed in related works.
+However, we use this general bound because we also consider the most general case where $\mathcal{A}_{\operatorname*{full}}$ is simply an index set, e.g., $\mathcal{A}_{\operatorname*{full}}:=\{\textnormal{treatment A},\textnormal{treatment B},\textnormal{treatment C}\}$., where $|\mathcal{A}|$ denotes the cardinality of the action set and $C>0$ is a constant.
 Note that if the policy has access to the full action space, the Bayesian bandit regret is instead bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$.
@@
 1.1 Related Works
-Multi-Armed Bandits (Lai and Robbins, 1985; Auer et al., 2002a) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution.
+Multi-Armed Bandits (lai1985asymptotically, ; auer2002finite, ) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution.
 The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time.
-Popular methods include the UCB (Auer et al., 2002a), TS (Agrawal and Goyal, 2012), and EXP3 (Auer et al., 2002b) for adversarial settings.
-Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few independent samples as possible (Jamieson and Nowak, 2014; Kaufmann et al., 2016).
-Popular methods in fixed confidence setting include Action Elimination Even-Dar et al.
-(2006); Karnin et al.
-(2013), UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum.
-In fixed budget setting, there is Successive halving Karnin et al.
-(2013), successive reject Audibert and Bubeck (2010).
-Our algorithm includes a step for identifying the optimal action.
-While this line of work assumes (non-centered) sub-Gaussian outcome distributions within a bandit instance, we model each instance with a linear structure, allowing for further simplification.
-Stochastic Linear Optimization assumes that the expected outcome of each action depends through the inner product between a context $\theta$ and an action $a\in\mathcal{A}_{\operatorname*{full}}$ Auer (2002); Dani et al.
-(2008); Rusmevichientong and Tsitsiklis (2010), and our model of each bandit instance align with this setting.
-This is a sequential method for maximizing a linear function—i.e., $f(a):=\mu_{a}(\theta)$ with fixed $\theta$—based on noisy observations of the function values.
-Soare et al.
-(2014) study the sample complexity of optimal action identification in this setting.
-However, this line of work assumes the action feature vectors are known.
-GP Optimization can deal with the case that the feature vectors of actions are unknown.
-Srinivas et al.
-(2009) models the outcome function as a sample from a GP prior with a kernel function (Williams and Rasmussen, 2006).
-Their Bayesian algorithm, GP-UCB, uses this GP prior to first select points $x_{t}$ that enable global estimation of the function, and then plays the point with the highest posterior mean.
-The widespread adoption of this method in bandit settings (Valko et al., 2013; Li and Scarlett, 2022), as well as in continuous action spaces (Chowdhury and Gopalan, 2017), highlights the practicality of assuming that action outcomes are correlated.
+Popular methods include the UCB (auer2002finite, ), TS (agrawal2012analysis, ), and EXP3 (auer2002nonstochastic, ) for adversarial settings.
+Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few samples as possible (jamieson2014best, ; kaufmann2016complexity, ).
+Popular methods in fixed confidence setting include Action Elimination even2006action ; karnin2013almost , UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum.
+In fixed budget setting, there is Successive halving karnin2013almost , successive reject audibert2010best .
+Stochastic Linear Optimization assumes that the expected outcome of each action depends through the inner product between a context $\theta$ and an action $a\in\mathcal{A}_{\operatorname*{full}}$ auer2002using ; dani2008stochastic ; rusmevichientong2010linearly .
+This line of work assumes that the action feature vectors are known, so the cardinality of the action space does not play a role and may even be infinite.
+soare2014best study the sample complexity of optimal action identification in this setting, which scales linearly with the action space dimension $n$.
+GP Optimization addresses the case where the feature vectors of actions are unknown but a kernel function is available, so the inputs to the kernel are required.
+srinivas2009gaussian model the outcome function as a sample from a GP prior with a kernel function (williams2006gaussian, ), achieving bandit regret that scales as $\sqrt{\log|\mathcal{A}_{\operatorname*{full}}|}$ for finite action space.
+The widespread adoption of this method in bandit settings (valko2013finite, ; li2022gaussian, ), as well as in continuous action spaces (chowdhury2017kernelized, ), highlights the practicality of assuming that action outcomes are correlated.
 This also motivates our extension, where the outcome function of each bandit instance is modeled as a sample from a RKHS.
-Before discussing two seemingly relevant lines of work that select a subset from the action space, we first highlight a advantage of our approach.
-These methods assume a fixed $|\mathcal{A}|=K$, which is often unclear in practice and requires restarting the algorithm when changed.
+Before discussing two seemingly relevant lines of work that select a subset from the action space, we first highlight an advantage of our approach.
+These methods assume a fixed subset cardinality $|\mathcal{A}|=K$, which is often unclear in practice and requires restarting the algorithm when changed.
 In contrast, our algorithm can adapt the subset size on the fly.
-For example, if 100 runs all return the same action, choosing this one may suffice ($|\mathcal{A}|=1$); if each run yields a different action, we may need more iterations to ensure a small regret ($|\mathcal{A}|>100$).
-[…]
+Top-K Action Identification aims to identify the $K$ actions with the highest expected outcomes using as few samples as possible (kalyanakrishnan2012pac, ; gabillon2012best, ; kaufmann2016complexity, ; chen2017nearly, ).
+This line of work assumes that all actions are independent and have distinct expected values, making its methods inapplicable to our framework.
+If one were to apply these methods regardless, the most reasonable approach, in our view, would be to treat the family of bandit instances as a super-bandit, where each bandit instance corresponds to a round, and the expected payoff of an action in that round is given by $\mu_{a}$.
+In this setting, top-$K$ identification would refer to selecting $K$ actions with the highest expected payoffs $\mathbb{E}[\mu_{a}]$.
+In contrast, our framework considers that the expected outcome of each action, averaged over the distribution of bandits, may be the same—i.e., $\mathbb{E}[\mu_{a}]\!=\!c$ for all $a\!\in\!\mathcal{A}_{\operatorname*{full}}$, where $c$ is a constant—so that the entire action space shares the same highest expected payoff.
+Further, even if $\mathbb{E}[\mu_{a}]$ varies across actions, ignoring correlations can be fatal in our framework:
 Example 1.
@@
 $\mathbb{E}\langle a_{1},\theta\rangle=\mathbb{E}\langle a_{2},\theta\rangle=0.5,\quad\mathbb{E}\langle a_{3},\theta\rangle=0.45.$
-So under the Best-222-Action perspective, $a_{1}$ and $a_{2}$ would be selected. However, this is suboptimal in our framework, since $\mu_{a_{1}}$ and $\mu_{a_{2}}$ are positively correlated:
+So under the Best-$2$-Action perspective, $a_{1}$ and $a_{2}$ would be selected. However, this is suboptimal in our framework, since $\mu_{a_{1}}$ and $\mu_{a_{2}}$ are positively correlated:
 $\mathbb{E}\max_{a\in\{a_{1},a_{2}\}}\langle a,\theta\rangle=0.55,\quad\mathbb{E}\max_{a\in\{a_{1},a_{3}\}}\langle a,\theta\rangle=1.$
-In contrast, our algorithm, if run until it selects two distinct actions, would output $a_{1}$ and $a_{3}$, which is the true optimal.
+Our algorithm, if run until it selects two distinct actions, would output $a_{1}$ and $a_{3}$—the true optimal.
 Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super arm $\mathcal{A}$, with $|\mathcal{A}|=K$.
-Popular methods include CUCB Chen et al.
-(2016), CTS Wang and Chen (2018).
-We argue that this line of work is not applicable to our framework: (1) It assumes that the expected outcome of a super arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition.
+Popular methods include CUCB chen2016combinatorial , CTS wang2018thompson .
+We argue that this line of work is not directly applicable to our framework, but we include it as a baseline in subsequent empirical evaluations: (1) It assumes that the expected outcome of a super arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition.
 In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions.
@@
 Ignoring these correlations misses the core challenge—an issue illustrated in Example 1.
-[…]
-Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$.
-It is an important topic in high-dimensional probability (Vershynin, 2018).
-The sharpest known bounds are due to Talagrand (2014).
 2 Subset Selection Framework
@@
 (This framework applies to the case $n=+\infty$, with the additional assumption $\sum_{i\geq 1}a^{2}_{i}\!<\!\infty$.) The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, and includes a constant $c\in\mathbb{R}$:
-$\mu_{a}:=\langle a,g\rangle+c,\quad\forall,a\in\mathcal{A}_{\operatorname*{full}},;g\sim\mathcal{N}(0,\Sigma),$ (2)
+$\mu_{a}:=\langle a,g\rangle+c,\quad\forall\;a\in\mathcal{A}_{\operatorname*{full}},\;g\sim\mathcal{N}(0,\Sigma),$ (2)
 where $\Sigma$ is a positive semi-definite matrix.
-Let $\theta\in\mathbb{R}^{n}$ follows a multivariate normal distribution with zero mean and identity covariance matrix.
+Let $\theta\in\mathbb{R}^{n}$ follows a multivariate normal distribution with zero mean and identity covariance matrix $I$.
 The distribution of $g$ is in fact equivalent to $\Sigma^{1/2}\theta$.
@@
 $\langle a,g\rangle=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
-Therefore, the setting in Equation (2) is equivalent to
-$\mu_{a}:=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}},\;\theta\sim\mathcal{N}(0,I),$
-where $\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}}$ denotes the image of $\mathcal{A}_{\operatorname*{full}}$ under the linear transformation $\Sigma^{1/2}$.
-Since the constant $c$ does not affect the regret (as defined in Equation (1)), we can, without loss of generality, focus on this canonical Gaussian process (Vershynin, 2018, Chapter 7) in the remainder:
+Therefore, the setting in equation 2 is equivalent to $\mu_{a}:=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}},\;\theta\sim\mathcal{N}(0,I),$ where $\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}}$ denotes the image of $\mathcal{A}_{\operatorname*{full}}$ under the linear transformation $\Sigma^{1/2}$.
+Since the constant $c$ does not affect the regret (as defined in equation 1), we can, without loss of generality, focus on this canonical Gaussian process (vershynin2018high, , Chapter 7) in the remainder:
 $\boxed{\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{\operatorname*{full}},\theta\sim\mathcal{N}(0,I).}$ (3)
@@
 $\begin{split}\langle a,\theta\rangle=\sum_{i=1}^{K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i=1}^{K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.\end{split}$
-By Equation (3), it yields
-$\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$
-where the left inequality uses $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$.
+By equation 3, it yields $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$ where the left inequality uses $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$.
 Thus, the two quantities $\max_{a\in\mathcal{A}}\mu_{a}$ and $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ are equal.
@@
 2.1 Epsilon Nets
-If the set of extreme points $\mathcal{A}_{\operatorname*{full}}$ is still large, a natural approach is to construct a grid over the action space, where the grid points serve as representative actions.
+If $\mathcal{A}_{\operatorname*{full}}$ is very large, a natural approach is to construct a grid over the action space, where the grid points serve as representative actions.
 This ensures that for every action in the full space, there exists a representative that is close to it.
 This idea is formally captured by the notion of a (geometric) $\epsilon$-net.
-To proceed, we clarify what we mean by an $\epsilon$-net, as there are at least two definitions: one from a geometric perspective (Vershynin, 2018, Chapter 4) and another from a measure-theoretic perspective (Matousek, 2013, Chapter 10).
+To proceed, we clarify what we mean by an $\epsilon$-net, as there are at least two definitions: one from a geometric perspective (vershynin2018high, , Chapter 4) and another from a measure-theoretic perspective (matousek2013lectures, , Chapter 10).
 Let $\|\cdot\|_{2}$ denote the Euclidean norm.
@@
     $r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
-A geometric $\epsilon$-net ensures small regret because if two actions $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ are close in the Euclidean sense, then the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., their expected squared difference is small):
-$\|\mu_{a}\!-\!\mu_{a^{\prime}}\|_{L^{2}}=\!\left(\mathbb{E}(a\!-\!a^{\prime})^{\top}\theta\theta^{\top}(a\!-\!a^{\prime})\right)^{1/2}\!=\|a\!-\!a^{\prime}\|_{2},$
-where $\theta^{\top}$ denotes the transpose of $\theta$.
-The equalities use Equation (3) and $\mathbb{E}\theta\theta^{\top}=I$.
+A geometric $\epsilon$-net ensures small regret because if two actions $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ are close in the Euclidean sense, then the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., their expected squared difference is small): $\|\mu_{a}\!-\!\mu_{a^{\prime}}\|_{L^{2}}=\!\left(\mathbb{E}(a\!-\!a^{\prime})^{\top}\theta\theta^{\top}(a\!-\!a^{\prime})\right)^{1/2}\!=\|a\!-\!a^{\prime}\|_{2},$ where $\theta^{\top}$ denotes the transpose of $\theta$.
+The equalities use equation 3 and $\mathbb{E}\theta\theta^{\top}=I$.
 Therefore, by definition, a geometric $\epsilon$-net guarantees the existence of an action $a\in\mathcal{A}$ whose expected outcome $\mu_{a}$ is close to that of the optimal action for any given bandit instance.
-However, this net suffers from the curse of dimension: e.g., for $[0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$.
+However, this net suffers from the curse of dimension: e.g., for $0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$.
 The measure-theoretic $\epsilon$-net addresses this issue.
@@
 2.2 Epsilon Net Algorithm
-We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by Haussler and Welzl (1986).
+We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by haussler1986epsilon .
 It selects $K$ i.i.d.
@@
 6:  Find optimal action $a^{*}(\theta)\!:=\!\operatorname*{arg\,max}_{a\in\mathcal{A}_{\operatorname*{full}}}\langle a,\theta\rangle$.
-7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$triangleright Repetition of actions is allowed
+7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$ $\triangleright$ Repetition of actions is allowed
 8:end for
@@
 Assumption 2 (Unique optimal action).
-The optimal action $a^{*}(\theta)$ is unique with probability 111 for all bandit instances.
+The optimal action $a^{*}(\theta)$ is unique with probability $1$ over all bandit instances.
 Define the Importance Measure $q$ over a partition $\mathcal{R}$:
 $\boxed{q(r):=\Pr[a^{*}(\theta)\in r]=\!\int\!\mathbb{1}\{a^{*}(\theta)\in r\}\;p(\theta)d\theta,}$ (4)
-[…]
+where $p(\theta)$ is the density of $\theta$ and $\int p(\theta)d\theta=1$.
+Under Assumption 2, measure $q$ is a probability distribution.
+It reflects the probability that a given cluster contains the optimal action and thus represents the potential contribution of that cluster to the expected regret.
 Assumption 3.
@@
 In fact, with high probability, this algorithm outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$ with respect to measure $q$.
-(This is a simplified version of Theorem 10.2.4 of Matousek (2013).)
+(This is a simplified version of Theorem 10.2.4 of matousek2013lectures .)
 Lemma 4.
@@
 Geometrically, $\epsilon$ is the largest cluster diameter $\epsilon:=\max_{r\in\mathcal{R}}\mathrm{diam}(r)$; measure-theoretically, $\epsilon$ is the smallest cluster measure $\epsilon:=\min_{r\in\mathcal{R}}q(r)$.
-Hence, we have the following corollary:
-Corollary 5.
-Under the same conditions as Lemma 4, let $\epsilon:=\min_{r\in\mathcal{R}}q(r)$. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, the following holds:
-$\forall a\in\mathcal{A}_{\operatorname*{full}},\ \exists a^{\prime}\in\mathcal{A}\text{ such that }\|a-a^{\prime}\|_{2}<\max_{r\in\mathcal{R}}\operatorname{diam}(r).$
 3 Regret Analysis
@@
 Definition 1 (Reference subsets).
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
-A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$, and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
-We impose the following assumption on the partition, used only for lower bounds: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it.
-This ensures clusters are well-separated.
-Assumption 6 (Effective partition).
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
+A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$ and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
+Only for lower bounds, we assume a well-separated partition: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it.
+Assumption 5.
 For any $r\in\mathcal{R}$, whenever the optimal action lies in cluster $r$, i.e., $a^{*}(\theta)\in r$, then $\mu_{a}\geq\max_{a^{\prime}\in\mathcal{A}_{\operatorname*{full}}\setminus r}\mu_{a^{\prime}},\;\forall a\in r$.
 Theorem 6 (Regret bounds of reference subsets).
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$.
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$.
 Then, there is an absolute constant $C>0$, such that
@@
 $\square$
+For each $\ell\leq m$, the term $\mathbb{E}_{\theta}[\max_{a\in r_{\ell}}\mu_{a}]$ equals $\mathbb{E}_{\theta}[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}]$, since $\mathbb{E}_{\theta}[\mu_{a}]=0$.
 3.1 Regret Bounds of Algorithm
@@
 The key difference is that whereas the reference subset includes a representative from each cluster, the algorithm may miss some clusters.
-However, although it may select a smaller subset, the algorithm achieves regret comparable to that of the reference subset, as it tends to miss clusters that contribute minimally to the expected regret.
+However, the algorithm still achieves regret comparable to that of the reference subset, as it tends to miss clusters that contribute minimally to the expected regret.
 The expected regret in previous results is taken over bandit instances $\theta$.
 In contrast, since the output of Algorithm 1 is random, the expected regret analyzed in this section is taken with respect to both the algorithm’s randomness (i.e., the sampled $\mathcal{A}$) and the distribution over $\theta$.
-Upper Bound.
 Theorem 7.
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
 Let $\mathcal{A}$ be output of Algorithm 1.
 For the same constant $C>0$ in Theorem 6,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
 Proof sketch.
@@
 The new set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset.
-Then, for each $\ell\leq m$, define a simple Gaussian process $\{Z_{a}\}_{a\in r_{\ell}}$, where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and let $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$.
-When $a^{*}(\theta)\in r_{\ell}$, we consider two cases:
-  • $r_{\ell}\cap\mathcal{A}\neq\emptyset$: The regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
-  • $r_{\ell}\cap\mathcal{A}=\emptyset$: The regret is bounded by $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\!\mu_{a}$. $\square$
-The term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ in Theorem 8 behaves similarly to entropy: it is maximized when $q$ is uniform, provided $|\mathcal{R}|\geq 2K+1$ (since we can choose any partition); see Lemma 18 in Supplementary.
+Then, for each $\ell\leq m$, define a Gaussian process $\{Z_{a}\}_{a\in r_{\ell}}$, where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and let $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$.
+When $a^{*}(\theta)\in r_{\ell}$, we consider two cases.
+If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, the regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
+If $r_{\ell}\cap\mathcal{A}=\emptyset$, the regret is bounded by $\max_{a\in\mathcal{A}_{\text{full}}}\mu_{a}$.
+$\square$
+The term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ in Theorem 7 is maximized by the uniform $q$, provided $|\mathcal{R}|\geq 2K+1$ (since we can choose any partition); see Supplementary.
 Intuitively, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$.
 This expression is small when $q(r)$ is small, and decays faster when $q(r)$ gets larger, making the overall term negligible if $q$ is highly concentrated.
-The connection between Theorem 7 and 8 is insightful:
+The connection between Theorem 6 and 7 is insightful:
 Remark 8.
@@
 In this case, the benefit of constructing an explicit grid—often a non-trivial task when the action feature vectors are unknown or high-dimensional—is limited.
-On the other hand, if the measure $q$ is close to uniform, then $\min_{r\in\mathcal{R}}q(r)$ is large.
-In this case, with relatively high probability, Algorithm 1 outputs a geometric $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$-net (see Corollary 5).
-Moreover, since the partition is constructed using a grid, the cluster diameters are similar, so $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$ closely matches the spacing of the original grid—making the output resemble a slightly coarser grid.
-Partition-Independent Upper Bound.
 Generally speaking, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified.
 We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$:
-$N(\mathcal{A}_{\operatorname*{full}},\!\epsilon)=\min\Big{\{}m\!\in\!\mathbb{N}:$$\;\exists\{a_{\ell}\}_{\ell\leq m}\subseteq\mathcal{A}_{\operatorname*{full}},$
-$\forall a\in\mathcal{A}_{\operatorname*{full}},\exists\ell,\|a\!-\!a_{\ell}\|_{2}\leq\epsilon\Big{\}}.$
+$N(\mathcal{A}_{\operatorname*{full}},\!\epsilon)=\min\Big\{m\!\in\!\mathbb{N}:\;\exists\{a_{\ell}\}_{\ell\leq m}\subseteq\mathcal{A}_{\operatorname*{full}},\forall a\in\mathcal{A}_{\operatorname*{full}},\exists\ell,\|a\!-\!a_{\ell}\|_{2}\leq\epsilon\Big\}.$
 Theorem 9.
@@
 For the same constant $C>0$ in Theorem 6, and another absolute constant $c>0$,
-$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;$ $2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},$
-$\text{where }K\geq$ $c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon).$
+$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\text{where }K\geq$ $c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon).$
 As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
-[…]
-Lower Bound.
+The partition-independent upper bound in Theorem 9 depends on a positive constant $\epsilon$, which serves as a tolerance parameter up to the diameter of the action space.
+Given a $\epsilon$, the required number of samples $K$ scales with the square of the diameter-to-$\epsilon$ ratio, multiplied by the covering number.
+The resulting expected regret is then bounded in terms of $\epsilon$, the dimensionality, and the logarithm of the covering number.
+As $\epsilon$ decreases, more samples are needed, but the regret bound becomes tighter.
+The matching lower bound for Theorem 7, as well as the upper bound under the general sub-Gaussian process assumption for Algorithm 1, are provided in the Supplementary.
+4 Generalization and Empirical Validation
+In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings.
+Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling.
+We therefore treat the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set, e.g., $\mathcal{A}_{\operatorname*{full}}:=\{\textnormal{treatment A},\textnormal{treatment B},\textnormal{treatment C}\}$.
+In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$: $f(a):=\mu_{a},\forall a\in\mathcal{A}_{\operatorname*{full}}$.
+We show that our framework applies when the outcome functions lie in a reproducing kernel Hilbert space (RKHS) (williams2006gaussian, ).
+Consider a positive semidefinite kernel $k:\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}\to\mathbb{R}$.
+By Mercer’s theorem, for a non-negative measure $\mathbb{P}$ over $\mathcal{A}_{\operatorname*{full}}$, if the kernel satisfies $\int_{\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}}k^{2}(a,a^{\prime})d\mathbb{P}(a)d\mathbb{P}(a^{\prime})<\infty,$ then it admits an eigenfunction expansion: $k(a,a^{\prime})=\sum\nolimits_{i\leq\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(a^{\prime}),$ where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues.
+Let outcome functionff be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{\operatorname*{full}}$ and a weight vector $\alpha\in\mathbb{R}^{N}$.
+The function $f$ can be rewritten as
+$\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
+where $\mathbf{f}$ is a vector of coefficients, and $\Phi(a)$ usually refereed as a feature map, has entries $\Phi_{i}(a)=\sqrt{\lambda_{i}}\phi_{i}(a)$.
+The full action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{\operatorname*{full}}$.
+Assuming the outcome function is generated from a kernel inherently implies that actions are correlated.
+To study the effect of varying correlation levels among actions, we use the parametric RBF kernel to sample outcome functions from a RKHS: $k(a,a^{\prime}):=\exp\left(-\frac{\|a-a^{\prime}\|^{2}}{2l^{2}}\right)$ , where $l$ is a length-scale parameter that control the dependence among actions.
+By (kanagawa2018gaussian, , Theorem 4.12), we sample functions from a RKHS by first constructing the kernel matrix $\mathbf{K}$ with entries $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, and then drawing $f\sim\mathcal{N}(0,\mathbf{K})$.
+Note that although the kernel is used to generate outcome functions, the kernel itself is not revealed; consequently, $a$ and $a^{\prime}$ are regarded merely as indices rather than as inputs to the kernel.
+Figure 1: Comparison on solving equation 5, selecting $K=5$ distinct actions from 15 grid points over $[0,2]$, with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales.
+Left: Expected regret over 50 repetitions of our method with exhaustive search (green), compared with Thompson Sampling (TS; yellow), Upper Confidence Interval (UCB; orange), and Successive Halving (SH, blue) applied to the set of super-arms.
+TS/UCB are run for 3000 rounds; SH is given a budget of 37,000 pulls.
+Right: SH’s expected regret (blue) and pull counts (yellow bars) per round, compared with our method (green).
+SH requires nearly 10,000 pulls—about three times the number of super-arms—to match our method.
+Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in equation 1.
+Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
+$\max_{\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (5)
+In general, the problem is non-trivial since the objective involves an expectation over random variables with an unspecified distribution, and the inner expression is a maximum over a subset of potentially correlated variables.
+Consequently, small changes in $\mathcal{A}$ can cause non-smooth variations in the maximum.
+Our method is computationally efficient, avoiding the combinatorial complexity.
+We assess performance in identifying near-optimal solutions by selecting $K=5$ actions from an action space of 15 grid points in $[0,2]$.
+The configuration yields a moderate number of super arms, allowing approximation of the optimal subset using Successive Halving (SH), which, however, becomes impractical for larger action spaces.
+Outcome functions are generated with an RBF kernel at length-scales $l\in\{0.5,1,\dots,4\}$.
+We compare the proposed method against Thompson Sampling (TS) and Upper Confidence Bound (UCB).
+In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where exhaustive search is used over the action space to find the optimal action $a^{*}(\theta)$.
+For other methods, we treat each $K$-tuple of actions as a super arm, and the payoff of each super arm in a round $\mathcal{A}$ is given by $\max_{a\in\mathcal{A}}\mu_{a}$.
+Thus, these methods are tailored to find the best super arm that maximizes $\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]$, aligning with the same objective.
+In TS/UCB methods, we adopt a bandit feedback setting: at each round (corresponding to a bandit instance $\theta$), the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}(\theta)$, and updates its policy accordingly, repeating this process for 3000 rounds, chosen to roughly match the number of super arms ($N=3003$).
+Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB, with the prior set to $\mathcal{N}(0,1)$.
+In the SH method, all super arms are evaluated using a fixed budget of arm pulls (we use the minimum budget required in karnin2013almost , $N\log_{2}N\approx 37{,}000$) over a few rounds.
+In each round (corresponding to a bandit instance $\theta$), the number of remaining super arms is halved.
+This process continues until only the best-performing super arm remains or the budget is exhausted.
+We evaluate expected regret of all methods over $10^{5}$ additional randomly sampled outcome functions.
+The left subplot of Figure 1 reports the expected regret of our method (green), TS (yellow), UCB (orange) and SH (blue), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions.
+As shown, the expected regret decreases as the length-scale increases.
+The length-scale $l$ controls the effective number of approximately independent actions: when $a$ and $a^{\prime}$ are far relative to $l$, $k(a,a^{\prime})$ is negligible, so $\mu_{a}$ and $\mu_{a^{\prime}}$ are nearly uncorrelated.
+In the limit, $\{\mu_{a}\}_{a\in\mathcal{A}_{\text{full}}}$ forms a collection of i.i.d.
+zero-mean, unit-variance variables.
+The SH in left subplot of Figure 1 is used to approximate the ground-truth optimal subset and our method does give similar performance as SH.
+The right subplot highlights SH’s impracticality: we ran our method (green) and SH (blue) for 50 repetitions with fixed $l=1$, tracking the current best super arm and total super arm pulls after each SH round.
+The blue curves and shades represent the mean $\pm$ one standard deviation of SH’ expected regret, while yellow bars (corresponding to the right y-axis) show the number of super arm pulls by SH at the end of each round.
+Green curves show our method’s expected regret.
+SH matches our performance after three rounds but requires 9005 pulls—roughly three times the number of super arms—making it intractable for large $\mathcal{A}_{\operatorname*{full}}$.
+Figure 2: Comparison of our method (green), CTS (brown), and CUCB (red) on solving equation 5 with $K=10$ actions chosen from 500 grid points over [5,5][-5,5], with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales.
+In our method, EpsilonNet+TS, the “argmax” step in Algorithm 1 is replaced by TS run for 300 rounds.
+CTS is run for 3000 rounds and CUCB for 2000 rounds.
+All methods use prior variance 11.
+Left: expected regret over 30 repetitions.
+Right: runtime (purple, left axis) and number of arm pulls (gray, right axis) over 30 repetitions.
+Further, we demonstrate the generality of our method (denoted EpsilonNet+TS), where the “argmax” step in Algorithm 1 is replaced by TS.
+We consider the same setup as in Figure 1, but with $K=10$ actions chosen from 500 grid points over [5,5][-5,5].
+In our method, for each bandit instance $\theta$, we run TS for 300 rounds, selecting an action $a$ and observing $\mu_{a}$ in each round; the final action is taken as an approximation of the optimal action $a^{*}(\theta)$.
+We repeat this process until $K=10$ distinct actions are identified.
+We compare against two heuristics: combinatorial TS (CTS) and combinatorial UCB (CUCB).
+CTS is run for 3000 rounds to match the runtime of our method, while CUCB is run for 2000 rounds due to its slower execution and additional initialization step.
+Both CTS and CUCB operate in the semi-bandit feedback setting: in each round (corresponding to one bandit instance $\theta$), an outcome function is sampled and $K=10$ actions $\mathcal{A}$ are selected, with the rewards $\mu_{a}(\theta)$ revealed for all $a\in\mathcal{A}$.
+The prior for TS, CTS, and CUCB is set to $\mathcal{N}(0,1)$, consistent with the ground truth.
+In Figure 2, the left subplot reports the expected regret and its standard deviation for EpsilonNet+TS (green), CTS (brown), and CUCB (red), averaged over 30 repetitions.
+The poor performance of CTS and CUCB is expected, as noted in related works.
+The variance of EpsilonNet+TS could be further reduced by running more TS rounds or using a more accurate action-identification method, since we currently use a Gaussian prior—modeling each action’s expected outcome $\mu_{a}$ as an independent Gaussian random variable—rather than a Gaussian process prior, in order to accommodate the most general case where $\mathcal{A}_{\operatorname*{full}}$ is an index set.
+In the right subplot, we fix the length-scale $l=1$ and evaluate runtime and arm pulls of the three methods over 30 repetitions.
+The left axis (purple) shows average runtime $\pm$ one standard deviation, and the right axis (gray) shows the corresponding number of arm pulls.
+EpsilonNet+TS requires 3000 pulls, since it selects $K=10$ optimal actions, with each action identified by running TS for 300 rounds.
+CTS has similar runtime but needs far more pulls, as it runs 3000 rounds with $K=10$ actions pulled per round.
+Conclusion and Future Work
+We proposed a framework for selecting a subset of correlated actions.
+A simple algorithm was introduced and shown to effectively identify near-optimal subsets.
+Future work could develop a stopping criterion.
+When subset size is flexible but sampling new bandit instances is costly, ideas from species discovery (roswell2021conceptual, ) may be useful—treating the discovery of a new action as analogous to discovering a new species.
+Appendix Summary: The appendix is divided into four parts.
+First, Sections A-B provide additional related works and the regret bound of Algorithm 1, which were omitted from the main body due to page limits.
+Second, Sections C to H contain the proofs of all the theorems and lemmas stated both in the main body and in Section B.
+Third, Sections I-J present auxiliary tools used throughout the paper.
+Fourth, Sections K-M include additional examples and a numerical study of Algorithm 1.
+This part features a low-dimensional ($n=3$) action space on the unit sphere to investigate the effect of cluster diameter on expected regret, in line with Theorem 6.
+It also studies two extensions of the varying dependence structure of actions introduced in the main body:
+  • Nonstationary dependence over the action space: Using a nonstationary Gibbs kernel, where the kernel value depends not only on the difference between two inputs $a,a^{\prime}$ but also on the location of $a$.
+  • Limiting case: The collection of expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ becomes i.i.d.
+LLM Usage: Large language models are used for polishing writing, and finding relevant research.
+Appendix A Extra Related Works
+Epsilon Nets have two standard definitions.
+The first, geometric definition [37], requires that radius-$\epsilon$ balls centered at net points cover the set.
+It relates to the covering number and extends to function classes, as in [31].
+The second, measure-theoretic definition [24], requires the net to intersect all subsets of sufficiently large measure.
+The classic $\epsilon$-net algorithm by [14] remains the simplest and most broadly applicable method.
+Later works aim to reduce net size [26, 27, 25] and address online settings [6].
+Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$.
+It is an important topic in high-dimensional probability [37].
+The sharpest known bounds are due to [34].
+and asymptotic convex geometry [29, Chapter 9]
+Appendix B Extra Regret Bounds of Algorithm 1
+We give a matching lower bound:
 Theorem 10.
 Under the same condition of Theorem 7. Let each cluster contains more than one action and satisfies Assumption 5. For the same constant $C$ in Theorem 6 and another absolute constant $c>0$,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\\ &\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
-3.2 Regret Bound for Sub-Gaussian Process
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
 In fact, the Gaussian process assumption can be relaxed to a sub-Gaussian one, where the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
 $\forall u>0,\quad\Pr\left[|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|_{2}^{2}}\right).$
-Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional (Talagrand, 2014) for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm.
+Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional [34] for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm.
 It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$.
@@
 Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process.
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space.
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space.
 Let $\mathcal{A}$ be output of Algorithm 1.
 Then, for a constant $C>0$,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
-4 Generalization and Empirical Validation
-In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings.
-Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling.
-We therefore adopt a dimension-free view by treating the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set $\mathcal{A}_{\operatorname*{full}}$.
-In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$:
-$f(a):=\mu_{a},\quad\forall a\in\mathcal{A}_{\operatorname*{full}}.$
-[…]
-$k(a,a^{\prime})=\sum\nolimits_{i\leq\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(a^{\prime}),$
-where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues.
-Let outcome $f$ be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{\operatorname*{full}}$ and a weight vector $\alpha\in\mathbb{R}^{N}$.
-The function $f$ can be rewritten as
-$\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
-where $\mathbf{f}$ is a vector of coefficients, and $\Phi(a)$ usually refereed as a feature map, has entries $\Phi_{i}(a)=\sqrt{\lambda_{i}}\phi_{i}(a)$.
-If the coefficients $\{\mathbf{f}_{i}\}_{i=1}^{\infty}$ are i.i.d.
-Gaussian, this formulation can be represented as the canonical Gaussian process model in Equation (3).
-Note that in this case, the actual action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{\operatorname*{full}}$.
-4.1 Varying-dependence Actions
-Assuming the outcome function is generated from a kernel inherently implies that the expected outcomes of actions are correlated.
-To study the effect of varying correlation levels among actions, we conduct the following experiments:
-Sampling Outcome Functions from a Kernel.
-We use the stationary RBF kernel (Williams and Rasmussen, 2006), and provide additional illustrations using the non-stationary Gibbs kernel in Supplementary:
-[…]
-where $l$ is a length-scale parameter that control the dependence between actions.
-By (Kanagawa et al., 2018, Theorem 4.12), we sample functions from a RKHS by first constructing the kernel matrix $\mathbf{K}$ with entries $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, and then drawing $f\sim\mathcal{N}(0,\mathbf{K})$.
-Figure 1: Performance comparison on stochastic combinatorial optimization in Equation (6), selecting $K=5$ distinct actions from 15 grid points, with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales.
-Left: Expected regret for our method (green), Thompson Sampling (yellow), UCB (orange), and Successive Halving (SH, blue), averaged over 50 repetitions at each length-scale.
-TS/UCB are run for 3000 rounds; SH is given a budget of 37,000 pulls.
-Right: SH’s expected regret (blue) and pull counts (yellow bars) per round.
-Our method’s performance (green) is shown for comparison.
-SH requires nearly 10,000 pulls—about 3× the number of super arms—to match our method, making it infeasible for large $\mathcal{A}_{\operatorname*{full}}$.
-Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in Equation (1).
-Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
-$\max_{\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (6)
-In general, this problem is non-trivial because the objective function involves an expectation over a collection of random variables whose distribution is neither specified nor directly accessible.
-Moreover, the expression inside the expectation is the maximum over a subset of potentially correlated random variables.
-As a result, small changes in $\mathcal{A}$ can lead to non-smooth variations in the maximum.
-Our method is computationally efficient, avoiding the combinatorial complexity of the optimization problem in Equation (6).
-To assess effectiveness in identifying near-optimal solutions, we compare our method against Thompson Sampling (TS) and Upper Confidence Bound (UCB) in a simple example, using Successive Halving (SH) as a reference for approximating the optimal subset—though SH becomes impractical when the full action space is large.
-We consider a fixed action space consisting of 15 grid points in [0,2][0,2][ 0 , 2 ], using an RBF kernel in Equation (5) to sample outcome functions while varying the length-scale $l\in\{0.5,1,\dots,4\}$.
-In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where the payoff of each action in a round is given by $\mu_{a}$.
-We use exhaustive search over the action space to find the optimal action $\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$.
-For other methods, we treat each $K$-tuple of actions as a super arm, and the payoff of each super arm in a round $\mathcal{A}$ is given by $\max_{a\in\mathcal{A}}\mu_{a}$.
-Thus, these methods are tailored to find the best super arm that maximizes $\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]$, aligning with the same objective.
-In TS/UCB methods, we adopt a bandit feedback setting: at each round, the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}$, and updates its policy accordingly, repeating this process for 3000 rounds, chosen to roughly match the number of super arms ($N=3003$).
-Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB.
-The algorithms are summarized in Supplementary, with the prior of TS set to $\mathcal{N}(0,1)$.
-In the SH method, all super arms are evaluated using a fixed budget of arm pulls (we use the minimum budget required in Karnin et al.
-(2013), $N\log_{2}N\approx 37{,}000$) over a few rounds.
-In each round, the number of remaining super arms is halved.
-This process continues until only the best-performing super arm remains or the budget is exhausted.
-We evaluate expected regret of all methods over $10^{5}$ randomly sampled outcome functions.
-The left subplot of Figure 1 reports the expected regret of our method (green), TS (yellow), UCB (orange) and SH (blue), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions.
-As shown in this subplot, the expected regret decreases as the length-scale increases.
-This is because the length-scale $l$ of the RBF kernel controls the number of approximately independent actions in the process $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$: When actions $a$ and $a^{\prime}$ are far apart relative to the length-scale, the kernel value becomes very small, so the covariance between the function values $\mu_{a}$ and $\mu_{a^{\prime}}$ is close to zero.
-Regarding computational complexity: in a given bandit instance $\theta$, identifying the optimal action $a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)$ is a well-studied problem in both best-arm identification and GP optimization.
-However, if computing $a^{*}(\theta)$ is not computationally efficient, then identifying the optimal super arm, as in SH—whose search space grows combinatorially—is infeasible.
-Next, in the right subplot of Figure 1, we examine sample complexity.
-We run our method (green) and Successive Halving (SH, blue) for 50 additional repetitions, using a fixed length-scale $l=1$.
-Since SH proceeds in rounds, we track the current best super arm and the total number of super arm pulls so far after each round.
-The expected regret of the current best super arm, is evaluated over another $10^{5}$ randomly sampled outcome functions.
-The blue solid curves and shades for SH represent the mean $\pm$ one standard deviation of expected regret over 50 repetitions over each SH round.
-The yellow bars (corresponding to the right y-axis) show the number of super arm pulls by SH at the end of each round.
-For comparison, the green curve and shaded region represent the mean $\pm$ one standard deviation of expected regret for our method over the same 50 repetitions.
-As shown, SH’s performance roughly matches ours after three rounds, requiring 9005 super arm pulls—approximately three times the number of super arms.
-If the action space $|\mathcal{A}_{\operatorname*{full}}|$ is already large, running SH is intractable.
-4.2 Conclusion and Future Work
-We proposed a framework for selecting a subset of correlated actions, modeling payoff correlations with a Gaussian process and extending to the sub-Gaussian case.
-A simple algorithm was introduced and shown to effectively identify near-optimal subsets.
-A key direction for future work is developing a stopping criterion.
-When the subset size is flexible but sampling new bandit instances is costly, ideas from species discovery Roswell et al.
-(2021) may be useful—treating the discovery of a new action as analogous to discovering a new species.
-One could use Turing’s formula to estimate the probability of finding a new action, or frame the process as a one-armed bandit that stops when the discovery probability becomes sufficiently low.
-Appendix A Connection to Bayesian Regret of Multi-Armed Bandits
-Given a subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$, let $\mu^{*}(\mathcal{A},\theta):=\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ be the highest expected outcome in a bandit instance $\theta$.
-We can equivalently rewrite the regret in Equation (1):
-$\textnormal{Regret}(\theta)=\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\mu^{*}(\mathcal{A},\theta).$
-Bandit Regret: In the classic multi-armed bandit setting [Auer et al., 2002a, Lattimore and Szepesvári, 2020], outcome of an arm $a\in\mathcal{A}_{\operatorname*{full}}$ in a bandit instance $\theta$ is generated i.i.d.
-according to an unknown but fixed law with expectation $\mu_{a}(\theta)$.
-A decision-maker would follow a policy $\pi$ that chooses the next arm based on the sequence of rounds and obtained outcomes.
-The bandit regret of policy $\pi$ in this bandit environment $\theta$ over $N$ rounds is defined as the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action and following a policy $\pi$ to choose actions.
-Bayesian bandit regret [Agrawal and Goyal, 2012, Lattimore and Szepesvári, 2020] is the expectation of this bandit regret, taken over the randomness in the distribution of bandit instances, and the action chosen by the policy:
-$\begin{split}\textnormal{BanditRegret}(\theta):=&\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta),\\ \textnormal{BayesianBanditRegret}:=&\mathbb{E}\!\left[\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta)\right],\end{split}$
-where $A_{t}$ is the action chosen by policy $\pi$ in round $t$.
-Since we consider settings where only a subset of the full action space is accessible, the action chosen by the policy can only from the subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$.
-This quantifies the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action from the full action set $\mathcal{A}_{\operatorname*{full}}$ and following policy $\pi$ to choose actions from the restricted subset $\mathcal{A}$.
-Next, we establish the connection between the regret defined in this paper and the Bayesian bandit regret.
-A.1 Optimal policy
-If $\pi$ is the optimal policy, it selects the best action in $\mathcal{A}$ in every round.
-Therefore, the regret of only having access to $\mathcal{A}$ instead of $\mathcal{A}_{\operatorname*{full}}$ for $N$ rounds, is determined by the optimal actions in $\mathcal{A}$ and $\mathcal{A}_{\operatorname*{full}}$:
-[…]
-A.2 Thompson sampling policy
-If $\pi$ is the Thompson sampling policy, the Bayesian regret is bounded by
-[…]
-$=$ $\mathbb{E}\left[\sum_{t=1}^{N}\left(\mu^{*}(\mathcal{A})-\mu_{A_{t}}+\mu^{*}(\mathcal{A}_{\operatorname*{full}})-\mu^{*}(\mathcal{A})\right)\right]$
-$\leq$ $C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}],$
-where $|\mathcal{A}|$ denotes the cardinality, and $C>0$ is a constant.
-The inequality uses the Bayesian (bandit) regret of Thompson sampling (Theorem 36.1 in Lattimore and Szepesvári [2020]).
-Note that with access to the full action space, the Bayesian regret is bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$.
-However, this can be worse than restricting to a subset $\mathcal{A}$ when $\mathbb{E}_{\theta}[\textnormal{Regret}]$ is sufficiently small and the time horizon $N$ is finite.
-Intuitively speaking, using a well-chosen subset of representative actions instead of the full action space can reduce Bayesian regret in multi-armed bandits.
-Appendix B Proof of Lemma 4
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+Proofs are in Appendices G and H, respectively.
+Appendix C Proof of Lemma 4
 Statement: Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$.
@@
 where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$.
-The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [TOPSØE1, 2007].
+The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [35].
 Therefore,
@@
 $=$ $1-|R_{\epsilon}|\exp(-K\epsilon)\geq 1-1/\epsilon\cdot\exp(-K\epsilon),$
-where the first inequality uses union bound, the second inequality uses Equation (7).
+where the first inequality uses union bound, the second inequality uses equation 6.
 The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$.
 ∎
-Appendix C Proof of Theorem 7
+Appendix D Proof of Theorem 6
+Under Assumption 5, the regret can be decomposed as follows; note that this Lemma is used only to derive lower bounds.
 Lemma 12 (Regret decomposition of reference subsets).
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space and an arbitrary reference subset $\mathcal{A}$.
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space and an arbitrary reference subset $\mathcal{A}$.
 For each $\ell\leq m$, define a Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$.
@@
 Proof.
-[…]
-[…]
-[…]
-[…]
+$\mathbb{E}X=$ $\int_{0}^{\infty}\Pr[X\geq u]\;du$
+$=$ $\int_{0}^{u_{0}}\Pr[X\geq u]\;du+\int_{u_{0}}^{\infty}\Pr[X\geq u]\;du$
+$\leq$ $u_{0}+\frac{1}{u_{0}}\int_{u_{0}}^{\infty}u\cdot\Pr[X\geq u]\;du$
+$\leq$ $u_{0}+\frac{c}{u_{0}}\int_{u_{0}}^{\infty}u\cdot\exp{\left(-\frac{u^{2}}{\epsilon^{2}}\right)}\;du$
 $=$ $u_{0}+\exp\left(-\frac{u_{0}^{2}}{\epsilon^{2}}\right)\cdot\frac{c\epsilon^{2}}{2u_{0}}$
 $=$ $\epsilon\sqrt{\log c}+\frac{\epsilon}{2\sqrt{\log c}}\leq C\epsilon\sqrt{\log c},$
-where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of Vershynin [2018]). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
-Lemma 15 (Borell-TIS inequality; Lemma 2.4.7 of [Talagrand, 2014]).
+where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of [37]). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
+Lemma 14 (Borell-TIS inequality; Lemma 2.4.7 of [34]).
 Given a set $\mathcal{S}$, and a zero-mean Gaussian process $\left(X_{a}\right)_{a\in\mathcal{S}}$.
@@
 Statement of Theorem 6:
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$.
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$.
 Then, for some constant $C>0$,
@@
 $\operatorname*{Regret}=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell}\leq\max_{\ell^{\prime}\leq m}Y_{\ell^{\prime}},$
-where the first equality follows from the definition of regret in Equation (1), and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$.
+where the first equality follows from the definition of regret in equation 1, and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$.
 The inequality holds because $a_{\ell}\in\mathcal{A}$, and hence $\max_{a\in\mathcal{A}}\mu_{a}\geq\mu_{a_{\ell}}$.
@@
 $\mathbb{E}\operatorname*{Regret}\leq\mathbb{E}\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\leq\max_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}+C\epsilon\sqrt{\log m},$ (9)
-where the first inequality uses Equation (8).
+where the first inequality uses equation 7.
 The second inequality uses $\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\leq\mathbb{E}Y_{\ell}+\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$.
-The last inequality uses Equation (9) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
+The last inequality uses equation 8 and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
 Lower bound:
 $\mathbb{E}\operatorname*{Regret}\geq\mathbb{E}\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq\min_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}-C\epsilon\sqrt{\log m},$ (10)
-where the first inequality uses Equation (8) under Assumption 6.
+where the first inequality uses equation 7 under Assumption 5.
 The second inequality uses $\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\geq\mathbb{E}Y_{\ell}-\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$.
-The last inequality uses Equation (9) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
+The last inequality uses equation 8 and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
 ∎
-Appendix D Proof of Theorem 8
+Appendix E Proof of Theorem 7
+An important consequence of assuming $\mathbb{E}_{\theta}\left[\mu_{a}\right]=0$ for all $a\in\mathcal{A}_{\operatorname*{full}}$ is the following property:
 Lemma 15 (Transformation invariance).
@@
 where the last equality uses $\mathbb{E}\left[\langle x,\theta\rangle\right]=\left\langle x,\mathbb{E}[\theta]\right\rangle=0$. ∎
+The key tool for decomposing the regret of Algorithm 1 into that from a reference subset and that from missing clusters is the following Lemma:
 Lemma 16.
@@
 Let $\mathcal{A}$ be output of Algorithm 1.
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
 Then, for the same constant $C>0$ in Theorem 6,
@@
 where the first equality uses tower rule.
-The first inequality uses Equations (12-13), and Lemma 17.
+The first inequality uses Equations (11-12), and Lemma 16.
 The last equality uses tower rule again.
 The last inequality uses Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$.
-Note that the Gaussian process assumption is not used in Equation (14); it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$.
-By applying Equation (10) in the proof of Theorem 7, we bound this term and complete the proof.
+Note that the Gaussian process assumption is not used in equation 13; it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$.
+By applying equation 9 in the proof of Theorem 6, we bound this term and complete the proof.
 ∎
-Appendix E Proof of Theorem 10
+Appendix F Proof of Theorem 9
+To analyze the worst-case behavior of Theorem 7, we first examine the term $\mathbb{E}_{q}\left[(1-q(r))^{K}\right]$ in the upper bound, summarized in the following Lemma:
 Lemma 17.
@@
 Proof.
-Let =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=|\mathcal{R}|$.
+Let $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=|\mathcal{R}|$.
 Let $q_{i}:=q(r_{i})$ denote the probability mass at each support point.
@@
 $f^{\prime}(q_{\ell})=(1-q_{\ell})^{K-1}\big(1-(K+1)q_{\ell}\big),$
-which is positive on the interval $\left[0,\frac{1}{K+1}\right)$ and negative on the interval $\left(\frac{1}{K+1},1\right)$. Therefore, $f(q_{\ell})$ attains its maximum over [0,1)[0,1)[ 0 , 1 ) at
+which is positive on the interval $\left[0,\frac{1}{K+1}\right)$ and negative on the interval $\left(\frac{1}{K+1},1\right)$. Therefore, $f(q_{\ell})$ attains its maximum over $[0,1)$ at
 $q^{*}_{\ell}=\frac{1}{K+1},$
@@
 where the right inequality uses that $f(q_{\ell})\geq 0$ for $q_{\ell}\in[0,1]$.
-Case II $m\geq K+1$: Consider the relaxed maximization problem of Equation (15):
+Case II $m\geq K+1$: Consider the relaxed maximization problem of equation 14:
 $\max_{q_{1},\dots,q_{m}\in[0,1)}\sum_{\ell=1}^{m}f(q_{\ell})\quad\text{subject to }\sum_{\ell=1}^{m}q_{\ell}\leq 1.$ (15)
@@
 $(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda,\quad\text{with }\lambda\geq 0.$
-To find critical points of Equation (16), we solve the system:
+To find critical points of equation 15, we solve the system:
 $(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda\geq 0\quad\forall\ell\leq m,\quad\text{and}\quad\sum_{\ell=1}^{m}q_{\ell}\leq 1.$
@@
 Due to the assumption of $m\geq K+1$, we have $\frac{1}{m}\leq\frac{1}{K+1}$, so any choice of $c\in\left[0,\frac{1}{m}\right]$ satisfies the constraint $\sum_{\ell=1}^{m}q_{\ell}\leq 1$ and is feasible for the system.
-Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of Equation (16). The corresponding objective value is:
+Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of equation 15. The corresponding objective value is:
 $\sum_{\ell=1}^{m}f(c)=m\cdot c\cdot(1-c)^{K},$
@@
 which is negative for $q_{\ell}\leq\frac{1}{K+1}$, because then $2-(K+1)q_{\ell}>0$. Hence, the Hessian is negative definite, and the critical point is a local (and thus global) maximum.
-Finally, note that Equation (16) is a relaxation of Equation (15).
+Finally, note that equation 15 is a relaxation of equation 14.
 While the optimum of the original problem is upper bounded by that of the relaxed problem, the optimal solution to the relaxed problem also lies within the feasible region of the original problem.
-Therefore, the maximum of Equation (15) is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$.
+Therefore, the maximum of equation 14 is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$.
 ∎
@@
 Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{\operatorname*{full}}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{\operatorname*{full}},\epsilon)$.
-Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition =$\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
+Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
 $r_{\ell}=\left\{a\in\mathcal{A}_{\operatorname*{full}}:\pi(a)=a_{\ell}\right\}.$
@@
 $\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}\right]\leq\mathbb{E}_{\theta}\left[\max_{a\in B(a_{\ell},\epsilon)}\mu_{a}\right]=\mathbb{E}_{\theta}\left[\max_{a\in\epsilon B_{2}^{n}}\mu_{a}\right]\leq\epsilon\sqrt{n},$ (17)
-where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 16 to shift; and the final bound uses Claim 3 from Supplementary K. Then,
+where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 15 to shift; and the final bound uses Claim 3 from Supplementary I. Then,
 $\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
@@
 where the first inequality follows from the algorithm-dependent upper bound in Theorem 7.
-The second inequality follows from Equations (17) and (18), and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma 18).
+The second inequality follows from Equations equation 16 and equation 17, and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma 17).
 The last inequality comes from
@@
 where the first inequality uses $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)=\tfrac{1}{2}(\log m-\log(2K+1))$.
-Also, $\log\left(\tfrac{2K}{2K+1}\right)=\log\left(1-\tfrac{1}{2K+1}\right)\leq-\tfrac{1}{2K+1}$ due to the inequality $\log(1-X)\leq-X$ for $X<1$, and for a large $K$, we approximate $-\tfrac{K}{2K+1}\approx-\frac{1}{2}$.
+Also, $\log\left(\tfrac{2K}{2K+1}\right)=\log\left(1-\tfrac{1}{2K+1}\right)\leq-\tfrac{1}{2K+1}$ due to the inequality $\log(1-X)\leq-X$ for $X<1$, and for a large $K$, we approximate $\tfrac{K}{2K+1}\approx-\frac{1}{2}$.
 The last inequality uses the assumption $K\geq\frac{1}{2}\left(\frac{mM^{2}}{\epsilon^{2}e}-1\right)$.
-Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of Vershynin [2018].
+Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of [37].
 Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$.
 ∎
-Appendix F Proof of Theorem 11
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 6.
+Appendix G Proof of Theorem 10
+To derive a matching lower bound, we decompose the regret of Algorithm 1 into that from a reference subset and from missing clusters, using Lemma 16:
+Statement of Theorem 10
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 5.
 Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
@@
 The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 16.
-The second inequality follows from Equation (19).
-The third inequality uses the Pólya-Szegő inequality [Dragomir, 2004], corresponding to a constant $c>0$:
+The second inequality follows from equation 18.
+The third inequality uses the Pólya-Szegő inequality [11], corresponding to a constant $c>0$:
 $c:=\frac{1}{2}\left(\sqrt{\frac{M_{1}M_{2}}{m_{1}m_{2}}}+\sqrt{\frac{m_{1}m_{2}}{M_{1}M_{2}}}\right),$
 where $m_{1},m_{2},M_{1},M_{2}$ are constants such that
-[…]
+$0<m_{1}\leq(1-q(r_{\ell}))^{K}\leq M_{1},\;0<m_{2}\leq\mathbb{E}\left[Y_{\ell}\Big|\,a^{*}(\theta)\in r_{\ell}\right]\leq M_{2},\;\;\forall\ell\leq m.$
 Note that $m_{2}$ is positive due to the assumption that each cluster contains more than one action.
@@
 ∎
-Appendix G Proof of Theorem 12
+Appendix H Proof of Theorem 11
 Definition 2.
@@
 $\forall u>0,\Pr[|\mu_{a}-\mu_{a^{\prime}}|\geq u]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|^{2}_{2}}\right).$
-Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional [Talagrand, 2014] for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm.
+Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional [34] for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm.
 It is well-known that under the assumption that $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a Gaussian process, this gives the tightest bound, for a universal constant $L$:
@@
 Within the proof of the upper bound for general sub-Gaussian process, it also derives the deviation bound for the term $\max_{a\in\mathcal{S}}\mu_{a}$:
-Lemma 19 (Theorem 2.2.22 of Talagrand [2014]).
+Lemma 18 (Theorem 2.2.22 of [34]).
 Let $\{\mu_{a}\}_{a\in\mathcal{S}}$ be a mean-zero sub-Gaussian process, then there exists a constant $c>0$:
@@
 Let $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ be a mean-zero sub-Gaussian process.
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, and an arbitrary reference subset $\mathcal{A}:=\{a_{\ell}\}_{\ell\leq m}$.
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, and an arbitrary reference subset $\mathcal{A}:=\{a_{\ell}\}_{\ell\leq m}$.
 Fix $\ell$, define a non-negative random variable
@@
 Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process.
-Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space.
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space.
 Then, for the same constant $C>0$,
@@
 $\leq$ $C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
-The first inequality follows from Equation (14) in the proof of Theorem 8. The last inequality uses Lemma 20. ∎
-Appendix H The effect of clustering structure on regret
+The first inequality follows from equation 13 in the proof of Theorem 7. The last inequality uses Lemma 19. ∎
+Appendix I Properties of Gaussian width
+Given a set $\mathcal{S}\in\mathbb{R}^{n}$, the term $\mathbb{E}[\max_{a\in\mathcal{S}}\mu_{a}]$ where $\theta\sim\mathcal{N}(0,I)$ is called Gaussian (mean) width.
+Claim 1: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
+$\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,-\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle-a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
+where the first equality uses $\theta$ and $\theta$ are identically distributed. The third equality uses for any $a\in\mathcal{S}$, it holds that $a\in-\mathcal{S}$.
+Claim 2: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]\leq\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$.
+Let $a^{*}(-\mathcal{S},\theta)$ denote the optimal action in $\mathcal{S}$ for bandit instance $\theta$.
+$2\cdot\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]+\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$
+$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta),\;\theta\rangle\right]+\mathbb{E}\left[\langle a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
+$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
+$\leq$ $\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right],$
+where the first equality uses Claim 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$.
+The third equality uses linearity of expectation.
+The inequality uses that the vector $a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta)$ belongs to the set of vectors $\{a-a^{\prime}:a,a^{\prime}\in\mathcal{S}\}$.
+In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need inequality to prove the next claim.
+Claim 3: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n}$.
+$\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]$
+$\leq$ $\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$
+$\leq$ $\frac{1}{2}\mathbb{E}\max_{a,a^{\prime}\in\mathcal{S}}\|\theta\|_{2}\|a-a^{\prime}\|_{2}$
+$\leq$ $\frac{1}{2}\mathbb{E}\operatorname{diam}(\mathcal{S})\|\theta\|_{2}\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n},$
+where the first equality uses the definition of $\mu_{a}$.
+The first inequality uses Claim 2.
+The second inequality uses Cauchy-Schwarz inequality.
+The third inequality uses the definition of $\operatorname{diam}(\cdot)$.
+The last inequality uses $\mathbb{E}\|\theta\|_{2}\leq\sqrt{n}$.
+Appendix J Bounds of expected maximum of Gaussian
+Let $X_{1},\dots,X_{N}$ be $N$ random Gaussian variables (no necessarily independent) with zero mean and variance of marginals smaller than $\sigma^{2}$, then
+$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]\leq\sigma\sqrt{2\log N}.$
+Proof.
+for any $\delta>0$,
+$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]=$ $\frac{1}{\delta}\mathbb{E}\left[\log\exp(\delta\max_{i=1,\dots,N}X_{i})\right]\leq\frac{1}{\delta}\log\mathbb{E}\left[\exp(\delta\max_{i=1,\dots,N}X_{i})\right]$
+$=$ $\frac{1}{\delta}\log\mathbb{E}\left[\max_{i=1,\dots,N}\exp(\delta X_{i})\right]\leq\frac{1}{\delta}\log\sum_{i=1}^{N}\mathbb{E}\left[\exp(\delta X_{i})\right]$
+$\leq$ $\frac{1}{\delta}\log\sum_{i=1}^{N}\exp(\sigma^{2}\delta^{2}/2)=\frac{\log N}{\delta}+\frac{\sigma^{2}\delta}{2},$
+where the first inequality uses Jensen’s inequality. Taking $\delta:=\sqrt{2(\log N)/\sigma^{2}}$ yields the results. ∎
+Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [17]:
+$\mathbb{E}_{\theta}\left[\max_{i=1,\dots,N}X_{i}\right]\geq\frac{\sigma\sqrt{\log N}}{\sqrt{\pi\log 2}}.$
+Appendix K The effect of clustering structure on regret
 In Figure 3, we study the effect of the cluster diameters (controlled by a spread parameter) on regret.
@@
 The middle and right subplots show example action spaces (blue dots) for spread values 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
-Appendix I Varying-dependence actions with RBF/Gibbs kernels
-We study the effect of varying action dependence using a Gaussian process with a kernel.
-To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel [Williams and Rasmussen, 2006].
-[…]
+Appendix L Varying-dependence actions with RBF/Gibbs kernels
+We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel [39].
+$\begin{split}\underset{\text{RBF}}{k}(a,a^{\prime})=&\exp\left(-\frac{\|a-a^{\prime}\|^{2}}{2l^{2}}\right),\\ \underset{\text{Gibbs}}{k}(a,a^{\prime})=&\sqrt{\frac{2\,l(a)l(a^{\prime})}{l(a)^{2}\!+\!l(a^{\prime})^{2}}}\exp\left(\!-\!\frac{\|a-a^{\prime}\|^{2}}{l(a)^{2}\!+\!l(a^{\prime})^{2}}\right),\end{split}$ (19)
 where $l$ is a length-scale parameter and $l(a):=0.1+0.9\cdot\exp(-\|a\|^{2})$ is a location-dependent length-scale function.
@@
 We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$.
-Under either kernels defined in Equation (20), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{\operatorname*{full}}$.
-In this way, we sample functions from a RKHS function class; See [Kanagawa et al., 2018, Theorem 4.12].
-Figure 3: Experiments with outcome functions sampled from RBF/Gibbs kernels in Equation (20).
-Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in [0,2][0,2][ 0 , 2 ] (blue curves, right y-axis).
+Under either kernels defined in equation 19, the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{\operatorname*{full}}$.
+In this way, we sample functions from a RKHS function class; See [18, Theorem 4.12].
+Figure 4: Experiments with outcome functions sampled from RBF/Gibbs kernels in equation 19.
+Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in $[0,2]$ (blue curves, right y-axis).
 The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points.
-To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval [0,2][0,2][ 0 , 2 ], using Gibbs kernel in Equation (20) to sample outcome functions.
+To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval $[0,2]$, using Gibbs kernel in equation 19 to sample outcome functions.
 To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid.
@@
 The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
-Appendix J Independent and identically distributed actions
+Appendix M Independent and identically distributed actions
 As an extreme case, suppose that $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ is a set of i.i.d.
@@
 This corresponds to the canonical process in which the action space is given by the orthonormal basis of $\mathbb{R}^{n}$, where $n=|\mathcal{A}_{\operatorname*{full}}|$.
-To see this, we associate each action $a$ with a unit vector $e_{a}$, which has a value of 111 at the $a$th coordinate and 0 elsewhere, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$.
+To see this, we associate each action $a$ with a unit vector $e_{a}$, which has a value of $1$ at the $a$th coordinate and 0 elsewhere, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$.
 With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ consists of mutually independent random variables.
@@
 $\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\mathbb{E}_{\theta}\max_{i=1,\dots,n}\theta_{i}=\mathbb{E}\max_{i=1,\dots,n}X_{i},$
-where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary L, we have
+where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary J, we have
 $\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\sqrt{2\log n},\quad\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}}\mu_{a}\leq\sqrt{2\log|\mathcal{A}|}.$
-Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in Equation (1),
+Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in equation 1,
 $\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}\leq\mathbb{E}_{\theta}[\textnormal{Regret}]\leq\sqrt{2\log n}-\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}.$
 As a result, any algorithm, including Algorithm 1, would perform poorly unless the subset size $|\mathcal{A}|$ is sufficiently large.
-Appendix K Properties of Gaussian width
-Given a set $\mathcal{S}\in\mathbb{R}^{n}$, the term $\mathbb{E}[\max_{a\in\mathcal{S}}\mu_{a}]$ where $\theta\sim\mathcal{N}(0,I)$ is called Gaussian (mean) width.
-Claim 1: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
-$\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,-\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle-a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
-where the first equality uses $-\theta$ and $\theta$ are identically distributed. The third equality uses for any $a\in\mathcal{S}$, it holds that $-a\in-\mathcal{S}$.
-Claim 2: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]\leq\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$.
-Let $a^{*}(-\mathcal{S},\theta)$ denote the optimal action in $\mathcal{S}$ for bandit instance $\theta$.
-$2\cdot\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]+\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$
-$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta),\;\theta\rangle\right]+\mathbb{E}\left[\langle a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
-$=$ $\mathbb{E}\left[\langle a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta),\;\theta\rangle\right]$
-$\leq$ $\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right],$
-where the first equality uses Claim 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$.
-The third equality uses linearity of expectation.
-The inequality uses that the vector $a^{*}(\mathcal{S},\theta)+a^{*}(-\mathcal{S},\theta)$ belongs to the set of vectors $\{a-a^{\prime}:a,a^{\prime}\in\mathcal{S}\}$.
-In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need this inequality to prove the next claim.
-Claim 3: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n}$.
-$\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]$
-$\leq$ $\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$
-$\leq$ $\frac{1}{2}\mathbb{E}\max_{a,a^{\prime}\in\mathcal{S}}\|\theta\|_{2}\|a-a^{\prime}\|_{2}$
-$\leq$ $\frac{1}{2}\mathbb{E}\operatorname{diam}(\mathcal{S})\|\theta\|_{2}\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n},$
-where the first equality uses the definition of $\mu_{a}$.
-The first inequality uses Claim 2.
-The second inequality uses Cauchy-Schwarz inequality.
-The third inequality uses the definition of $\operatorname{diam}(\cdot)$.
-The last inequality uses $\mathbb{E}\|\theta\|_{2}\leq\sqrt{n}$.
-Appendix L Bounds of expected maximum of Gaussian
-Let $X_{1},\dots,X_{N}$ be $N$ random Gaussian variables (no necessarily independent) with zero mean and variance of marginals smaller than $\sigma^{2}$, then
-$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]\leq\sigma\sqrt{2\log N}.$
-Proof.
-for any $\delta>0$,
-$\mathbb{E}\left[\max_{i=1,\dots,N}X_{i}\right]=$ $\frac{1}{\delta}\mathbb{E}\left[\log\exp(\delta\max_{i=1,\dots,N}X_{i})\right]\leq\frac{1}{\delta}\log\mathbb{E}\left[\exp(\delta\max_{i=1,\dots,N}X_{i})\right]$
-$=$ $\frac{1}{\delta}\log\mathbb{E}\left[\max_{i=1,\dots,N}\exp(\delta X_{i})\right]\leq\frac{1}{\delta}\log\sum_{i=1}^{N}\mathbb{E}\left[\exp(\delta X_{i})\right]$
-$\leq$ $\frac{1}{\delta}\log\sum_{i=1}^{N}\exp(\sigma^{2}\delta^{2}/2)=\frac{\log N}{\delta}+\frac{\sigma^{2}\delta}{2},$
-where the first inequality uses Jensen’s inequality. Taking $\delta:=\sqrt{2(\log N)/\sigma^{2}}$ yields the results. ∎
-Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [Kamath, 2015]:
-$\mathbb{E}_{\theta}\left[\max_{i=1,\dots,N}X_{i}\right]\geq\frac{\sigma\sqrt{\log N}}{\sqrt{\pi\log 2}}.$

### v4

@@
-Representative Action Selection for Large Action Space Meta-Bandits
+Representative Action Selection for Large Action Space Bandit Families
 Abstract
 We study the problem of selecting a subset from a large action space shared by a family of bandits, with the goal of achieving performance nearly matching that of using the full action space.
-We assume that similar actions tend to have related payoffs, modeled by a Gaussian process.
-To exploit this structure, we propose a simple $\epsilon$-net algorithm to select a representative subset.
-We provide theoretical guarantees for its performance and compare it empirically to Thompson Sampling and Upper Confidence Bound.
+Indeed, in many natural situations, while the nominal set of actions may be large, there also exist significant correlations between the rewards of different actions.
+In this paper we propose an algorithm that can significantly reduce the action space when such correlations are present, without the need to a-priori know the correlation structure.
+We provide theoretical guarantees on the performance of the algorithm and demonstrate its practical effectiveness through empirical comparisons with Thompson Sampling and Upper Confidence Bound methods.
 1 Introduction
-We study a family of bandits that share a common but extremely large action space.
-We aim to understand whether it is possible—and how—to select a smaller set of representative actions that performs nearly as well as the full action space across all bandit instances.
-To build intuition, imagine a pharmacy preparing its inventory for the upcoming season.
-The available drugs (actions) are nearly infinite, and each customer (bandit) has unique characteristics.
-If two drugs share similar ingredients, their effects on a patient are likely to be similar.
-Likewise, if two patients have comparable health indices, a drug is likely to have similar effects on both.
-By modeling the expected outcome of each drug for each patient as a Gaussian process, we can capture these correlations.
-This correlation structure makes subset selection particularly interesting: if two drugs treat the same disease, choosing one may suffice, while drugs for distinct diseases might both be kept.
-In addition, demand also matters: drugs for rare illnesses can be excluded to save space, while flu medications should be stocked during flu season.
-Different from prior approaches in multi-armed bandits (MAB) (lai1985asymptotically, ) that aim for identifying either a single best action or a subset that achieves high cumulative outcomes for a fixed bandit, our objective focuses on selecting a subset that is likely to contain the best action, or one whose best element performs nearly as well for a family of bandits.
-This problem can be seen as a large-scale combinatorial optimization under uncertainty, with applications where decisions involve a vast number of possibilities but are constrained by computational or time limitations for evaluating all options, e.g., inventory management, online recommendations.
-Consider the following setting: In a bandit, if a decision-maker plays an action with a fixed but unknown feature vector $a\in\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, they observe a random outcome taking values in $\mathbb{R}$.
-We define the expected outcome of playing action $a$ in this bandit as $\mu_{a}(\theta):=\langle a,\theta\rangle$ where this bandit instance $\theta\in\mathbb{R}^{n}$ is drawn from an unknown multivariate Gaussian distribution.
-Thus, the collection of random variables $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ forms a Gaussian process (GP) (vershynin2018high, , Chapter 7), while a more general sub-Gaussian assumption is considered in the Supplementary.
-Our setting aligns with that of contextual bandits (dani2008stochastic, ), and the restriction to linear functions may be partially mitigated by allowing $n$ to be unbounded.
-Consider that the decision-maker has access to a fixed action subset $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$.
-For a given bandit instance $\theta$, if the optimal action over the full space lies within $\mathcal{A}$, the decision-maker benefits from reduced suboptimal actions to explore.
-Conversely, if the optimal action lies outside $\mathcal{A}$, regret arises from being unable to select this best action.
-Thus, for a bandit instance $\theta$, we define the regret as the difference in expected outcome between having access to the full action space versus being restricted to the subset:
-$\boxed{\textnormal{Regret}(\theta):=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}}(\theta),}$ (1)
-which depends on the sampled bandit instance and is therefore a random variable.
-Our objective is to identify a small subset $\mathcal{A}$ that minimizes the expected regret $\mathbb{E}_{\theta}\left[\operatorname*{Regret}\right]$ over all possible bandit instances, making the underlying optimization both stochastic and combinatorial.
-This objective is motivated by practical considerations.
-The classic Bayesian regret in the bandit literature (agrawal2012analysis, ) typically scales with the number of available actions.
-However, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower.
-This is especially beneficial when the action space is large, as even the initialization phase can be computationally expensive.
-To see this, we can decompose the bandit regret as follows:
-$\textnormal{BayesianBanditRegret}:=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\mu_{A_{t}}(\theta)\right]$
-$=\mathbb{E}\sum_{t=1}^{N}\left[\max_{a\in\mathcal{A}}\mu_{a}(\theta)\!-\!\mu_{A_{t}}+\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)\!-\!\max_{a\in\mathcal{A}}\mu_{a}(\theta)\right]$
-$\leq C\sqrt{|\mathcal{A}|\cdot N\log N}+N\cdot\mathbb{E}_{\theta}[\textnormal{Regret}].$
-In the first equality, $A_{t}\in\mathcal{A}$ denotes the action chosen by a policy, e.g., Thompson Sampling (agrawal2012analysis, ), in round $t$, and $N$ is the number of rounds.
-The expectation is taken over the randomness in the distribution of bandit instances and in actions selected by the policy.
-The inequality follows from the well-known regret bounds for Thompson Sampling (lattimore2020bandit, ) [^1] This bound could be relaxed if the feature vector $a$ is known, or if $a$ can be incorporated as the input of a kernel function, as discussed in related works.
-However, we use this general bound because we also consider the most general case where $\mathcal{A}_{\operatorname*{full}}$ is simply an index set, e.g., $\mathcal{A}_{\operatorname*{full}}:=\{\textnormal{treatment A},\textnormal{treatment B},\textnormal{treatment C}\}$., where $|\mathcal{A}|$ denotes the cardinality of the action set and $C>0$ is a constant.
-Note that if the policy has access to the full action space, the Bayesian bandit regret is instead bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$.
+The multi-armed bandit (MAB) problem (Lai & Robbins, 1985; Auer et al., 2002a) models sequential decision-making under uncertainty: at each round, an agent chooses an arm from a fixed set and receives a stochastic reward drawn from an unknown distribution.
+The objective is to maximize cumulative reward by balancing exploration and exploitation.
+In modern applications, action spaces are often extremely large, yet many related decision problems share this same set of actions.
+We ask a fundamental question: Can we select a small, representative subset of actions that performs nearly as well as the full action space across all such related problems?
+If possible, this would dramatically reduce computational and statistical complexity with only controllable performance loss.
+In this work, we formalize this question and investigate when and how such a small, universally effective action subset can be constructed.
+As an example, consider a pharmacy preparing its inventory for the upcoming season.
+The set of available drugs will correspond to actions, and can be extremely large.
+Each patient will be represented by a bandit instance, i.e., a unique mapping from drugs to health outcomes based on their physiology and history.
+However, when multiple drugs contain similar ingredients, their effects on patients are likely to be similar.
+Therefore, we can retain just one or a few representative drugs from such a group without meaningfully reducing coverage or effectiveness.
+In the classical multi-armed bandit problem (MAB) (Lai & Robbins, 1985; Lattimore & Szepesvári, 2020), the objective is to maximize cumulative reward over a time horizon $T$ for a single, fixed bandit.
+This is achieved by exploiting a single optimal action after exploration over the action space.
+In contrast, our objective is to select an action subset that likely contains near-optimal actions jointly for a family of bandits.
+This problem is a large-scale combinatorial optimization under uncertainty (Van Hentenryck & Bent, 2006; Li & Yuan, 2013).It arises in settings with too many options to evaluate exhaustively— such as route planning (Feng et al., 2025) and inventory management (Namir et al., 2022)—due to computational or time constraints.
+Consider a Bayesian multi armed bandits setting, (Slivkins et al., 2019; Lattimore & Szepesvári, 2020): Let $\mathcal{F}$ be a family of bandits which share an action space $\mathcal{A}_{\operatorname*{full}}$, and let $\mathcal{P}$ be a probability distribution over $\mathcal{F}$.
+For a bandit instance $\theta\in\mathcal{F}$, let $\mu_{a}=\mu_{a}(\theta)$ denote the expected reward of an action $a\in\mathcal{A}_{\operatorname*{full}}$ given $\theta$.
+Each $\mu_{a}$ is then a random variable as $\theta$ is sampled from $\mathcal{P}$, and the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ constitutes a stochastic process indexed by the action space.
+In this paper we consider the similarity structure on the actions $\mathcal{A}_{\operatorname*{full}}$ that is induced by the correlations of the variables $\mu_{a}$.
+In this setting, the actions $a$ and $a^{\prime}$ will be considered similar if the rewards $\mu_{a}$ and $\mu_{a}^{\prime}$ are similar for most bandits sampled from $\mathcal{P}$.
+Note that this notion does not depend on the nominal properties of the set $\mathcal{A}_{\operatorname*{full}}$, such as the dimension in which $\mathcal{A}_{\operatorname*{full}}$ happens to be represented, but depends only on the statistical properties of the family of bandits $\mathcal{F}$ and the distribution $\mathcal{P}$.
+This provides a robust practical modeling framework which in turn may be used as the basis for the action subset selection — in a cluster of highly positively correlated actions, one may retain just one, without considerably reducing the performance.
+1.1 Reward Approximation
+Now, suppose that the decision-maker has access to a certain subset of actions, $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$, rather than to the full set $\mathcal{A}_{\operatorname*{full}}$.
+For a bandit instance $\theta$, define the regret of being restricted to subset $\mathcal{A}$ as the gap between the best outcome achievable with the full action space and the best outcome within $\mathcal{A}$:
+$\textnormal{Regret}(\theta):=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\max_{a^{\prime}\in\mathcal{A}}\;\mu_{a^{\prime}}(\theta),$} (1)
+Our objective then is to identify a small subset $\mathcal{A}$ that minimizes the expected regret $\mathbb{E}_{\theta\sim\mathcal{P}}\left[\operatorname*{Regret}\right]$, where the expectation is taken over all possible bandit instances.
+This objective may be motivated in several ways: First, as discussed above, in problems such as inventory management it may be of natural interest to find a representative subset of actions, due to the intrinsic difficulties associated with maintaining the full set of actions.
+Another possible reason would be to reduce the exploration phase when encountering a new bandit instance, especially when the time horizon $T$ is small.
+Indeed, the classic Bayesian bandit regret, (Agrawal & Goyal, 2012), generally scales as $\Omega(\sqrt{\left|\mathcal{A}_{\operatorname*{full}}\right|T})$, and may become prohibitively large for large $|\mathcal{A}_{\operatorname*{full}}|$.
+On the other hand, if a subset $\mathcal{A}$ is carefully chosen, the resulting Bayesian bandit regret can be significantly lower.
+To see this, for a given $\mathcal{A}\subset\mathcal{A}_{\operatorname*{full}}$ we can decompose the bandit regret as follows:
+$\textnormal{BayesianBanditRegret}:=\mathbb{E}\sum_{t=1}^{T}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)-\mu_{A_{t}}(\theta)\right]$
+$=\mathbb{E}\sum_{t=1}^{T}\left[\max_{a\in\mathcal{A}}\mu_{a}(\theta)\!-\!\mu_{A_{t}}+\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)\!-\!\max_{a\in\mathcal{A}}\mu_{a}(\theta)\right]$
+$\leq C\sqrt{|\mathcal{A}|\cdot T\log T}+T\cdot\mathbb{E}_{\theta}[\textnormal{Regret}].$ (2)
+[…]
+A further motivation is to reduce the computational burden that persists even in structured bandits; see Remark 4.8 for a discussion of linear and Gaussian process (GP) bandits.
+1.2 Subset Selection Algorithm
+We now introduce the subset selection algorithm itself, Algorithm 1.
+Conceptually, the algorithm is simple: Sample a bandit instance $\theta$ from a distribution $\mathcal{P}$ and solve it using an oracle (or practically, any of the standard bandit algorithms, such as UCB or TS), to find the optimal action $a^{*}=a^{*}(\theta)$.
+Add the action to the set or representation actions $\mathcal{A}$.
+Repeat the process $K$ times.
+In the rest of the paper we show that the action subset generated by this simple algorithm will be diverse enough to appropriately cover $\mathcal{A}_{\operatorname*{full}}$ by $\mathcal{A}$.
+Our bounds will depend on the metric structure of the set $\mathcal{A}_{\operatorname*{full}}$ under the $L_{2}$ metric, and on the number of samples $K$.
+Note that Algorithm 1 does not require the knowledge of the metric, only the ability to sample bandit instances at random, and to solve them.
+Algorithm 1 Epsilon Net Algorithm
+1:Input: Action space $\mathcal{A}_{\operatorname*{full}}$, Sample size $K$.
+2:Output: A subset of actions $\mathcal{A}$.
+[…]
+4:for $1,\dots,K$ do
+5:  Sample a bandit instance $\theta$ from a distribution $\mathcal{P}$ (unknown to the learner).
+6:  Find optimal action $a^{*}(\theta)\!:=\!\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)$.
+7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$.
+8:end for
 Our main contributions are organized as follows:
-  • Meta-Bandits Framework.
-We propose a meta-bandits framework that specifically tackles combinatorial action selection by leveraging correlations across similar actions and bandit instances.
-To the best of our knowledge, this is the first such framework.
-  • $\epsilon$-Net Algorithm.
-We introduce a simple algorithm within this framework.
-It starts with the intuitive idea of placing a grid over the action space, then refines it using an importance-based selection mechanism.
+  • Problem Formalization and An Algorithm Although we believe the subset selection problem is natural and of practical interest, to the best of our knowledge it has not been studied in the literature.
+In this paper we provide the first formalization of the problem, in the framework of the statistical similarities between actions.
+We also introduce a simple but effective subset selection algorithm, that does not require apriori knowledge of the correlations.
   • Regret Analysis.
-We provide theoretical guarantees for both the grid and the algorithm’s output, including upper and lower bounds on expected regret, along with results under general sub-Gaussian processes.
-We also discuss the cost of not using a grid, which depends on the importance-structure of the action space.
+When the action sets exhibits sufficient correlation structure (ex.
+clustering or a low dimensional manifold structure), we provide upper and lower bounds on the expected error of the algorithm, i.e.
+the regret (1) (Section 4).
+The analysis is done under the standard assumption that the reward process $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ is Gaussian, (Srinivas et al., 2009; Garnett, 2023) and shows the basic relations between the induced geometry of the action set and the expected rewards.
+Our arguments can also be applied virtually with no changes to the sub-Gaussian processes, and this extension is carried out in Supplementary I.
+An extension of these arguments to other types of tail behavior may be implemented using standard techniques.
   • Generalization and Empirical Validation.
-We extend the analysis to settings where outcome functions are sampled from a reproducing kernel Hilbert space (RKHS), and empirically compare our algorithm to Thompson Sampling (TS) and Upper Confidence Bound (UCB).
-1.1 Related Works
-Multi-Armed Bandits (lai1985asymptotically, ; auer2002finite, ) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution.
-The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time.
-Popular methods include the UCB (auer2002finite, ), TS (agrawal2012analysis, ), and EXP3 (auer2002nonstochastic, ) for adversarial settings.
-Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few samples as possible (jamieson2014best, ; kaufmann2016complexity, ).
-Popular methods in fixed confidence setting include Action Elimination even2006action ; karnin2013almost , UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum.
-In fixed budget setting, there is Successive halving karnin2013almost , successive reject audibert2010best .
-Stochastic Linear Optimization assumes that the expected outcome of each action depends through the inner product between a context $\theta$ and an action $a\in\mathcal{A}_{\operatorname*{full}}$ auer2002using ; dani2008stochastic ; rusmevichientong2010linearly .
-This line of work assumes that the action feature vectors are known, so the cardinality of the action space does not play a role and may even be infinite.
-soare2014best study the sample complexity of optimal action identification in this setting, which scales linearly with the action space dimension $n$.
-GP Optimization addresses the case where the feature vectors of actions are unknown but a kernel function is available, so the inputs to the kernel are required.
-srinivas2009gaussian model the outcome function as a sample from a GP prior with a kernel function (williams2006gaussian, ), achieving bandit regret that scales as $\sqrt{\log|\mathcal{A}_{\operatorname*{full}}|}$ for finite action space.
-The widespread adoption of this method in bandit settings (valko2013finite, ; li2022gaussian, ), as well as in continuous action spaces (chowdhury2017kernelized, ), highlights the practicality of assuming that action outcomes are correlated.
-This also motivates our extension, where the outcome function of each bandit instance is modeled as a sample from a RKHS.
-Before discussing two seemingly relevant lines of work that select a subset from the action space, we first highlight an advantage of our approach.
-These methods assume a fixed subset cardinality $|\mathcal{A}|=K$, which is often unclear in practice and requires restarting the algorithm when changed.
-In contrast, our algorithm can adapt the subset size on the fly.
-Top-K Action Identification aims to identify the $K$ actions with the highest expected outcomes using as few samples as possible (kalyanakrishnan2012pac, ; gabillon2012best, ; kaufmann2016complexity, ; chen2017nearly, ).
+We extend the analysis to settings where outcome functions lie in a reproducing kernel Hilbert space (RKHS).
+We empirically evaluate the subset selection algorithm on a number of problems (Section 5).
+Although there are no existing algorithms that could serve as direct performance comparison benchmarks, one could naively apply standard bandit algorithms, such as TS and UCB, via the notions of “super-arms”.
+As expected, since these methods are not sensitive to the underlying geometry of the subsets, we show that their performance is much weaker than that of our approach.
+2 Related Work
+For a general overview of multi-armed bandits (MAB), see (Lattimore & Szepesvári, 2020; Slivkins et al., 2019).
+Background on optimal action identification—which addresses the sample complexity of the inner optimization step in Algorithm 1—is provided in Supplementary A, along with regret bounds for structured (linear/GP) bandits.
+While regret in general Bayesian bandits scales with $|\mathcal{A}_{\operatorname*{full}}|$, linear/GP bandits achieve $|\mathcal{A}_{\operatorname*{full}}|$-independent regret that depends on structural complexity (e.g., feature dimension or information gain).
+Their computational complexity, however, typically remains $|\mathcal{A}_{\operatorname*{full}}|$-dependent.
+Further discussion of computational aspects appears in Remark 4.8.
+Before discussing two related approaches that select a fixed-size subset $|\mathcal{A}|=K$–a choice often unclear in practice and requiring algorithm restarts–we highlight a key advantage of our method: it can adapt the subset cardinality on the fly.
+Top-K Action Identification aims to identify the $K$ actions with the highest expected outcomes using as few samples as possible (Kalyanakrishnan et al., 2012; Gabillon et al., 2012; Kaufmann et al., 2016; Chen et al., 2017).
 This line of work assumes that all actions are independent and have distinct expected values, making its methods inapplicable to our framework.
@@
 In this setting, top-$K$ identification would refer to selecting $K$ actions with the highest expected payoffs $\mathbb{E}[\mu_{a}]$.
-In contrast, our framework considers that the expected outcome of each action, averaged over the distribution of bandits, may be the same—i.e., $\mathbb{E}[\mu_{a}]\!=\!c$ for all $a\!\in\!\mathcal{A}_{\operatorname*{full}}$, where $c$ is a constant—so that the entire action space shares the same highest expected payoff.
+In contrast, our framework considers that the expected outcome of each action, averaged over the distribution of bandits, may be the same—i.e., $\mathbb{E}[\mu_{a}]\!=\!c$ for all $a\!\in\!\mathcal{A}_{\operatorname*{full}}$, where $c$ is a constant.
 Further, even if $\mathbb{E}[\mu_{a}]$ varies across actions, ignoring correlations can be fatal in our framework:
-Example 1.
+Example 2.1.
 Consider three actions: $a_{1}=[1,0]$, $a_{2}=[0.9,0.1]$, and $a_{3}=[-0.1,1]$, and suppose bandits are sampled uniformly from $\theta_{1}=[1,0]$ and $\theta_{2}=[0,1]$. Then,
@@
 Our algorithm, if run until it selects two distinct actions, would output $a_{1}$ and $a_{3}$—the true optimal.
-Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super arm $\mathcal{A}$, with $|\mathcal{A}|=K$.
-Popular methods include CUCB chen2016combinatorial , CTS wang2018thompson .
-We argue that this line of work is not directly applicable to our framework, but we include it as a baseline in subsequent empirical evaluations: (1) It assumes that the expected outcome of a super arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition.
-In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions.
+Combinatorial Bandits considers that the decision maker selects $K$ of base arms from $\mathcal{A}_{\operatorname*{full}}$ in each round, forming a super-arm $\mathcal{A}$, with $|\mathcal{A}|=K$.
+Popular methods include CUCB (Chen et al., 2016), CTS (Wang & Chen, 2018).
+We include this line of work as an empirical baseline, though it is not directly applicable to our framework for two reasons: (1) It assumes that the expected outcome of a super-arm depends only on the expected outcomes of its individual base arms, or imposes a stricter monotonicity condition.
+In our case, even though $\mathbb{E}[\mu_{a}]=c$ for all $a\in\mathcal{A}_{\operatorname*{full}}$, super-arm expected outcomes $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ can differ significantly due to correlations among actions.
 (2) It assumes independence across base arms, whereas we explicitly model correlations.
-Ignoring these correlations misses the core challenge—an issue illustrated in Example 1.
-2 Subset Selection Framework
-We consider the problem of selecting a small number of representative actions from a large action space $\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$, where $n\in\mathbb{N}$.
-(This framework applies to the case $n=+\infty$, with the additional assumption $\sum_{i\geq 1}a^{2}_{i}\!<\!\infty$.) The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, and includes a constant $c\in\mathbb{R}$:
-$\mu_{a}:=\langle a,g\rangle+c,\quad\forall\;a\in\mathcal{A}_{\operatorname*{full}},\;g\sim\mathcal{N}(0,\Sigma),$ (2)
-where $\Sigma$ is a positive semi-definite matrix.
-Let $\theta\in\mathbb{R}^{n}$ follows a multivariate normal distribution with zero mean and identity covariance matrix $I$.
-The distribution of $g$ is in fact equivalent to $\Sigma^{1/2}\theta$.
-Now, let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$.
-We have $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$ and
-$\langle a,g\rangle=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
-Therefore, the setting in equation 2 is equivalent to $\mu_{a}:=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}},\;\theta\sim\mathcal{N}(0,I),$ where $\Sigma^{1/2}\mathcal{A}_{\operatorname*{full}}$ denotes the image of $\mathcal{A}_{\operatorname*{full}}$ under the linear transformation $\Sigma^{1/2}$.
-Since the constant $c$ does not affect the regret (as defined in equation 1), we can, without loss of generality, focus on this canonical Gaussian process (vershynin2018high, , Chapter 7) in the remainder:
-$\boxed{\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{\operatorname*{full}},\theta\sim\mathcal{N}(0,I).}$ (3)
-We define the extreme points as those $x\in\mathcal{A}_{\operatorname*{full}}$ for which there do not exist distinct $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ and $\lambda\in(0,1)$ such that $x=\lambda a+(1-\lambda)a^{\prime}$.
-By the extreme point theorem, if we select all extreme points—denoted $\mathcal{A}=\{a_{1},\dots,a_{K}\}$—as representatives of the full action space, the regret is zero.
-This is because any $a\in\mathcal{A}_{\operatorname*{full}}$ can be expressed as a convex combination of the extreme points: $a=\lambda_{1}a_{1}+\dots+\lambda_{K}a_{K}$, where $\lambda_{i}\geq 0$ and $\sum_{i=1}^{K}\lambda_{i}=1$.
-Thus, for any $\theta\in\mathbb{R}^{n}$:
-$\begin{split}\langle a,\theta\rangle=\sum_{i=1}^{K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i=1}^{K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.\end{split}$
-By equation 3, it yields $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$ where the left inequality uses $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$.
-Thus, the two quantities $\max_{a\in\mathcal{A}}\mu_{a}$ and $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ are equal.
-This example highlights how a geometric approach can be used to solve the stochastic combination problem.
-However, even if one only needs the extreme points, the set of extreme points may still be large, e.g, the extreme points of a Euclidean ball is infinite.
-To address this, we will later introduce the notions of $\epsilon$-nets.
-Without loss of generality, we assume $\mathcal{A}_{\operatorname*{full}}$ consists only of the extreme points of $\mathcal{A}_{\operatorname*{full}}$, as they are the only points of interest.
+Ignoring these correlations misses the core challenge—an issue illustrated in Example 2.1.
+3 Subset Selection Algorithm
+In this Section, we introduce the formal problem setup and the technical tools used in the analysis.
+The optimal action in a bandit instance $\theta$ is defined as
+$a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta).$ (3)
+Thus $a^{*}$ is a random variable taking values in the set $\mathcal{A}_{\operatorname*{full}}$.
+We denote by $q$ the induced probability distribution of $a^{*}$ on $\mathcal{A}_{\operatorname*{full}}$.
+Specifically, for any subset $r\subset\mathcal{A}_{\operatorname*{full}}$, we have
+$q(r):=\mathbb{P}\left(a^{*}(\theta)\in r\right)=\mathbb{E}_{\theta\sim\mathcal{P}}\mathbbm{1}_{\{a^{*}(\theta)\in r\}}.$ (4)
+In particular, note that the samples $a^{*}$ in lines 5&65\&6 of Algorithm 1 may be viewed as i.i.d samples from $q$. The distribution $q$ will be named as the importance measure $q$.
+3.1 Gaussian Processes
+As discussed earlier, in our analysis we consider the setting where the reward expectations $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ form a Gaussian process.
+An extension of our main bound to sub Gaussian processes is given in Supplementary I.
+Let $\|\cdot\|_{2}$ denote the standard Euclidean norm on $\mathbb{R}^{n}$, and for a set $S\subset\mathbb{R}^{n}$ define the diameter of $S$ as $\mathrm{diam}(S):=\sup_{a,b\in S}\|a-b\|_{2}$.
+To simplify the notation, we will consider the canonical representation of the Gaussian process: That is, we assume that $\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$ for some dimension $n>0$, and that
+$\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{\operatorname*{full}},\theta\sim\mathcal{N}(0,I).$} (5)
+Here, $\left$$\cdot$$\cdot\right$> is the standard inner product on $\mathbb{R}^{n}$, and $\mathcal{N}(0,I)$ is the standard Gaussian distribution on $\mathbb{R}^{n}$.
+Thus, in terms of the notation in Section 1, the family of bandits $\mathcal{F}$ may be identified as $\mathbb{R}^{n}$, and the distribution $\mathcal{P}$ as the standard Gaussian.
+Clearly, the family of variables $\left\{\mu_{a}\right\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ given by (5) is a Gaussian process.
+Conversely, we note that any finite dimensional Gaussian process can be represented as (5), and any infinite dimensional Gaussian process [^1] Under very mild technical assumptions can be arbitrarily approximated by (5), in a sufficiently high dimension $n$.
+See (Vershynin, 2018, Chapter 7).
+Thus, we may assume the form (5) without loss of generality.
+Moreover, crucially, the dimension $n$ will not be important in our bounds, and can be large.
+On the other hand, however, it may be also of interest to note that when $n$ is fixed, the setup (5) is equivalent to the Bayesian linear bandit setup, (Lattimore & Szepesvári, 2020) (although we solve a different problem).
+The relation with linear bandits will be discussed further in Remark 4.8.
+We define the $L_{2}$ distance induced on the $\mathcal{A}_{\operatorname*{full}}$ by the process to be the distance
+$\left\|\mu_{a}-\mu_{a^{\prime}}\right\|^{2}_{L_{2}}:=\mathbb{E}_{\theta}\left|\left<a,\theta\right>-\left<a^{\prime},\theta\right>\right|^{2}.$ (6)
+Note that under the definition (5), the $L_{2}$ distance and the standard Euclidean distance in $\mathbb{R}^{n}$ coincide, $\left\|\mu_{a}-\mu_{a^{\prime}}\right\|^{2}_{L_{2}}=\left\|a-a^{\prime}\right\|_{2}^{2}$ for all $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$.
 3.2 Epsilon Nets
-If $\mathcal{A}_{\operatorname*{full}}$ is very large, a natural approach is to construct a grid over the action space, where the grid points serve as representative actions.
-This ensures that for every action in the full space, there exists a representative that is close to it.
-This idea is formally captured by the notion of a (geometric) $\epsilon$-net.
-To proceed, we clarify what we mean by an $\epsilon$-net, as there are at least two definitions: one from a geometric perspective (vershynin2018high, , Chapter 4) and another from a measure-theoretic perspective (matousek2013lectures, , Chapter 10).
-Let $\|\cdot\|_{2}$ denote the Euclidean norm.
-Define the diameter of a compact set $r\in\mathbb{R}^{n}$ as $\mathrm{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$.
+There will be two senses in which we would be interested in approximating $\mathcal{A}_{\operatorname*{full}}$: First, for every $a\in\mathcal{A}_{\operatorname*{full}}$ we would like to have $a^{\prime}\in\mathcal{A}$ such that $a^{\prime}$ is close to $a$ in the distance (6) (or, equivalently, in $\left\|\cdot\right\|_{2}$).
+Second, in order for the subset selection algorithm to able to discover the points $a^{\prime}$, we need an additional, probabilistic notion of approximation.
+The following two notions of an $\epsilon$-net, the geometric and the measure theoretic nets, formalize these approximations.
+We refer to (Vershynin, 2018, Chapter 4) for additional details on geometric nets, and to (Matousek, 2013, Chapter 10) for measure theoretic nets.
+Definition 3.1.
 A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a geometric $\epsilon$-net if, for all $a\in\mathcal{A}_{\operatorname*{full}}$, there exists $a^{\prime}\in\mathcal{A}$ such that
 $\|a-a^{\prime}\|_{2}<\epsilon.$
-  • Let $\mathcal{R}$ be a finite partition of the extreme points into disjoint clusters such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{\operatorname*{full}}$.
-Given a measure $q$ assigning a value to each cluster $r\in\mathcal{R}$.
-A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a Measure-Theoretic $\epsilon$-net with respect to measure $q$ if, for any cluster $r\in\mathcal{R}$, we have:
+As discussed earlier, geometric $\epsilon$-nets will be used as approximations of the set $\mathcal{A}_{\operatorname*{full}}$.
+While we’ll assume that good approximating nets exist in the data, these nets will never need to be explicitly constructed.
+Instead, part of such representative points will be implicitly chosen by the subset selection algorithm.
+The $\epsilon$-covering number of the set $\mathcal{A}_{\operatorname*{full}}$, denoted $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$, is the size of the smallest $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$:
+[…]
+Now, Let $\mathcal{R}$ be a finite partition of $\mathcal{A}_{\operatorname*{full}}$ into disjoint clusters $r$, such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{\operatorname*{full}}$.
+Let $q$ be a probability distribution on $\mathcal{A}_{\operatorname*{full}}$, which therefore assigns a weight $q(r)$ to each cluster $r\in\mathcal{R}$.
+Definition 3.2.
+A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a measure-theoretic $\epsilon$-net with respect to measure $q$ and partition $\mathcal{R}$, if, for any cluster $r\in\mathcal{R}$, we have:
 $r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
-A geometric $\epsilon$-net ensures small regret because if two actions $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ are close in the Euclidean sense, then the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., their expected squared difference is small): $\|\mu_{a}\!-\!\mu_{a^{\prime}}\|_{L^{2}}=\!\left(\mathbb{E}(a\!-\!a^{\prime})^{\top}\theta\theta^{\top}(a\!-\!a^{\prime})\right)^{1/2}\!=\|a\!-\!a^{\prime}\|_{2},$ where $\theta^{\top}$ denotes the transpose of $\theta$.
-The equalities use equation 3 and $\mathbb{E}\theta\theta^{\top}=I$.
-Therefore, by definition, a geometric $\epsilon$-net guarantees the existence of an action $a\in\mathcal{A}$ whose expected outcome $\mu_{a}$ is close to that of the optimal action for any given bandit instance.
-However, this net suffers from the curse of dimension: e.g., for $0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$.
-The measure-theoretic $\epsilon$-net addresses this issue.
-Put simply, the measure-theoretic $\epsilon$-net restricts the grid construction to only the most important clusters $r\in\mathcal{R}$, as determined by the $q$-measure.
-2.2 Epsilon Net Algorithm
-We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by haussler1986epsilon .
-It selects $K$ i.i.d.
-random actions, aligned with the distribution of bandit instances.
-Since repetitions are allowed, the resulting subset $\mathcal{A}$ may have fewer than $K$ distinct actions.
-Algorithm 1 Epsilon Net Algorithm
-1:Input: Action space $\mathcal{A}_{\operatorname*{full}}$, Sample size $K$.
-2:Output: A subset of actions $\mathcal{A}$.
-[…]
-4:for $1,\dots,K$ do
-5:  Sample a bandit instance $\theta$.
-6:  Find optimal action $a^{*}(\theta)\!:=\!\operatorname*{arg\,max}_{a\in\mathcal{A}_{\operatorname*{full}}}\langle a,\theta\rangle$.
-7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$ $\triangleright$ Repetition of actions is allowed
-8:end for
-We define the optimal action in a bandit instance $\theta$ as
-$a^{*}(\theta):=\operatorname*{arg\,max}_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta).$
-Assumption 2 (Unique optimal action).
-The optimal action $a^{*}(\theta)$ is unique with probability $1$ over all bandit instances.
-Define the Importance Measure $q$ over a partition $\mathcal{R}$:
-$\boxed{q(r):=\Pr[a^{*}(\theta)\in r]=\!\int\!\mathbb{1}\{a^{*}(\theta)\in r\}\;p(\theta)d\theta,}$ (4)
-where $p(\theta)$ is the density of $\theta$ and $\int p(\theta)d\theta=1$.
-Under Assumption 2, measure $q$ is a probability distribution.
-It reflects the probability that a given cluster contains the optimal action and thus represents the potential contribution of that cluster to the expected regret.
-Assumption 3.
-The support of measure $q$ is compact.
-The compactness assumption ensures that the term $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\right]$ is finite and guarantees the attainment of a unique optimal action.
-Without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}$ is the support of the measure $q$.
-By definition, Algorithm 1 samples $K$ i.i.d.
-extreme points from clusters in $\mathcal{R}$ according to measure $q$.
-If a cluster $r\in\mathcal{R}$ has a higher measure $q(r)$, its elements are more likely to be included in the output.
-In fact, with high probability, this algorithm outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$ with respect to measure $q$.
-(This is a simplified version of Theorem 10.2.4 of matousek2013lectures .)
-Lemma 4.
-Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$.
+The measure theoretic $\epsilon$-net ensures that in a given partition $\mathcal{R}$, either $\mathcal{A}$ intersects a cluster $r$, or the cluster $r$ is small in terms of the distribution $q$.
+Roughly speaking, this will be used to show that if the subset selection algorithm doesn’t find a cluster $r$ of actions, then the cluster’s contribution to the total expected regret is small and can be ignored.
+An important property of random sampling is that with high probability it will produce a measure theoretic $\epsilon$-net for any fixed partition of the support of the distribution.
+Specifically, we have the following result:
+Lemma 3.3.
+Fix $\epsilon>0$.
+Let $\mathcal{R}$ be a partition of $\mathcal{A}_{\operatorname*{full}}$, and let $q$ be an importance measure given by (4).
 Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ samples.
@@
 $r\cap\mathcal{A}\neq\emptyset\quad\text{whenever}\quad q(r)>\epsilon.$
-The partition $\mathcal{R}$ bridges the two definitions of $\epsilon$-net: choosing one point from each cluster gives an $\epsilon$-net in both senses, though with different values of $\epsilon$.
-Geometrically, $\epsilon$ is the largest cluster diameter $\epsilon:=\max_{r\in\mathcal{R}}\mathrm{diam}(r)$; measure-theoretically, $\epsilon$ is the smallest cluster measure $\epsilon:=\min_{r\in\mathcal{R}}q(r)$.
+This result is a special case of Theorem 10.2.4 in (Matousek, 2013), which in turn follows the analysis in (Haussler & Welzl, 1986).
+The algorithm in (Haussler & Welzl, 1986) is called the $\epsilon$-net Algorithm, and we have retained that name.
+The full proof of the Lemma is given in Supplementary D.
 4 Regret Analysis
-In this section, we begin by analyzing a special class of geometric $\epsilon$-nets, constructed by partitioning the action space into clusters and selecting a single representative action from each cluster.
-We then extend the analysis to obtain algorithm-dependent bounds for the output of Algorithm 1.
-Definition 1 (Reference subsets).
-Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
-A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$ and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
-Only for lower bounds, we assume a well-separated partition: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it.
-Assumption 5.
+In this Section we introduce and discuss the regret bounds for subset selection Algorithm 1.
+Our main result is Theorem 4.5, which bounds the regret in terms of the covering numbers of the set $\mathcal{A}_{\operatorname*{full}}$.
+Typically, sets which exhibit clustering, or sets with low dimensional manifold structure, would have low covering numbers, and we expect such sets to commonly occur in practical problems (the manifold hypothesis, (Bengio et al., 2013)).
+Roughly speaking, suppose that the set $\mathcal{A}_{\operatorname*{full}}$ can be represented as a union of a few small clusters, as illustrated in Figure 1.
+The main term in regret bound bound in Theorem 4.5 is related to the complexity of each individual cluster, and typically would controlled by its diameter $\epsilon$.
+Additional terms involve a logarithmic dependence on the total number of clusters (times $\epsilon$), and an error term which accounts for the probability of Algorithm 1 to find the relevant clusters.
+This term decays exponentially with $K$.
+Figure 1: Illustration of Theorem 4.5.
+The full action space (blue dots) lies on the unit sphere in $\mathbb{R}^{3}$ and can be partitioned into three clusters.
+The regret bound established in the theorem depends not on the diameter of the entire action space—indicated by the red dashed line—but rather on the diameters of the individual clusters, represented by the green dashed lines, each of which is significantly smaller.
+Perhaps the most central technical notion in our analysis is the notion of partition $\mathcal{R}=\left\{r_{\ell}\right\}_{\ell\leq m}$, such that $\mathcal{A}_{\operatorname*{full}}=\cup_{l\leq m}r_{\ell}$, and we will typically be interested in partitions where each subset $r_{\ell}$ also have a relatively small diameter $\epsilon$.
+There are two main ways in which such partitions may be present in $\mathcal{A}_{\operatorname*{full}}$.
+First, some realistic datasets may be naturally clustered, as depicted in Figure 1.
+Typically, in such situations there will be a characteristic scale $\epsilon$, such that there are few clusters with diameter $\epsilon$, which are well separated.
+However, inside each cluster, on scales smaller then $\epsilon$, there may be no additional structure, i.e.
+the clusters themselves would look dense and high dimensional.
+In such situations, we still may benefit from a regret of order $\epsilon$.
+The second scenario is when the whole set $\mathcal{A}_{\operatorname*{full}}$ is a low dimensional manifold, typically in a high dimensional space.
+In that case, for any scale $\epsilon$ we construct relatively small $\epsilon$-nets, which can be used to derive appropriate partitions.
+Section 4.1 introduces bounds on the regret in terms of the partitions and their diameters.
+These are the bounds that one could obtain for a fixed a-priori known partitions.
+In Section 4.2 we obtain bounds for the subset selected by Algorithm 1, which finds (an approximation of) the reference points for the best partition available in $\mathcal{A}_{\operatorname*{full}}$, without knowing this partition.
+Theorem 4.5 provides bounds in terms of a general partition given scale $\epsilon$, while in Theorem 4.9, by taking $\epsilon$ to 0 we provide worst case diminishing bounds.
+4.1 Partitions and Reference Set Bounds
+For a point $a\in\mathbb{R}^{n}$ and scalar $\epsilon\geq 0$ denote by $B(a,\epsilon):=\left\{a^{\prime}\in\mathbb{R}^{n}\hskip 5.69054pt|\hskip 5.69054pt\left\|a-a^{\prime}\right\|_{2}\leq\epsilon\right\}$ the closed Euclidean ball of radius $\epsilon$ around $a$.
+Definition 4.1 ($\epsilon$-reference sets).
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of $\mathcal{A}_{\operatorname*{full}}$.
+For an $\epsilon>0$, a set $\mathcal{A}^{\epsilon}:=\{a_{1},\dots,a_{m}\}$ is an $\epsilon$-reference set for $\mathcal{R}$ if for every $\ell\leq m$, $r_{\ell}\subseteq B(a_{\ell},\epsilon)$.
+The point $a_{\ell}$ is called the reference point of the subset $r_{\ell}$.
+Note that the $\epsilon$-reference set is in particular a (geometric) $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$.
+However, for our purposes it will be more convenient to discuss partitions, which we do with the notion of reference sets.
+Note also that given a partition $\mathcal{R}$, one can always find a reference set with $\epsilon=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
+Theorem 4.2 (Regret bounds of $\epsilon$-reference subsets).
+Let $\mathcal{R}$ be a partition of $\mathcal{A}_{\operatorname*{full}}$ and let $\mathcal{A}^{\epsilon}$ be an $\epsilon$-reference subset for $\mathcal{R}$.
+There is an absolute constant $C>0$, such that
+$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\left(\mu_{a}-\mu_{a_{\ell}}\right)\right]+C\epsilon\sqrt{\log m}.$
+Proof sketch.
+For each cluster $\ell\leq m$, define a simple GP $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$.
+Define a non-negative random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$.
+When $a^{*}(\theta)\in r_{\ell}$, regret is upper bounded by $Y_{\ell}$.
+Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$.
+Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded via appropriate concentration arguments.
+$\square$
+For the lower bounds, we’ll require an additional separation assumption: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it.
+Assumption 4.3.
 For any $r\in\mathcal{R}$, whenever the optimal action lies in cluster $r$, i.e., $a^{*}(\theta)\in r$, then $\mu_{a}\geq\max_{a^{\prime}\in\mathcal{A}_{\operatorname*{full}}\setminus r}\mu_{a^{\prime}},\;\forall a\in r$.
-Theorem 6 (Regret bounds of reference subsets).
-Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$.
-Then, there is an absolute constant $C>0$, such that
-$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
-If the partition $\mathcal{R}$ satisfies Assumption 5, then
-$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
-Proof sketch.
-For each cluster $\ell\leq m$, define a simple Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$.
-Define a non-negative random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$.
-When $a^{*}(\theta)\in r_{\ell}$, regret is upper bounded by $Y_{\ell}$ (or equal to it under Assumption 5).
-Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$.
-Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded via concentration property of Gaussian process.
-$\square$
-For each $\ell\leq m$, the term $\mathbb{E}_{\theta}[\max_{a\in r_{\ell}}\mu_{a}]$ equals $\mathbb{E}_{\theta}[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}]$, since $\mathbb{E}_{\theta}[\mu_{a}]=0$.
-3.1 Regret Bounds of Algorithm
-The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds.
-The key difference is that whereas the reference subset includes a representative from each cluster, the algorithm may miss some clusters.
-However, the algorithm still achieves regret comparable to that of the reference subset, as it tends to miss clusters that contribute minimally to the expected regret.
-The expected regret in previous results is taken over bandit instances $\theta$.
-In contrast, since the output of Algorithm 1 is random, the expected regret analyzed in this section is taken with respect to both the algorithm’s randomness (i.e., the sampled $\mathcal{A}$) and the distribution over $\theta$.
-Theorem 7.
-Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
+Theorem 4.4.
+If the partition $\mathcal{R}$ satisfies Assumption 4.3, then
+$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\left(\mu_{a}-\mu_{a_{\ell}}\right)\right]-C\epsilon\sqrt{\log m}.$
+4.2 Regret Bounds for the Algorithm
+As discussed above, Theorem 4.5 provides regret bounds for the regret of the set $\mathcal{A}$ selected by Algorithm 1.
+To get an intuition about the argument, suppose that $\mathcal{A}_{\operatorname*{full}}$ admits a good partition $\mathcal{R}$ (i.e.
+relatively few clusters, of small diameter $\epsilon$).
+We then show that for any fixed partition $\mathcal{R}$, samples in Algorithm 1 will hit most of the clusters of that partition, and can consequently be used as reference set points in Theorem 4.2.
+Indeed, by Lemma 3.3, either Algorithm 1 will hit the cluster, or the probability of that cluster under $q$ is small, in which case its contribution to regret will be small too.
+This gives rise to the additional error term in Theorem 4.5 compared to Theorem 4.2, but this term decays exponentially with $K$.
+Theorem 4.5.
+Let $\mathcal{R}$ be a partition of $\mathcal{A}_{\operatorname*{full}}$ and set $\epsilon=\max_{r\in\mathcal{R}}diam(r)$. Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ iterations. Then
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\left(\mu_{a}-\mu_{a_{\ell}}\right)\right]+C\epsilon\sqrt{\log m}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},\end{split}$
+where $C>0$ is an absolute constant.
+Note that the expectation $\mathbb{E}_{\theta,\mathcal{A}}$ in this result is taken with respect to both the algorithm’s randomness (i.e.
+the sampled set $\mathcal{A}$) and the distribution over $\theta$.
+In the following remarks we discuss the different error terms appearing in Theorem 4.5.
+Remark 4.6 (The Cluster Complexity Term).
+The term $\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right]$ in Theorem 4.5 measures the size of the cluster, and is also known as the Gaussian width of $r_{\ell}$ (see (Vershynin, 2018), Chapter 7).
+There are two standard ways in which we can bound this term: First, for any finite $r_{\ell}$ we have
+$\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right]\leq\frac{\operatorname{diam}(r_{\ell})}{2}\sqrt{\log|r_{\ell}|}\leq\frac{\epsilon}{2}\sqrt{\log\left|\mathcal{A}_{\operatorname*{full}}\right|},$ (8)
+thus yielding a scaling with the diameter and a square-root-logarithmic dependence on the size of the actions set. See Supplementary K additional details on (8). In addition, we have the bound
+$\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right]\leq\operatorname{diam}(r_{\ell})\frac{\sqrt{n}}{2},$ (9)
+where $n$ is the ambient dimension of the action set, i.e.
+$\mathcal{A}_{\operatorname*{full}}\subset\mathbb{R}^{n}$ (see (Vershynin, 2018), Chapter 7).
+This bound is independent of the cardinality of the action set and may be applied if $n$ is known to be small.
+However, as discussed in Section 3.1, we do not assume that $n$ is small.
+We use both bounds in Theorem 4.9 below.
+Remark 4.7 (Sampling Correction Term).
+The sampling correction term $\mathbb{E}_{q}\big[(1-q(r))^{2K}\big]$ in Theorem 4.5 decays exponentially with $K$. Here we analyze this decay in more detail. Write
+$\mathbb{E}_{q}\big[(1-q(r))^{2K}\big]$ $=\sum_{r}q(r)(1-q(r))^{2K}$
+$\leq\sum_{r}q(r)e^{-2Kq(r)},$
+where we have used $1-x\leq e^{-x}$ for $x\in[0,1]$.
+For each cluster $r\in\mathcal{R}$, the summand $q(r)e^{-2Kq(r)}$ is small when $q(r)$ is small and decays rapidly when $q(r)$ is large.
+For any threshold $\tau\in(0,1)$, we split the sum into contributions from “large” and “small” probabilities:
+$\sum_{r}q(r)e^{-2Kq(r)}\leq e^{-2K\tau}+\sum_{r:q(r)\leq\tau}q(r).$
+Thus, to ensure $\mathbb{E}_{q}[(1-q(r))^{2K}]\leq\varepsilon$, it suffices to choose $\tau$ such that $\sum_{r:q(r)\leq\tau}q(r)\leq\varepsilon/2$ and set
+$K\geq\frac{1}{2\tau}\log\!\left(\frac{2}{\varepsilon}\right).$
+Remark 4.8 (Linear Bandits and Known Gaussian Process Regret Bounds).
+As discussed in Section 3.1, when dimension $n$ is fixed, we can view the problem setup (5) as that of Bayesian linear bandits.
+In Section 1.1 we have mentioned two motivations for the action subset selections problem: The intrinsic interest in reducing the size of the action set for purposes such inventory maintenance, and a possibility of faster learning in a new bandit instances, when the action set is reduced from $\mathcal{A}_{\operatorname*{full}}$ to $\mathcal{A}$, see (2).
+While the first scenario holds for linear bandits as well, the second scenario may be weaker, since in linear bandits the regret may be bounded either as $O(\sqrt{n\cdot\log\left|\mathcal{A}_{\operatorname*{full}}\right|}\sqrt{T})$, (Russo & Van Roy, 2016), or as $O(n\sqrt{T})$, (Dani et al., 2008).
+Thus, in these situations the regret depends at most only logarithmically on the size of $\left|\mathcal{A}_{\operatorname*{full}}\right|$, if the dimension $n$ is known to be small.
+However, as discussed earlier, we do not require the dimension to be small, and our bounds depend on the actual geometry of the set instead of relying on global constraints such as the dimension.
+Thus our bounds may remain small even when $n$ is arbitrarily large.
+Similar considerations also apply more generally to existing regret bounds for Bayesian bandits with Gaussian Process reward distributions (a problem also known as Bayesian Optimization), (Srinivas et al., 2009; Garnett, 2023).
+While these bounds are applicable with infinite $\mathcal{A}_{\operatorname*{full}}$, they involve global parameters of the model and assumptions on the kernel.
+For instance, for linear kernels the bound is $O(n\sqrt{T})$, while for RBF kernels one has $O(\sqrt{T\left(\log T\right)^{n+1}})$ - exponential in the underlying dimension, (Srinivas et al., 2009).
+Thus such bounds do not reflect the underlying clustering or non-linear low dimensionality of the data.
+Even for small $n$, each round incurs non-trivial optimization costs.
+Both UCB and TS require solving an inner maximization over $\mathcal{A}_{\text{full}}$, whose difficulty depends on the action space structure.
+In linear bandits, UCB solves $\max_{a\in\mathcal{A}_{\text{full}}}\max_{\theta\in\mathcal{C}_{t}}\langle\theta,a\rangle$, which is NP-hard in general—even for convex $\mathcal{A}_{\text{full}}$ (Zhang et al., 2025)—but trivial for small finite sets.
+TS similarly requires $\max_{a\in\mathcal{A}_{\text{full}}}\langle\tilde{\theta}_{t},a\rangle$; while linear in $a$, this is costly for large finite sets, though near-sublinear costs $\tilde{O}(|\mathcal{A}_{\text{full}}|^{1-o(1)})$ are achievable under structural assumptions (Jun et al., 2017; Yang et al., 2022).
+For GP bandits, the inner problem maximizes a non-concave acquisition function (e.g., GP-UCB), which is generally intractable for continuous domains.
+With finite $\mathcal{A}$ of size $|\mathcal{A}|$, the per-step cost is $\tilde{O}(|\mathcal{A}|h^{2}+h^{3})$ for batch size $h$ (Calandriello et al., 2022).
+Continuous domains typically require discretizations for approximate maximization (Chowdhury & Gopalan, 2017; Salgia et al., 2021).
+Figure 2: Comparison on solving (11) by selecting $K=5$ actions from 15 grid points in $[0,2]$, with outcome functions $f(a)=\mu_{a}$ sampled via an RBF kernel at varying length-scales.
+Left: Expected regret over 50 repetitions: our method with exhaustive search (green) vs.
+super-arm TS (yellow), UCB (orange), and Successive Halving (SH; blue).
+TS/UCB run for 3,000 rounds; SH uses a budget of 37,000 pulls.
+Right: SH’s expected regret (blue) and cumulative super-arm pulls (yellow bars) per round vs.
+our method (green).
+SH requires nearly 10,000 pulls—about three times the number of super-arms—to match our performance.
+Finally, as discussed earlier, while Theorem 4.5 is stated for a fixed partition, for any $\epsilon>0$ we can construct a partition derived from an $\epsilon$-net, and to obtain a corresponding regret bound.
+We carry this construction out in the following result.
+Recall the covering number $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$ from (7).
+Theorem 4.9.
+Set $M:=\operatorname{diam}(\mathcal{A}_{\operatorname*{full}})$ and fix $\epsilon>0$. Let $\mathcal{A}$ be the output of Algorithm 1. Then, for
+$K\geq c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon),$ (10)
+we have
+$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;$ $C\epsilon\min\{\sqrt{n},\sqrt{\log|\mathcal{A}_{\operatorname*{full}}|}\}$
+$+C^{\prime}\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},$
+where $C,C^{\prime},c>0$ are absolute constants.
+In particular, by taking $\epsilon\to 0^{+}$, and taking $K$ to satisfy (10), we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$.
+A matching lower bound and an extension to the sub-Gaussian case are provided in Supplementary B, along with the full proofs of the results in this section.
+5 Generalization and Empirical Validation
+We now treat the expected rewards $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as evaluations of an unknown outcome function $f:\mathcal{A}_{\operatorname*{full}}\to\mathbb{R}$ over the abstract action set, where $f(a):=\mu_{a}$.
+Although our framework is general, we instantiate it in the experiments using outcome functions from an RKHS (Williams & Rasmussen, 2006), a function class for which our method is applicable and theoretically sound.
+The synthetic outcome functions are sampled from a Gaussian process with an RBF kernel (with length-scale parameter $\ell$), yielding outcome functions consistent with an underlying RKHS (Kanagawa et al., 2018).
+The action space corresponds to the input domain of the RBF kernel, while the kernel itself is not revealed to the learner.
+Details justifying the applicability of our framework to RKHS, as well as the data generation procedure, are provided in Supplementary C.
+Returning to the objective stated in the Introduction, we seek a subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ of cardinality $K$ that minimizes the expected regret defined in (1).
+Since $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is constant with respect to $\mathcal{A}$, this is equivalent to solving:
+$\max_{\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}}\mathbb{E}\!\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (11)
+This problem is generally non-trivial: the objective involves an expectation over random variables with an unspecified (and potentially correlated) distribution, and the inner $\max$ operator induces non-smooth dependence on $\mathcal{A}$.
+Nevertheless, our method efficiently circumvents the combinatorial complexity inherent in (11).
+Figure 3: Comparison of EpsilonNet+TS (green), CTS (brown), and CUCB (red) on solving (11) with $K=10$ actions from 500 grid points in [5,5][-5,5], using outcome functions $f(a)=\mu_{a}$ sampled via an RBF kernel at varying length-scales.
+In EpsilonNet+TS, the argmax oracle in Algorithm 1 is replaced by TS; CTS and CUCB run for 3,000 rounds.
+All methods use a $\mathcal{N}(0,1)$ prior.
+Left: expected regret (30 repetitions).
+Right: runtime (purple, left axis) and arm pulls (gray, right axis).
+5.1 Comparisons at the Super-Arm Level
+We compare our method against TS and UCB in terms of super-arm identification performance: we aim to identify near-optimal subsets of (11) by selecting $K=5$ actions from an action space consisting of 15 grid points in $[0,2]$.
+Each 5-action tuple is treated as a super-arm.
+This yields $\binom{15}{5}=3003$ super-arms—a moderate number that allows us to approximate the optimal super-arm using Successive Halving (SH), although SH becomes computationally infeasible for larger action spaces.
+Outcome functions are sampled from via an RBF kernel at length scales $l\in\{0.5,1,\dots,4\}$.
+In our approach, we run Algorithm 1 until $K=5$ distinct actions are selected, using exhaustive search over the action space to compute the optimal action $a^{*}(\theta)$ for each bandit instance $\theta$.
+TS and UCB operate in a bandit feedback setting: at each round (i.e., for a new $\theta$), a super-arm $\mathcal{A}$ is selected, the payoff $\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ is observed, and the policy is updated.
+We run both methods for 3,000 rounds—matching the total number of super-arms ($\binom{15}{5}=3003$).
+The payoff of a super-arm $\mathcal{A}$ is defined as $\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ to align with the objective in (11); thus, these methods directly optimize $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$.
+Since this payoff is unbounded, we use a $\mathcal{N}(0,1)$ prior for both TS and UCB.
+For SH, we allocate a fixed budget of $N\log_{2}N\approx 37{,}000$ arm pulls (following (Karnin et al., 2013)), where $N=3003$ is the number of super-arms.
+In each round (corresponding to a new $\theta$), the set of surviving super-arms is halved based on empirical payoffs, continuing until one remains or the budget is exhausted.
+Finally, we estimate the expected regret of all methods over $10^{5}$ additional sampled outcome functions.
+The left subplot of Figure 2 shows the expected regret of our method (green), TS (yellow), UCB (orange), and SH (blue).
+Solid curves and shaded regions denote the mean $\pm$ one standard deviation over $50$ repetitions.
+As observed, the expected regret decreases with increasing length-scale $l$.
+The length-scale $l$ controls the effective number of independent actions: for $|a-a^{\prime}|\gg l$, $\mu_{a}$ and $\mu_{a^{\prime}}$ become nearly uncorrelated.
+In the limit of large $l$, the outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\text{full}}}$ behave like i.i.d.
+$\mathcal{N}(0,1)$ variables.
+The right subplot of Figure 2 highlights SH’s impracticality.
+We ran both our method (green) and SH (blue) for 50 repetitions with fixed $l=1$, tracking the best super-arm found and total super-arm pulls after each SH round.
+The blue curves (with error shades) show SH’s expected regret (mean $\pm$ std.), while the yellow bars (right y-axis) indicate cumulative super-arm pulls per round.
+SH matches our performance only after three rounds—but at roughly three times the total number of super-arms.
+5.2 Oracle-Free Comparisons with CTS and CUCB
+While Algorithm 1 relies on an oracle argmax step, this is often infeasible in practice.
+We replace it with a simple TS rule, yielding a fully implementable, oracle-free variant (EpsilonNet+TS).
+We compare this method against combinatorial TS (CTS) and combinatorial UCB (CUCB)—two practical approaches for choosing an action subset with $|\mathcal{A}|=K$—under the same setup as in Section 5.1, but now selecting $K=10$ actions from an action space consisting of 500 grid points over [5,5][-5,5].
+Our approach runs Algorithm 1 until 10 distinct actions are identified.
+The oracle argmax step is approximated by running TS for 300 rounds; the action chosen in the final round serves as an approximation of $a^{*}(\theta)$.
+CTS and CUCB operate in a semi-bandit setting for 3000 rounds.
+In each round $t$, an instance $\theta_{t}$ is sampled, an action subset $\mathcal{A}_{t}$ of cardinality $K=10$ is selected, and rewards $\mu_{a}(\theta_{t})$ are observed for all $a\in\mathcal{A}_{t}$.
+The subset selected in the final round is taken as the algorithm’s output.
+All methods use a $\mathcal{N}(0,1)$ prior, consistent with the ground truth.
+The left subplot of Figure 3, the left subplot shows the expected regret (mean $\pm$ standard deviation over 30 repetitions) for EpsilonNet+TS (green), CTS (brown), and CUCB (red).
+As observed, the poor performance of CTS and CUCB aligns with discussions in related work.
+The right subplot reports runtime and arm pulls across 30 repetitions with fixed $l=1$: the left axis (purple) shows average runtime $\pm$ one standard deviation, while the right axis (gray) shows total base arm pulls.
+EpsilonNet+TS uses exactly 3,000 arm pulls (300 TS rounds each $\times$ 10 base arms).
+CTS incurs comparable runtime but requires far more pulls (3,000 CTS rounds $\times$ 10 base arm pulls).
+CUCB is significantly slower and incurs additional pulls due to its mandatory initialization.
+6 Conclusions and Future Work
+We proposed a framework for selecting a small representative subset of actions from a very large shared action space in a family of bandit problems.
+Our simple subset-selection algorithm exploits reward correlations to identify representative actions with controllable performance loss.
+Natural directions for future work include developing data-dependent stopping rules for choosing the subset size K, and extending the setting from bandits to Markov decision processes to capture sequential decision making.
+Supplementary Summary: The appendix is divided into four parts.
+First, Sections AC provide additional related work, the regret bound analysis for Algorithm 1, and experimental details omitted from the main text due to space constraints.
+Second, Sections D to I contain the proofs of all the theorems and lemmas stated both in the main body and in Section B.
+Third, Sections J-K present auxiliary tools used throughout the paper.
+Fourth, Sections L-N include additional examples and a numerical study of Algorithm 1.
+This part features a low-dimensional ($n=3$) action space on the unit sphere to investigate the effect of cluster diameter on expected regret, in line with Theorem 4.2.
+It also studies two extensions of the varying dependence structure of actions introduced in the main body:
+  • Nonstationary dependence over the action space: Using a nonstationary Gibbs kernel, where the kernel value depends not only on the difference between two inputs $a,a^{\prime}$ but also on the location of $a$.
+  • Limiting case: The collection of expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ becomes i.i.d.
+LLM Usage: Large language models are used for polishing writing, and finding relevant research.
+Appendix A Extra Related Works
+Multi-Armed Bandits (Lai & Robbins, 1985; Auer et al., 2002a) is defined by a set of actions (arms), each deliver outcomes that are independently drawn from a fixed and unknown distribution.
+The decision-maker sequentially selects an action, observes its outcome, and aims to maximize cumulative outcomes over time.
+The policy is to choose the next action based on the sequence of past plays and outcomes.
+Popular methods include the UCB (Auer et al., 2002a), TS (Agrawal & Goyal, 2012), and EXP3 (Auer et al., 2002b) for adversarial settings.
+Optimal Action Identification focuses on identifying the action with the highest expected outcome in a MAB setting using as few samples as possible (Jamieson & Nowak, 2014; Kaufmann et al., 2016).
+Popular methods in fixed confidence setting include Action Elimination (Even-Dar et al., 2006; Karnin et al., 2013), UCB, and LUCB, all of which achieve sample complexity within a $\log(|\mathcal{A}_{\operatorname*{full}}|)$ factor of the optimum.
+In fixed budget setting, there is Successive Halving (Karnin et al., 2013), Successive Reject (Audibert & Bubeck, 2010).
+Linear Bandits assumes that the expected outcome of each action depends linearly on the inner product between an unknown parameter vector $\theta$ and the feature representation of an action $a\in\mathcal{A}_{\operatorname*{full}}$ (Auer, 2002; Dani et al., 2008; Rusmevichientong & Tsitsiklis, 2010).
+This line of work assumes that action feature vectors are known a priori, so the cardinality of the action space is not a limiting factor and may even be infinite.
+The bound scales as $O(n\sqrt{T})$, where $n$ is the dimension of the feature space.
+GP Bandits address the case where explicit feature representations are unavailable, but outcomes are assumed to be correlated according to a known kernel function $k(\cdot,\cdot)$ (Williams & Rasmussen, 2006).
+Srinivas et al.
+(2009) model the reward function as a sample from a Gaussian process prior and propose the GP-UCB algorithm, which achieves cumulative regret $O\big(\sqrt{T\,\gamma_{T}\log|\mathcal{A}_{\operatorname*{full}}|}\big)$ for finite action spaces, where $\gamma_{T}$ denotes the maximum information gain after $T$ rounds.
+The widespread adoption of this approach—both for finite action sets (Valko et al., 2013; Li & Scarlett, 2022) and continuous domains (Chowdhury & Gopalan, 2017) highlights the practicality of assuming that action outcomes are correlated.
+Epsilon Nets have two standard definitions.
+The first, geometric definition (Vershynin, 2018), requires that radius-$\epsilon$ balls centered at net points cover the set.
+It relates to the covering number and extends to function classes, as in Russo & Van Roy (2013).
+The second, measure-theoretic definition (Matousek, 2013), requires the net to intersect all subsets of sufficiently large measure.
+The classic $\epsilon$-net algorithm by Haussler & Welzl (1986) remains the simplest and most broadly applicable method.
+Later works aim to reduce net size (Pach & Tardos, 2011; Rabani & Shpilka, 2009; Mustafa, 2019) and address online settings (Bhore et al., 2024).
+Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$.
+It is an important topic in high-dimensional probability (Vershynin, 2018) and asymptotic convex geometry (Rothvoss, 2021, Chapter 9).
+The sharpest known bounds are due to (Talagrand, 2014).
+The expectation of regret defined in (1) is indeed the difference between the expected suprema of two Gaussian processes, $\{\mu_{a}\}_{a\in\mathcal{A}{\operatorname*{full}}}$ and $\{\mu_{a}\}_{a\in\mathcal{A}}$.
+This connection motivates our ideas and analysis around the concept of an $\epsilon$-net, which is a central tool for bounding the expected supremum.
+Appendix B Extra Regret Bounds of Algorithm 1
+We give a matching lower bound:
+Theorem B.1.
+Under the same condition of Theorem 4.5. Let each cluster contains more than one action and satisfies Assumption 4.3. For the same constant $C$ in Theorem 4.2 and another absolute constant $c>0$,
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(\min_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right]-C\epsilon\sqrt{\log m}\right)\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
+The Gaussian process assumption can be relaxed to a sub-Gaussian one, where the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
+$\forall u>0,\quad\Pr\left[|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|_{2}^{2}}\right).$
+Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional (Talagrand, 2014) for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm.
+It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$.
+Using this functional, we establish a regret bound for our algorithm:
+Theorem B.2.
+Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process.
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space.
 Let $\mathcal{A}$ be output of Algorithm 1.
-For the same constant $C>0$ in Theorem 6,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
-Proof sketch.
-If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, choose $a_{\ell}\in r_{\ell}\cap\mathcal{A}$ as the representative point.
-If $r_{\ell}\cap\mathcal{A}=\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$.
-The new set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset.
-Then, for each $\ell\leq m$, define a Gaussian process $\{Z_{a}\}_{a\in r_{\ell}}$, where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and let $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$.
-When $a^{*}(\theta)\in r_{\ell}$, we consider two cases.
-If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, the regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
-If $r_{\ell}\cap\mathcal{A}=\emptyset$, the regret is bounded by $\max_{a\in\mathcal{A}_{\text{full}}}\mu_{a}$.
-$\square$
-The term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ in Theorem 7 is maximized by the uniform $q$, provided $|\mathcal{R}|\geq 2K+1$ (since we can choose any partition); see Supplementary.
-Intuitively, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$.
-This expression is small when $q(r)$ is small, and decays faster when $q(r)$ gets larger, making the overall term negligible if $q$ is highly concentrated.
-The connection between Theorem 6 and 7 is insightful:
-Remark 8.
-Consider placing a grid over the action space and defining a partition $\mathcal{R}$ by assigning each point in the space to its nearest grid point.
-In this way, the grid acts as a reference subset with respect to the partition $\mathcal{R}$.
-The regret upper bound in Theorem 6 applies to this grid, as it holds for any reference subset.
-Meanwhile, Theorem 7 applies to any partition, including $\mathcal{R}$.
-As a result, we obtain a regret bound for Algorithm 1 that exceeds the grid’s bound by only one additional term:
-$\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
-where the part $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]$ is simply a constant, as the action space $\mathcal{A}_{\operatorname*{full}}$ is fixed.
-If the distribution $q$ is highly concentrated, then the expectation $\mathbb{E}_{q}[(1-q(r))^{2K}]$ is small, making the extra term negligible.
-In this case, the benefit of constructing an explicit grid—often a non-trivial task when the action feature vectors are unknown or high-dimensional—is limited.
-Generally speaking, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified.
-We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$:
-$N(\mathcal{A}_{\operatorname*{full}},\!\epsilon)=\min\Big\{m\!\in\!\mathbb{N}:\;\exists\{a_{\ell}\}_{\ell\leq m}\subseteq\mathcal{A}_{\operatorname*{full}},\forall a\in\mathcal{A}_{\operatorname*{full}},\exists\ell,\|a\!-\!a_{\ell}\|_{2}\leq\epsilon\Big\}.$
-Theorem 9.
-Under Assumption 3, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$.
-Let the action space have dimension $n$, and fix a constant $0<\epsilon<M$.
-Let $\mathcal{A}$ be the output of Algorithm 1.
-For the same constant $C>0$ in Theorem 6, and another absolute constant $c>0$,
-$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\text{where }K\geq$ $c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon).$
-As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
-The partition-independent upper bound in Theorem 9 depends on a positive constant $\epsilon$, which serves as a tolerance parameter up to the diameter of the action space.
-Given a $\epsilon$, the required number of samples $K$ scales with the square of the diameter-to-$\epsilon$ ratio, multiplied by the covering number.
-The resulting expected regret is then bounded in terms of $\epsilon$, the dimensionality, and the logarithm of the covering number.
-As $\epsilon$ decreases, more samples are needed, but the regret bound becomes tighter.
-The matching lower bound for Theorem 7, as well as the upper bound under the general sub-Gaussian process assumption for Algorithm 1, are provided in the Supplementary.
-4 Generalization and Empirical Validation
+Then, for a constant $C>0$,
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+Proofs are in Appendices H and I, respectively.
+Appendix C Extra Details of Experiments
 In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings.
 Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling.
-We therefore treat the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set, e.g., $\mathcal{A}_{\operatorname*{full}}:=\{\textnormal{treatment A},\textnormal{treatment B},\textnormal{treatment C}\}$.
+We therefore treat the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set.
 In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$: $f(a):=\mu_{a},\forall a\in\mathcal{A}_{\operatorname*{full}}$.
-We show that our framework applies when the outcome functions lie in a reproducing kernel Hilbert space (RKHS) (williams2006gaussian, ).
+We show that our framework applies when the outcome functions lie in a reproducing kernel Hilbert space (RKHS) (Williams & Rasmussen, 2006).
+Consider a positive semidefinite kernel $k:\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}\to\mathbb{R}$.
 Consider a positive semidefinite kernel $k:\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}\to\mathbb{R}$.
@@
 The full action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{\operatorname*{full}}$.
-Assuming the outcome function is generated from a kernel inherently implies that actions are correlated.
-To study the effect of varying correlation levels among actions, we use the parametric RBF kernel to sample outcome functions from a RKHS: $k(a,a^{\prime}):=\exp\left(-\frac{\|a-a^{\prime}\|^{2}}{2l^{2}}\right)$ , where $l$ is a length-scale parameter that control the dependence among actions.
-By (kanagawa2018gaussian, , Theorem 4.12), we sample functions from a RKHS by first constructing the kernel matrix $\mathbf{K}$ with entries $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, and then drawing $f\sim\mathcal{N}(0,\mathbf{K})$.
-Note that although the kernel is used to generate outcome functions, the kernel itself is not revealed; consequently, $a$ and $a^{\prime}$ are regarded merely as indices rather than as inputs to the kernel.
-Figure 1: Comparison on solving equation 5, selecting $K=5$ distinct actions from 15 grid points over $[0,2]$, with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales.
-Left: Expected regret over 50 repetitions of our method with exhaustive search (green), compared with Thompson Sampling (TS; yellow), Upper Confidence Interval (UCB; orange), and Successive Halving (SH, blue) applied to the set of super-arms.
-TS/UCB are run for 3000 rounds; SH is given a budget of 37,000 pulls.
-Right: SH’s expected regret (blue) and pull counts (yellow bars) per round, compared with our method (green).
-SH requires nearly 10,000 pulls—about three times the number of super-arms—to match our method.
-Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in equation 1.
-Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
-$\max_{\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (5)
-In general, the problem is non-trivial since the objective involves an expectation over random variables with an unspecified distribution, and the inner expression is a maximum over a subset of potentially correlated variables.
-Consequently, small changes in $\mathcal{A}$ can cause non-smooth variations in the maximum.
-Our method is computationally efficient, avoiding the combinatorial complexity.
-We assess performance in identifying near-optimal solutions by selecting $K=5$ actions from an action space of 15 grid points in $[0,2]$.
-The configuration yields a moderate number of super arms, allowing approximation of the optimal subset using Successive Halving (SH), which, however, becomes impractical for larger action spaces.
-Outcome functions are generated with an RBF kernel at length-scales $l\in\{0.5,1,\dots,4\}$.
-We compare the proposed method against Thompson Sampling (TS) and Upper Confidence Bound (UCB).
-In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where exhaustive search is used over the action space to find the optimal action $a^{*}(\theta)$.
-For other methods, we treat each $K$-tuple of actions as a super arm, and the payoff of each super arm in a round $\mathcal{A}$ is given by $\max_{a\in\mathcal{A}}\mu_{a}$.
-Thus, these methods are tailored to find the best super arm that maximizes $\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]$, aligning with the same objective.
-In TS/UCB methods, we adopt a bandit feedback setting: at each round (corresponding to a bandit instance $\theta$), the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}(\theta)$, and updates its policy accordingly, repeating this process for 3000 rounds, chosen to roughly match the number of super arms ($N=3003$).
-Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB, with the prior set to $\mathcal{N}(0,1)$.
-In the SH method, all super arms are evaluated using a fixed budget of arm pulls (we use the minimum budget required in karnin2013almost , $N\log_{2}N\approx 37{,}000$) over a few rounds.
-In each round (corresponding to a bandit instance $\theta$), the number of remaining super arms is halved.
-This process continues until only the best-performing super arm remains or the budget is exhausted.
-We evaluate expected regret of all methods over $10^{5}$ additional randomly sampled outcome functions.
-The left subplot of Figure 1 reports the expected regret of our method (green), TS (yellow), UCB (orange) and SH (blue), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions.
-As shown, the expected regret decreases as the length-scale increases.
-The length-scale $l$ controls the effective number of approximately independent actions: when $a$ and $a^{\prime}$ are far relative to $l$, $k(a,a^{\prime})$ is negligible, so $\mu_{a}$ and $\mu_{a^{\prime}}$ are nearly uncorrelated.
-In the limit, $\{\mu_{a}\}_{a\in\mathcal{A}_{\text{full}}}$ forms a collection of i.i.d.
-zero-mean, unit-variance variables.
-The SH in left subplot of Figure 1 is used to approximate the ground-truth optimal subset and our method does give similar performance as SH.
-The right subplot highlights SH’s impracticality: we ran our method (green) and SH (blue) for 50 repetitions with fixed $l=1$, tracking the current best super arm and total super arm pulls after each SH round.
-The blue curves and shades represent the mean $\pm$ one standard deviation of SH’ expected regret, while yellow bars (corresponding to the right y-axis) show the number of super arm pulls by SH at the end of each round.
-Green curves show our method’s expected regret.
-SH matches our performance after three rounds but requires 9005 pulls—roughly three times the number of super arms—making it intractable for large $\mathcal{A}_{\operatorname*{full}}$.
-Figure 2: Comparison of our method (green), CTS (brown), and CUCB (red) on solving equation 5 with $K=10$ actions chosen from 500 grid points over [5,5][-5,5], with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales.
-In our method, EpsilonNet+TS, the “argmax” step in Algorithm 1 is replaced by TS run for 300 rounds.
-CTS is run for 3000 rounds and CUCB for 2000 rounds.
-All methods use prior variance 11.
-Left: expected regret over 30 repetitions.
-Right: runtime (purple, left axis) and number of arm pulls (gray, right axis) over 30 repetitions.
-Further, we demonstrate the generality of our method (denoted EpsilonNet+TS), where the “argmax” step in Algorithm 1 is replaced by TS.
-We consider the same setup as in Figure 1, but with $K=10$ actions chosen from 500 grid points over [5,5][-5,5].
-In our method, for each bandit instance $\theta$, we run TS for 300 rounds, selecting an action $a$ and observing $\mu_{a}$ in each round; the final action is taken as an approximation of the optimal action $a^{*}(\theta)$.
-We repeat this process until $K=10$ distinct actions are identified.
-We compare against two heuristics: combinatorial TS (CTS) and combinatorial UCB (CUCB).
-CTS is run for 3000 rounds to match the runtime of our method, while CUCB is run for 2000 rounds due to its slower execution and additional initialization step.
-Both CTS and CUCB operate in the semi-bandit feedback setting: in each round (corresponding to one bandit instance $\theta$), an outcome function is sampled and $K=10$ actions $\mathcal{A}$ are selected, with the rewards $\mu_{a}(\theta)$ revealed for all $a\in\mathcal{A}$.
-The prior for TS, CTS, and CUCB is set to $\mathcal{N}(0,1)$, consistent with the ground truth.
-In Figure 2, the left subplot reports the expected regret and its standard deviation for EpsilonNet+TS (green), CTS (brown), and CUCB (red), averaged over 30 repetitions.
-The poor performance of CTS and CUCB is expected, as noted in related works.
-The variance of EpsilonNet+TS could be further reduced by running more TS rounds or using a more accurate action-identification method, since we currently use a Gaussian prior—modeling each action’s expected outcome $\mu_{a}$ as an independent Gaussian random variable—rather than a Gaussian process prior, in order to accommodate the most general case where $\mathcal{A}_{\operatorname*{full}}$ is an index set.
-In the right subplot, we fix the length-scale $l=1$ and evaluate runtime and arm pulls of the three methods over 30 repetitions.
-The left axis (purple) shows average runtime $\pm$ one standard deviation, and the right axis (gray) shows the corresponding number of arm pulls.
-EpsilonNet+TS requires 3000 pulls, since it selects $K=10$ optimal actions, with each action identified by running TS for 300 rounds.
-CTS has similar runtime but needs far more pulls, as it runs 3000 rounds with $K=10$ actions pulled per round.
-Conclusion and Future Work
-We proposed a framework for selecting a subset of correlated actions.
-A simple algorithm was introduced and shown to effectively identify near-optimal subsets.
-Future work could develop a stopping criterion.
-When subset size is flexible but sampling new bandit instances is costly, ideas from species discovery (roswell2021conceptual, ) may be useful—treating the discovery of a new action as analogous to discovering a new species.
-Appendix Summary: The appendix is divided into four parts.
-First, Sections A-B provide additional related works and the regret bound of Algorithm 1, which were omitted from the main body due to page limits.
-Second, Sections C to H contain the proofs of all the theorems and lemmas stated both in the main body and in Section B.
-Third, Sections I-J present auxiliary tools used throughout the paper.
-Fourth, Sections K-M include additional examples and a numerical study of Algorithm 1.
-This part features a low-dimensional ($n=3$) action space on the unit sphere to investigate the effect of cluster diameter on expected regret, in line with Theorem 6.
-It also studies two extensions of the varying dependence structure of actions introduced in the main body:
-  • Nonstationary dependence over the action space: Using a nonstationary Gibbs kernel, where the kernel value depends not only on the difference between two inputs $a,a^{\prime}$ but also on the location of $a$.
-  • Limiting case: The collection of expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ becomes i.i.d.
-LLM Usage: Large language models are used for polishing writing, and finding relevant research.
-Appendix A Extra Related Works
-Epsilon Nets have two standard definitions.
-The first, geometric definition [37], requires that radius-$\epsilon$ balls centered at net points cover the set.
-It relates to the covering number and extends to function classes, as in [31].
-The second, measure-theoretic definition [24], requires the net to intersect all subsets of sufficiently large measure.
-The classic $\epsilon$-net algorithm by [14] remains the simplest and most broadly applicable method.
-Later works aim to reduce net size [26, 27, 25] and address online settings [6].
-Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$.
-It is an important topic in high-dimensional probability [37].
-The sharpest known bounds are due to [34].
-and asymptotic convex geometry [29, Chapter 9]
-Appendix B Extra Regret Bounds of Algorithm 1
-We give a matching lower bound:
-Theorem 10.
-Under the same condition of Theorem 7. Let each cluster contains more than one action and satisfies Assumption 5. For the same constant $C$ in Theorem 6 and another absolute constant $c>0$,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
-In fact, the Gaussian process assumption can be relaxed to a sub-Gaussian one, where the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
-$\forall u>0,\quad\Pr\left[|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|_{2}^{2}}\right).$
-Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional [34] for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm.
-It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$.
-Using this functional, we establish a regret bound for our algorithm:
-Theorem 11.
-Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process.
-Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space.
-Let $\mathcal{A}$ be output of Algorithm 1.
-Then, for a constant $C>0$,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
-Proofs are in Appendices G and H, respectively.
-Appendix C Proof of Lemma 4
+Data Generation:
+Assuming the outcome function is sampled from a kernel implies action correlations. To study varying correlation levels, we use the RBF kernel
+$k(a,a^{\prime}):=\exp\!\left(-\frac{\|a-a^{\prime}\|^{2}}{2l^{2}}\right),$
+where the length-scale $l$ controls dependence among actions.
+Following (Kanagawa et al., 2018) (Theorem 4.12), we sample functions from the RKHS by forming the kernel matrix $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ over $\mathcal{A}_{\operatorname*{full}}$ and drawing $f\sim\mathcal{N}(0,\mathbf{K})$.
+Although the kernel generates the outcomes, it is not revealed to the learner; thus, actions are treated as abstract indices rather than kernel inputs.
+Appendix D Proof of Lemma 3.3
 Statement: Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$.
@@
 where the equality is the probability of missing the cluster $r$ for $K$ times, the first inequality uses the definition of $R_{\epsilon}$.
-The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ [35].
+The last inequality uses $\log(1-\epsilon)^{K}=K\log(1-\epsilon)$ and the inequality $\log(1-\epsilon)\leq-\epsilon$ for $\epsilon<1$ (TOPSØE1, 2007).
 Therefore,
@@
 $=$ $1-|R_{\epsilon}|\exp(-K\epsilon)\geq 1-1/\epsilon\cdot\exp(-K\epsilon),$
-where the first inequality uses union bound, the second inequality uses equation 6.
-The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$.
-∎
-Appendix D Proof of Theorem 6
-Under Assumption 5, the regret can be decomposed as follows; note that this Lemma is used only to derive lower bounds.
-Lemma 12 (Regret decomposition of reference subsets).
+where the first inequality uses union bound, the second inequality uses (12). The last inequality uses the fact that there could be at most $\lfloor 1/\epsilon\rfloor$ clusters in $R_{\epsilon}$. ∎
+Appendix E Proof of Theorem 4.2
+Under Assumption 4.3, the regret can be decomposed as follows; note that this Lemma is used only to derive lower bounds.
+Lemma E.1 (Regret decomposition of reference subsets).
 Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space and an arbitrary reference subset $\mathcal{A}$.
@@
 $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
-Under Assumption 5, $\operatorname*{Regret}=Y_{\ell}$ where the cluster $r_{\ell}$ contains the optimal action. Hence,
+Under Assumption 4.3, $\operatorname*{Regret}=Y_{\ell}$ where the cluster $r_{\ell}$ contains the optimal action. Hence,
 $\mathbb{E}_{\theta}\left[\operatorname*{Regret}\right]=\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}_{\theta}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right].$
@@
 $\operatorname*{Regret}=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell},$
-where the middle equality uses that when $a^{*}(\theta)\in r_{\ell}$, $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\max_{a\in r_{\ell}}\mu_{a}$ and further with Assumption 5, $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\geq\mu_{a_{\ell}}$ when $a^{*}(\theta)\in r_{\ell}$.
+where the middle equality uses that when $a^{*}(\theta)\in r_{\ell}$, $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\max_{a\in r_{\ell}}\mu_{a}$ and further with Assumption 4.3, $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\geq\mu_{a_{\ell}}$ when $a^{*}(\theta)\in r_{\ell}$.
 Then,
@@
 To bound the expected regret of reference subsets, we need a few more lemmas:
-Lemma 13 (Expectation integral identity).
+Lemma E.2 (Expectation integral identity).
 Given a non-negative random variables $X$.
@@
 $=$ $\epsilon\sqrt{\log c}+\frac{\epsilon}{2\sqrt{\log c}}\leq C\epsilon\sqrt{\log c},$
-where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of [37]). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
-Lemma 14 (Borell-TIS inequality; Lemma 2.4.7 of [34]).
+where the first equality uses integrated tail formula of expectation (cf. Lemma 1.6.1 of (Vershynin, 2018)). The last equality set $u_{0}:=\epsilon\sqrt{\log c}$. ∎
+Lemma E.3 (Borell-TIS inequality; Lemma 2.4.7 of (Talagrand, 2014)).
 Given a set $\mathcal{S}$, and a zero-mean Gaussian process $\left(X_{a}\right)_{a\in\mathcal{S}}$.
@@
 $\Pr\left[\left|\sup_{a\in\mathcal{S}}X_{a}-\mathbb{E}\sup_{a\in\mathcal{S}}X_{a}\right|\geq u\right]\leq 2\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
-It means that the size of the fluctuations of $\mathbb{E}\sup_{a\in\mathcal{S}}X_{a}$ is governed by the size of the individual random variables $X_{a}$.
-Statement of Theorem 6:
+It means that the size of the fluctuations of $\sup_{a\in\mathcal{S}}X_{a}$ is governed by the size of the individual random variables $X_{a}$.
+Statement of Theorem 4.2:
 Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$.
 Then, for some constant $C>0$,
-$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
-If the partition $\mathcal{R}$ satisfies Assumption 5, then
-$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
+$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}.$
+If the partition $\mathcal{R}$ satisfies Assumption 4.3, then
+$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log m}.$
 Proof.
@@
 $\operatorname*{Regret}=\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}=Y_{\ell}\leq\max_{\ell^{\prime}\leq m}Y_{\ell^{\prime}},$
-where the first equality follows from the definition of regret in equation 1, and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$.
+where the first equality follows from the definition of regret in (1), and the second equality follows from the assumption that $a^{*}(\theta)\in r_{\ell}$.
 The inequality holds because $a_{\ell}\in\mathcal{A}$, and hence $\max_{a\in\mathcal{A}}\mu_{a}\geq\mu_{a_{\ell}}$.
 The final equality follows from the definition of $Y_{\ell}$.
-On the other hand, according to Lemma 12, under Assumption 5, $\operatorname*{Regret}=Y_{\ell}$, and thus $\operatorname*{Regret}\geq\min_{\ell^{\prime}\leq m}Y_{\ell^{\prime}}$, where the cluster $r_{\ell}$ contains the optimal action.
+On the other hand, according to Lemma E.1, under Assumption 4.3, $\operatorname*{Regret}=Y_{\ell}$, and thus $\operatorname*{Regret}\geq\min_{\ell^{\prime}\leq m}Y_{\ell^{\prime}}$, where the cluster $r_{\ell}$ contains the optimal action.
 Therefore, we reach the important conclusion that
 $\min_{\ell\leq m}Y_{\ell}\leq\operatorname*{Regret}\leq\max_{\ell\leq m}Y_{\ell},$ (13)
-where the left inequality holds under Assumption 5.
+where the left inequality holds under Assumption 4.3.
 Further, by definition $r_{\ell}\subset B(a_{\ell},\epsilon)$, such that $\mathbb{E}Z_{a}^{2}=\mathbb{E}(\mu_{a}-\mu_{a_{\ell}})^{2}=\|a-a_{\ell}\|^{2}_{2}\leq\epsilon^{2}$.
-Using Lemma 14 on the process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$, we have
+Using Lemma E.3 on the process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$, we have
 $\Pr\left[\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq u\right]\leq 2\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
@@
 $\Pr\left[\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq u\right]\leq 2m\exp{\left(-\frac{u^{2}}{2\epsilon^{2}}\right)}.$
-Using Lemma 13, we have for some absolute constant $C>0$
+Using Lemma E.2, we have for some absolute constant $C>0$
 $\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\leq C\epsilon\sqrt{\log m}.$ (14)
@@
 $\mathbb{E}\operatorname*{Regret}\leq\mathbb{E}\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\leq\max_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}+C\epsilon\sqrt{\log m},$ (15)
-where the first inequality uses equation 7.
+where the first inequality uses (13).
 The second inequality uses $\max_{\ell\leq m}Y_{\ell}\leq\max_{\ell\leq m}\mathbb{E}Y_{\ell}+\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\leq\mathbb{E}Y_{\ell}+\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$.
-The last inequality uses equation 8 and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
+The last inequality uses (14) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
 Lower bound:
 $\mathbb{E}\operatorname*{Regret}\geq\mathbb{E}\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\mathbb{E}\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|\geq\min_{\ell\leq m}\mathbb{E}\max_{a\in r_{\ell}}\mu_{a}-C\epsilon\sqrt{\log m},$ (16)
-where the first inequality uses equation 7 under Assumption 5.
+where the first inequality uses (13) under Assumption 4.3.
 The second inequality uses $\min_{\ell\leq m}Y_{\ell}\geq\min_{\ell\leq m}\mathbb{E}Y_{\ell}-\max_{\ell\leq m}\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$, since $Y_{\ell}\geq\mathbb{E}Y_{\ell}-\left|Y_{\ell}-\mathbb{E}Y_{\ell}\right|$.
-The last inequality uses equation 8 and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
+The last inequality uses (14) and the identity $\mathbb{E}[Y_{\ell}]=\mathbb{E}[\max_{a\in r_{\ell}}\mu_{a}]$.
 ∎
-Appendix E Proof of Theorem 7
+Appendix F Proof of Theorem 4.5
 An important consequence of assuming $\mathbb{E}_{\theta}\left[\mu_{a}\right]=0$ for all $a\in\mathcal{A}_{\operatorname*{full}}$ is the following property:
-Lemma 15 (Transformation invariance).
+Lemma F.1 (Transformation invariance).
 Given any vector $x\in\mathbb{R}^{n}$, let $S+x:=\{s+x:s\in\mathcal{S}\}$.
@@
 The key tool for decomposing the regret of Algorithm 1 into that from a reference subset and that from missing clusters is the following Lemma:
-Lemma 16.
+Lemma F.2.
 Consider a partition $\mathcal{R}$ of the full action space.
@@
 Using $\Pr[r\cap\mathcal{A}\!\neq\!\emptyset]\leq 1$, we complete the proof. ∎
-Statement of Theorem 7:
+Statement of Theorem 4.5:
 Let $\mathcal{A}$ be output of Algorithm 1.
 Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
-Then, for the same constant $C>0$ in Theorem 6,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+Then, for the same constant $C>0$ in Theorem 4.2,
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
 Proof.
@@
 If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative.
-The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1.
+The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 4.1.
 The cluster $r_{\ell}$ is contained in a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
@@
 On the other hand, we assume that $0\in\operatorname*{Conv}(\mathcal{A})$ because even if it doesn’t hold, we can always find a vector $x\in\mathbb{R}^{n}$ such that $0\in\operatorname*{Conv}(\mathcal{A}+x)$, without changing the value of $\mathbb{E}[\max_{a\in\mathcal{A}}\mu_{a}]$ (c.f.
-Lemma 15).
+Lemma F.1).
 Therefore,
@@
 where the first equality uses tower rule.
-The first inequality uses Equations (11-12), and Lemma 16.
+The first inequality uses Equations (17-18), and Lemma F.2.
 The last equality uses tower rule again.
 The last inequality uses Cauchy-Schwarz inequality, which states that for any two random variables $X$ and $Y$, we have $|\mathbb{E}[XY]|\leq\sqrt{\mathbb{E}[X^{2}]\mathbb{E}[Y^{2}]}$.
-Note that the Gaussian process assumption is not used in equation 13; it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$.
-By applying equation 9 in the proof of Theorem 6, we bound this term and complete the proof.
+Note that the Gaussian process assumption is not used in (19); it is only needed to bound the term $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$.
+By applying (15) in the proof of Theorem 4.2, we bound this term and complete the proof.
 ∎
-Appendix F Proof of Theorem 9
-To analyze the worst-case behavior of Theorem 7, we first examine the term $\mathbb{E}_{q}\left[(1-q(r))^{K}\right]$ in the upper bound, summarized in the following Lemma:
-Lemma 17.
+Appendix G Proof of Theorem 4.9
+To analyze the worst-case behavior of Theorem 4.5, we first examine the term $\mathbb{E}_{q}\left[(1-q(r))^{K}\right]$ in the upper bound, summarized in the following Lemma:
+Lemma G.1.
 Let $q$ denote a discrete probability distribution over a finite support $\mathcal{R}$. Define
 $\begin{split}M:=\max_{q}&\;\mathbb{E}_{q}\left[(1-q(r))^{K}\right]\\ \textnormal{s.t.}&\sum_{r\in\mathcal{R}}q(r)=1,\;q(r)\in[0,1)\;\;\forall r\in\mathcal{R}.\end{split}$ (20)
-When $|\mathcal{R}|\geq K+1$, the maximum is $M=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$, and it is attained when $q$ is uniform.
-When $|\mathcal{R}|<K+1$, the maximum is upper bounded by $M\leq\frac{|\mathcal{R}|}{K+1}\left(\frac{K}{K+1}\right)^{K}$, and there exists a feasible distribution $q^{\prime}$ such that $\mathbb{E}_{q^{\prime}}\left[(1-q^{\prime}(r))^{K}\right]\geq\frac{|\mathcal{R}|-1}{K+1}\left(\frac{K}{K+1}\right)^{K}$.
+When $m\geq K+1$, the maximum is $M=\left(1-\frac{1}{m}\right)^{K}$, and it is attained when $q$ is uniform.
+When $m<K+1$, the maximum is upper bounded by $M\leq\frac{m}{K+1}\left(\frac{K}{K+1}\right)^{K}$, and there exists a feasible distribution $q^{\prime}$ such that $\mathbb{E}_{q^{\prime}}\left[(1-q^{\prime}(r))^{K}\right]\geq\frac{m-1}{K+1}\left(\frac{K}{K+1}\right)^{K}$.
 Proof.
-Let $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=|\mathcal{R}|$.
+Let $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be the finite support of the measure $q$, where $m:=m$.
 Let $q_{i}:=q(r_{i})$ denote the probability mass at each support point.
@@
 where the right inequality uses that $f(q_{\ell})\geq 0$ for $q_{\ell}\in[0,1]$.
-Case II $m\geq K+1$: Consider the relaxed maximization problem of equation 14:
+Case II $m\geq K+1$: Consider the relaxed maximization problem of (20):
 $\max_{q_{1},\dots,q_{m}\in[0,1)}\sum_{\ell=1}^{m}f(q_{\ell})\quad\text{subject to }\sum_{\ell=1}^{m}q_{\ell}\leq 1.$ (21)
@@
 $(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda,\quad\text{with }\lambda\geq 0.$
-To find critical points of equation 15, we solve the system:
+To find critical points of (21), we solve the system:
 $(1-q_{\ell})^{K-1}(1-(K+1)q_{\ell})=\lambda\geq 0\quad\forall\ell\leq m,\quad\text{and}\quad\sum_{\ell=1}^{m}q_{\ell}\leq 1.$
@@
 Due to the assumption of $m\geq K+1$, we have $\frac{1}{m}\leq\frac{1}{K+1}$, so any choice of $c\in\left[0,\frac{1}{m}\right]$ satisfies the constraint $\sum_{\ell=1}^{m}q_{\ell}\leq 1$ and is feasible for the system.
-Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of equation 15. The corresponding objective value is:
+Therefore, $q_{\ell}=c$ for all $\ell\leq m$ is a feasible critical point of (21). The corresponding objective value is:
 $\sum_{\ell=1}^{m}f(c)=m\cdot c\cdot(1-c)^{K},$
 which is increasing in $c$ over the interval $\left[0,\frac{1}{m}\right]$. Hence, the maximum is attained at $c=\frac{1}{m}$, and the optimal value is:
-$\sum_{\ell=1}^{m}f\left(\frac{1}{m}\right)=\left(1-\frac{1}{m}\right)^{K}=\left(1-\frac{1}{|\mathcal{R}|}\right)^{K},$
+$\sum_{\ell=1}^{m}f\left(\frac{1}{m}\right)=\left(1-\frac{1}{m}\right)^{K}=\left(1-\frac{1}{m}\right)^{K},$
 achieved when $q_{\ell}=\frac{1}{m}$ for all $\ell$.
@@
 which is negative for $q_{\ell}\leq\frac{1}{K+1}$, because then $2-(K+1)q_{\ell}>0$. Hence, the Hessian is negative definite, and the critical point is a local (and thus global) maximum.
-Finally, note that equation 15 is a relaxation of equation 14.
+Finally, note that (21) is a relaxation of (20).
 While the optimum of the original problem is upper bounded by that of the relaxed problem, the optimal solution to the relaxed problem also lies within the feasible region of the original problem.
-Therefore, the maximum of equation 14 is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}$.
+Therefore, the maximum of (20) is also attained when $q$ is uniform, with the maximum value being $\left(1-\frac{1}{m}\right)^{K}$.
 ∎
-Statement of Theorem 9
-Under Assumption 3, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$.
+Assumption G.2.
+The support of measure $q$ is compact.
+The compactness assumption ensures that the term $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\right]$ is finite and guarantees the attainment of a unique optimal action.
+Without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}$ is the support of the measure $q$.
+Statement of Theorem 4.9
+Under Assumption G.2, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$.
 Let the action space have dimension $n\in\mathbb{N}$, and fix a constant $0<\epsilon<M$.
 Let $\mathcal{A}$ be the output of Algorithm 1.
-For the same constant $C>0$ in Theorem 6, we have:
-$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\quad\text{where }K\geq\frac{1}{2}\left(\frac{M^{2}N(\mathcal{A}_{\operatorname*{full}},\epsilon)}{\epsilon^{2}e}-1\right).$
+For the same constant $C>0$ in Theorem 4.2, we have:
+$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\quad\text{where }K\geq\frac{1}{2}\left(\frac{M^{2}N(\mathcal{A}_{\operatorname*{full}},\epsilon)}{\epsilon^{2}e}-1\right).$
 As $\epsilon\to 0^{+}$, we have $\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\to 0$ as $K\to\infty$.
 Proof.
-By Lemma 15, we can shift the action space to be centered at the origin.
+Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{\operatorname*{full}}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{\operatorname*{full}},\epsilon)$.
+Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
+$r_{\ell}=\left\{a\in\mathcal{A}_{\operatorname*{full}}:\pi(a)=a_{\ell}\right\}.$
+Step 1 bounding by dimension $n$:
+By Lemma F.1, we can shift the action space to be centered at the origin.
 So, without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}\subset M\cdot B_{2}^{n}$, the scaled unit Euclidean ball in $\mathbb{R}^{n}$.
@@
 where the inequality follows from $\mu_{a}=\langle\theta,a\rangle\leq M\|\theta\|_{2}$ for all $a\in MB_{2}^{n}$, and the equality uses that $\theta\sim\mathcal{N}(0,I)$, so each coordinate has variance 1 and $\|\theta\|_{2}^{2}=\sum_{i=1}^{n}\theta_{i}^{2}$ has expectation $n$.
-Let $\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{\operatorname*{full}}$ be a minimal geometric $\epsilon$-net under the Euclidean norm, so that $m=N(\mathcal{A}_{\operatorname*{full}},\epsilon)$.
-Define $\pi(a)$ as the closest point in the $\epsilon$-net to $a$, and let the partition $\mathcal{R}=\{r_{1},\dots,r_{m}\}$ be given by
-$r_{\ell}=\left\{a\in\mathcal{A}_{\operatorname*{full}}:\pi(a)=a_{\ell}\right\}.$
 Then,
 $\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}\right]\leq\mathbb{E}_{\theta}\left[\max_{a\in B(a_{\ell},\epsilon)}\mu_{a}\right]=\mathbb{E}_{\theta}\left[\max_{a\in\epsilon B_{2}^{n}}\mu_{a}\right]\leq\epsilon\sqrt{n},$ (23)
-where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma 15 to shift; and the final bound uses Claim 3 from Supplementary I. Then,
+where the first inequality follows from $r_{\ell}\subset B(a_{\ell},\epsilon)$ by construction; the equality uses Lemma F.1 to shift; and the final bound uses Property 3 from Supplementary J.
+Step 2 bounding by cardinality $|\mathcal{A}_{\operatorname*{full}}|$:
+By the definition of variance:
+$\begin{split}\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}^{2}\right]=&\left(\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\right]\right)^{2}+\text{Var}\left(\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\right)\\ \leq&\left(\frac{\operatorname{diam}(\mathcal{A}_{\operatorname*{full}})}{2}\sqrt{2\log|\mathcal{A}_{\operatorname*{full}}|}\right)^{2}+4\operatorname{diam}(\mathcal{A}_{\operatorname*{full}})^{2}\\ =&M^{2}\left(\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\right)^{2},\end{split}$ (24)
+where the inequality uses Supplementary K to bound the square of Gaussian width, and Lemma E.3 to bound the variance. The last equality uses the definition of $M$ and simple rearrangement.
+Also,
+$\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\mu_{a}\right]\leq\max_{\ell\leq m}\frac{\operatorname{diam}(r_{\ell})}{2}\sqrt{2\log|r_{\ell}|}\leq\frac{\epsilon}{2}\sqrt{2\log|\mathcal{A}_{\operatorname*{full}}|},$ (25)
+where the left inequality uses Supplementary K, the right inequality uses the definition of $\epsilon$.
+Then,
 $\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq$ $\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log m}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
-$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}\cdot M\sqrt{n}$
-$\leq$ $\epsilon\sqrt{n}+C\epsilon\sqrt{\log m}+\epsilon\sqrt{n},$
-where the first inequality follows from the algorithm-dependent upper bound in Theorem 7.
-The second inequality follows from Equations equation 16 and equation 17, and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma 17).
-The last inequality comes from
+$\leq$ $\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}\}+C\epsilon\sqrt{\log m}+\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}M\;\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\}$
+$\leq$ $\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}\}+C\epsilon\sqrt{\log m}+\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\}$
+$\leq$ $2\epsilon\min\{\sqrt{n},\sqrt{\frac{1}{2}\log|\mathcal{A}_{\operatorname*{full}}|}+4\}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},$
+where the first inequality follows from the algorithm-dependent upper bound in Theorem 4.5.
+The second inequality follows from Equations (22)-(25), and the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ is bounded by $\frac{m}{2K+1}\left(\frac{2K}{2K+1}\right)^{2K}$ (see Lemma G.1).
+The third inequality comes from
 $\log\left(\sqrt{\frac{m}{2K+1}}\left(\frac{2K}{2K+1}\right)^{K}M\right)=$ $\tfrac{1}{2}\log\left(\tfrac{m}{2K+1}\right)+K\log\left(\tfrac{2K}{2K+1}\right)+\log M$
@@
 The last inequality uses the assumption $K\geq\frac{1}{2}\left(\frac{mM^{2}}{\epsilon^{2}e}-1\right)$.
-Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of [37].
+Finally, since $\mathcal{A}_{\operatorname*{full}}\subset MB_{2}^{n}$, the covering number satisfies $N(\mathcal{A}_{\operatorname*{full}},\epsilon)\leq C^{\prime}(M/\epsilon)^{n}$ for some constant $C^{\prime}>0$; see Proposition 4.2.12 of (Vershynin, 2018).
 Since $\epsilon\sqrt{\log(M/\epsilon)}\to 0$ as $\epsilon\to 0^{+}$, it follows that $\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)}\to 0$ as $\epsilon\to 0^{+}$.
 ∎
-Appendix G Proof of Theorem 10
-To derive a matching lower bound, we decompose the regret of Algorithm 1 into that from a reference subset and from missing clusters, using Lemma 16:
-Statement of Theorem 10
-Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 5.
+Appendix H Proof of Theorem B.1
+To derive a matching lower bound, we decompose the regret of Algorithm 1 into that from a reference subset and from missing clusters, using Lemma F.2:
+Statement of Theorem B.1
+Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 4.3.
 Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
 Let $\mathcal{A}$ be the output of Algorithm 1.
-Then, for the same constant $C$ in Theorem 6 and another constant $c>0$, for any reference subset $\mathcal{A}^{\prime}$:
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).\end{split}$
+Then, for the same constant $C$ in Theorem 4.2 and another constant $c>0$, for any reference subset $\mathcal{A}^{\prime}$:
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log m}\right).\end{split}$
 Proof.
-Choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative, such that the subset $\mathcal{A}^{\prime}=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1.
+Choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative, such that the subset $\mathcal{A}^{\prime}=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 4.1.
 Fix $\ell$.
-As in Lemma 12, define a Gaussian process $(Z_{a})_{a\in r_{\ell}}$ by setting $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and define random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$.
+As in Lemma E.1, define a Gaussian process $(Z_{a})_{a\in r_{\ell}}$ by setting $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and define random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$.
 The key idea is that $\operatorname*{Regret}\geq Y_{\ell}$ whenever $r_{\ell}\cap\mathcal{A}=\emptyset$ and $a^{*}(\theta)\in r_{\ell}$, such that
 $\begin{split}\mathbb{E}\left[\operatorname*{Regret}\Big|r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]=&\mathbb{E}\left[\max_{a\in r_{\ell}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\Big|r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ \geq&\mathbb{E}\left[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\Big|r_{\ell}\cap\mathcal{A}\!=\!\emptyset,a^{*}(\theta)\in r_{\ell}\right]\\ =&\mathbb{E}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right],\end{split}$ (26)
-where the inequality follows from Assumption 5. The last equality follows from $Y_{\ell}$ is independent from $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
-$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\right]=&\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]\\ &+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]\\ \geq&\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right],\\ \geq&\sum_{\ell\leq m}q(r_{\ell})\cdot(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}^{2}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right]\right)^{1/2}\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\end{split}$
+where the inequality follows from Assumption 4.3. The last equality follows from $Y_{\ell}$ is independent from $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$. Further,
+$\begin{split}\mathbb{E}\left[\operatorname*{Regret}\right]=&\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}\neq\emptyset,a^{*}(\theta)\in r\right]\\ &+\sum_{r\in\mathcal{R}}\Pr[r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r]\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right]\\ \geq&\sum_{r\in\mathcal{R}}q(r)\cdot(1-q(r))^{K}\cdot\mathbb{E}\left[\operatorname*{Regret}\Big|r\cap\mathcal{A}=\emptyset,a^{*}(\theta)\in r\right],\\ \geq&\sum_{\ell\leq m}q(r_{\ell})\cdot(1-q(r_{\ell}))^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}^{2}\left[Y_{\ell}\Big|a^{*}(\theta)\in r_{\ell}\right]\right)^{1/2}\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]\\ \geq&c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log m}\right)\end{split}$
 where the equality uses the tower rule.
-The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 16.
-The second inequality follows from equation 18.
-The third inequality uses the Pólya-Szegő inequality [11], corresponding to a constant $c>0$:
+The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma F.2.
+The second inequality follows from (26).
+The third inequality uses the Pólya-Szegő inequality (Dragomir, 2004), corresponding to a constant $c>0$:
 $c:=\frac{1}{2}\left(\sqrt{\frac{M_{1}M_{2}}{m_{1}m_{2}}}+\sqrt{\frac{m_{1}m_{2}}{M_{1}M_{2}}}\right),$
@@
 Note that $m_{2}$ is positive due to the assumption that each cluster contains more than one action.
-The fourth inequality follows from Jensen’s inequality that for a random variable $X$, $\mathbb{E}[X^{2}]\geq\mathbb{E}[X]^{2}$, and Lemma 12.
+The fourth inequality follows from Jensen’s inequality that for a random variable $X$, $\mathbb{E}[X^{2}]\geq\mathbb{E}[X]^{2}$, and Lemma E.1.
 Since the reference subset $\mathcal{A}^{\prime}$ is arbitrary, the lower bound $c\cdot\left(\mathbb{E}_{q}(1-q(r))^{2K}\right)^{1/2}\cdot\mathbb{E}_{\theta}\left[\operatorname*{Regret}\textnormal{of }\mathcal{A}^{\prime}\right]$ holds for any reference subset.
-The last inequality uses Theorem 6.
+The last inequality uses Theorem 4.2.
 ∎
-Appendix H Proof of Theorem 11
-Definition 2.
+Appendix I Regret Bound for Sub-Gaussian Processes
+The Gaussian process assumption can be relaxed to a sub-Gaussian one, where the linear structure in Equation (5) is not required, and the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
+$\forall u>0,\;\Pr\left[|\mu_{a}\!-\!\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{u^{2}}{2\|a\!-\!a^{\prime}\|_{2}^{2}}\right).$ (27)
+Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s $\gamma_{2}$-functional (Talagrand, 2014) for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm.
+It is known that for any sub-Gaussian process indexed by $\mathcal{S}$, there exists a universal constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\,\gamma_{2}(\mathcal{S}).$
+Using this functional, we establish the following regret bound for our algorithm:
+Theorem I.1.
+Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\mathrm{full}}}$ be a mean-zero sub-Gaussian process satisfying Equation (27). Let $\mathcal{A}$ be the output of Algorithm 1. Then, for some constant $C>0$,
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\mathrm{Regret}]\leq\;C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\mathrm{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+Definition I.2.
 The random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a mean-zero sub-Gaussian process if the process $\mathbb{E}\mu_{a}=0$ and has the increment condition:
 $\forall u>0,\Pr[|\mu_{a}-\mu_{a^{\prime}}|\geq u]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|^{2}_{2}}\right).$
-Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional [34] for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm.
+Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional (Talagrand, 2014) for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm.
 It is well-known that under the assumption that $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a Gaussian process, this gives the tightest bound, for a universal constant $L$:
@@
 Within the proof of the upper bound for general sub-Gaussian process, it also derives the deviation bound for the term $\max_{a\in\mathcal{S}}\mu_{a}$:
-Lemma 18 (Theorem 2.2.22 of [34]).
+Lemma I.3 (Theorem 2.2.22 of (Talagrand, 2014)).
 Let $\{\mu_{a}\}_{a\in\mathcal{S}}$ be a mean-zero sub-Gaussian process, then there exists a constant $c>0$:
 $\Pr\left[\sup_{a,a^{\prime}\in\mathcal{S}}|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{cu^{2}}{\gamma^{2}_{2}(\mathcal{S})}\right).$
-Lemma 19.
+Lemma I.4.
 Let $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ be a mean-zero sub-Gaussian process.
@@
 $\left|\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right|\leq\sup_{a\in r_{\ell}}|\mu_{a}-\mu_{a_{\ell}}|\leq\sup_{a,a^{\prime}\in r_{\ell}}|\mu_{a}-\mu_{a^{\prime}}|.$
-The second inequality uses the fact that $\{\mu_{a}\}_{a\in r_{\ell}}$ is a mean-zero sub-Gaussian process and applies Lemma 18.
+The second inequality uses the fact that $\{\mu_{a}\}_{a\in r_{\ell}}$ is a mean-zero sub-Gaussian process and applies Lemma I.3.
 The last inequality uses the definition of $\epsilon$, and $c>0$ is a constant.
@@
 $\Pr\left[\max_{\ell\leq m}|Y_{\ell}|\geq u\right]\leq 2m\exp\left(-\frac{cu^{2}}{\epsilon^{2}}\right).$
-Further, using Lemma 13, we have for some absolute constant $C>0$:
+Further, using Lemma E.2, we have for some absolute constant $C>0$:
 $\mathbb{E}\left[\max_{\ell\leq m}|Y_{\ell}|\right]\leq C\epsilon\sqrt{\log m}.$
-Statement of Theorem 11:
+Statement of Theorem B.2:
 Let $\mathcal{A}$ be output of Algorithm 1.
@@
 Then, for the same constant $C>0$,
-$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
+$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
 Proof.
@@
 If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative.
-The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1.
+The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 4.1.
 For each $\ell\leq m$, define a random variable
 $Y_{\ell}:=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
-Following the same reasoning used in the proof of Theorem 7:
+Following the same reasoning used in the proof of Theorem 4.5:
 $\mathbb{E}[\operatorname*{Regret}]\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}$
 $\leq$ $C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
-The first inequality follows from equation 13 in the proof of Theorem 7. The last inequality uses Lemma 19. ∎
-Appendix I Properties of Gaussian width
+The first inequality follows from (19) in the proof of Theorem 4.5. The last inequality uses Lemma I.4. ∎
+Appendix J Properties of Gaussian width
 Given a set $\mathcal{S}\in\mathbb{R}^{n}$, the term $\mathbb{E}[\max_{a\in\mathcal{S}}\mu_{a}]$ where $\theta\sim\mathcal{N}(0,I)$ is called Gaussian (mean) width.
-Claim 1: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
+Property 1: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right]$.
 $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,-\theta\rangle\right]=\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle-a,\theta\rangle\right]=\mathbb{E}\left[\max_{a^{\prime}\in-\mathcal{S}}\langle a^{\prime},\theta\rangle\right],$
 where the first equality uses $\theta$ and $\theta$ are identically distributed. The third equality uses for any $a\in\mathcal{S}$, it holds that $a\in-\mathcal{S}$.
-Claim 2: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]\leq\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$.
+Property 2: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]\leq\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$.
 Let $a^{*}(-\mathcal{S},\theta)$ denote the optimal action in $\mathcal{S}$ for bandit instance $\theta$.
@@
 $\leq$ $\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right],$
-where the first equality uses Claim 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$.
+where the first equality uses Property 1, the second equality uses the definition of $a^{*}(-\mathcal{S},\theta)$.
 The third equality uses linearity of expectation.
@@
 In fact, one can prove the equality that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]=\frac{1}{2}\mathbb{E}\left[\max_{a,a^{\prime}\in\mathcal{S}}\langle a-a^{\prime},\theta\rangle\right]$, but we only need inequality to prove the next claim.
-Claim 3: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n}$.
+Property 3: $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq\frac{\operatorname{diam}(\mathcal{S})}{2}\cdot\sqrt{n}$.
 $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]=$ $\mathbb{E}\left[\max_{a\in\mathcal{S}}\langle a,\theta\rangle\right]$
@@
 where the first equality uses the definition of $\mu_{a}$.
-The first inequality uses Claim 2.
+The first inequality uses Property 2.
 The second inequality uses Cauchy-Schwarz inequality.
@@
 The last inequality uses $\mathbb{E}\|\theta\|_{2}\leq\sqrt{n}$.
-Appendix J Bounds of expected maximum of Gaussian
+Appendix K Bounds of expected maximum of Gaussian
 Let $X_{1},\dots,X_{N}$ be $N$ random Gaussian variables (no necessarily independent) with zero mean and variance of marginals smaller than $\sigma^{2}$, then
@@
 where the first inequality uses Jensen’s inequality. Taking $\delta:=\sqrt{2(\log N)/\sigma^{2}}$ yields the results. ∎
-Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [17]:
+Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to (Kamath, 2015):
 $\mathbb{E}_{\theta}\left[\max_{i=1,\dots,N}X_{i}\right]\geq\frac{\sigma\sqrt{\log N}}{\sqrt{\pi\log 2}}.$
-Appendix K The effect of clustering structure on regret
+Appendix L The effect of clustering structure on regret
 In Figure 4, we study the effect of the cluster diameters (controlled by a spread parameter) on regret.
@@
 The middle and right subplots show example action spaces (blue dots) for spread values 0.01 and 0.5, with representative actions (purple stars) selected by Algorithm 1 with $K=10$.
-Appendix L Varying-dependence actions with RBF/Gibbs kernels
-We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel [39].
+Appendix M Varying-dependence actions with RBF/Gibbs kernels
+We study the effect of varying action dependence using a Gaussian process with a kernel.
+To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel (Williams & Rasmussen, 2006).
 $\begin{split}\underset{\text{RBF}}{k}(a,a^{\prime})=&\exp\left(-\frac{\|a-a^{\prime}\|^{2}}{2l^{2}}\right),\\ \underset{\text{Gibbs}}{k}(a,a^{\prime})=&\sqrt{\frac{2\,l(a)l(a^{\prime})}{l(a)^{2}\!+\!l(a^{\prime})^{2}}}\exp\left(\!-\!\frac{\|a-a^{\prime}\|^{2}}{l(a)^{2}\!+\!l(a^{\prime})^{2}}\right),\end{split}$ (28)
@@
 We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$.
-Under either kernels defined in equation 19, the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{\operatorname*{full}}$.
-In this way, we sample functions from a RKHS function class; See [18, Theorem 4.12].
-Figure 4: Experiments with outcome functions sampled from RBF/Gibbs kernels in equation 19.
+Under either kernels defined in (28), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{\operatorname*{full}}$.
+In this way, we sample functions from a RKHS function class; See (Kanagawa et al., 2018, Theorem 4.12).
+Figure 5: Experiments with outcome functions sampled from RBF/Gibbs kernels in (28).
 Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in $[0,2]$ (blue curves, right y-axis).
 The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points.
-To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval $[0,2]$, using Gibbs kernel in equation 19 to sample outcome functions.
+To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval $[0,2]$, using Gibbs kernel in (28) to sample outcome functions.
 To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid.
@@
 The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
-Appendix M Independent and identically distributed actions
+Appendix N Independent and identically distributed actions
 As an extreme case, suppose that $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ is a set of i.i.d.
@@
 $\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\mathbb{E}_{\theta}\max_{i=1,\dots,n}\theta_{i}=\mathbb{E}\max_{i=1,\dots,n}X_{i},$
-where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary J, we have
+where the equivalence comes from that each entry $\theta_{i}$ is i.i.d. sample from $\mathcal{N}(0,1)$. Then, according to the bounds of expected maximum of Gaussian in Supplementary K, we have
 $\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\sqrt{2\log n},\quad\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}}\mu_{a}\leq\sqrt{2\log|\mathcal{A}|}.$
-Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in equation 1,
+Bounds of expected regret of arbitrary $\mathcal{A}$: By definition of regret in (1),
 $\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}\leq\mathbb{E}_{\theta}[\textnormal{Regret}]\leq\sqrt{2\log n}-\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}.$